### v1

# From shape to fate: making bacterial swarming expansion predictable

Abstract
Microbial swarming on mucosal surfaces reshapes microbial communities and influences mucosal healing and antibiotic tolerance. Yet even with time-lapse microscopy and deep learning, analyses of swarming colonies remain descriptive and cannot forecast how their fronts reorganize in time. This limitation is significant because the advancing edge determines access to nutrients, host tissue and competing microbes. We recast the expansion of Enterobacter sp. SM3 swarms as a problem of morphological forecasting, and assemble SwarmEvo, a time-lapse dataset represented as boundary-resolved segmentations. TexPol–Net, a texture- and geometry-aware segmentation model, sharpens diffuse edges and preserves fingered fronts, creating a stable substrate for dynamics. On this representation, we develop Morpher, an autoregressive forecasting network with a “Morphon” memory that links local curvature to long-range temporal dependencies. Morpher outperforms leading video-prediction models in maintaining front localization and anisotropic branching, and modest segmentation improvements yield noticeably more stable forecasts. Ablations across sequence models, inference strategies and observation ratios show that attention-based architectures with structural memory best preserve dense-finger propagation. By uniting geometry-aware segmentation with morphology-level forecasting, this framework turns swarming expansion into a predictive dynamical system, enabling quantitative interrogation and potential control of microbial collectives during mucosal repair and gut ecosystem engineering.
1 Introduction
Microbial communities form dense and continually reorganizing ecological networks within and around animal hosts [27, 7, 2, 50, 38, 36]. Their spatial organization and collective movement govern population expansion, interspecies interactions and tissue homeostasis under both healthy and diseased conditions [17, 26, 22]. Among these behaviors, collective surface motility on semi–solid substrates—classically termed swarming [1, 5]—produces continually reconfiguring fronts and patterned colony architectures that both reflect coordinated cellular behaviors and reshape local host environments [11, 49, 32]. Yet despite extensive descriptive work, the spatiotemporal evolution of such structures remains poorly characterized. In particular, forecasting the future morphology of a swarming colony—one of the most dynamic forms of bacterial collective motion—is currently out of reach. A predictive framework that links present colony architecture to its future configurations would transform morphology from a descriptive endpoint into a quantitatively interrogable variable, enabling mechanistic studies of how microbial collectives respond to environmental and physiological cues.
Attempts to analyze colony growth have historically been constrained by low imaging frequency and qualitative interpretation [24, 28]. Classical computer–vision pipelines partially automated colony detection and quantification [4, 9, 37, 51], but they tend to fail when boundaries blur, textures reorganize or colonies overlap. More recent analyses of active suspensions and deformable colony–fluid interfaces have revealed turbulence–like flows and curvature–dependent edge dynamics [18, 48]. However, these approaches often presume approximate symmetry or reduce colony evolution to low–dimensional order parameters, limiting their applicability to the irregular, anisotropic and burst–like propagation characteristic of swarming fronts. In parallel, deep learning has expanded biological image analysis, yielding convolutional neural network (CNN)–based colony detectors [14, 47], temporal classifiers of motility states [31] and hybrid pipelines that combine growth–curve inference with object detection [29]. Advances in imaging, such as coherent time–lapse microscopy for early species identification [43] and engineered swarming biosensors [12], further underscore the morphological richness encoded in colony structure. Yet these developments remain essentially descriptive: they identify colonies, classify motility or extract bulk parameters, but they do not predict how a colony’s contour will evolve. Even models trained to infer motility type from single frames [23] collapse inherently dynamic processes into static labels and therefore offer limited insight into future behavior.
A central obstacle is that swarming colonies violate the appearance continuity implicitly assumed by most natural–scene video models. Fronts propagate through intermittent bursts, transient asymmetries and rapid reorganizations of multi–scale texture. Curvature modulates local velocity; protrusions emerge and retract discontinuously; textures shift without preserving pixel–level coherence. Under these conditions, extrapolating visual appearance is neither stable nor biologically grounded. Predicting future morphology instead requires reasoning over geometric and biophysical representations of the colony boundary and its local environment. Such forecasts would enable in silico perturbation experiments, real–time anticipation of responses to surface biochemistry or antimicrobial stress, and strategies to steer collective migration in applications ranging from mucosal repair to microbial therapeutics and microbiome engineering. Predictions of front position and branching patterns implicitly reflected experimental responses to mucin gradients and antimicrobial stress, indicating that the learned representations encode biologically meaningful adaptation rather than merely geometric similarity.
These questions crystallize in Enterobacter sp. SM3, a representative swarming commensal from the murine gut [33, 8]. SM3 reshapes intestinal microbial organization and promotes mucosal healing, whereas mutants impaired in swarming lose these beneficial functions [11]. More broadly, swarming has been implicated in antibiotic tolerance, regulation of virulence factors and robust colonization of host–associated surfaces [21, 30], with its dynamics strongly modulated by surface biochemical cues such as mucin [32]. In inflamed intestinal environments, the advancing swarming front is the actionable interface: its future position specifies where oxygen is depleted and anaerobic niches emerge, shaping downstream microbial reorganization and recovery, as shown in Fig. 1. Anticipating this position in advance therefore provides a principled basis for spatially and temporally targeted intervention, through pharmacological or environmental modulation of the local microenvironment. Despite this conceptual clarity, translating it into predictive capability remains an open challenge. No framework currently exists for forecasting the evolving contours of swarming colonies at the resolution of individual fronts. This gap limits our ability to quantify how environmental or physiological signals shape collective migration and to connect microscopic motility programs to macroscopic patterns in a predictive manner.
Here we address this gap by formulating swarming colony expansion as a problem of morphological forecasting in a geometric state space. We first assemble the Swarming Morphogenesis Evolution (SwarmEvo) dataset, a curated collection of high–resolution time–lapse recordings of Enterobacter sp. SM3 spreading across semi–solid substrates under systematically varied conditions. Each recording is processed into boundary–resolved segmentation sequences that capture the evolving colony contour, providing both a biologically faithful account of swarming dynamics and a quantitative benchmark for evaluating whether computational models can maintain geometric fidelity over time. Building on this representation, we develop TexPol–Net, a segmentation framework that integrates texture- and geometry-aware priors to recover high–fidelity fronts even when boundaries are diffuse or heavily reorganized, thereby establishing a stable morphological state space for temporal prediction. We then introduce Morpher, a spatiotemporal model that treats forecasting as contour evolution within this state space, combining local geometric cues with long–range temporal dependencies to capture both smooth curvature modulations and rapid reorganization of protrusions.
By establishing a geometric representation of colony expansion and a forecasting model capable of operating within it, SwarmEvo, TexPol–Net and Morpher together recast swarming from a descriptive phenomenon into a tractable dynamical system with predictable trajectories. This framework enables quantitative interrogation of how microbial collectives respond to environmental and physiological cues with a temporal resolution beyond the reach of current experimental tools, and it lays an analytical foundation for integrating morphological forecasting with real-time imaging and for interpreting and steering microbial collective behavior in vivo.
2 Results
2.1 A two-stage formulation of swarming morphology
The swarming expansion of Enterobacter sp. SM3 is characterized by pronounced morphological plasticity. Under identical assay conditions, colonies may evolve through irregular, dense finger-like protrusions driven by curvature instabilities at the advancing front, or settle into near-concentric growth characterized by a stabilized rim and weak directional anisotropy. These recurring regimes define a structured spectrum of morphogenetic behavior, spanning fine-scale boundary fluctuations and colony-scale organization (Fig. 1).
Figure 1: From swarming dynamics to predictive guidance. In a mouse model of colitis, Enterobacter sp. SM3 swarms along the inflamed mucosal surface. Under pharmacological, environmental, or microbiota-based modulation, forecasting the future position of the active swarming front may enable spatially targeted and temporally coordinated interventions. Such predictive capability could be coupled with experimental strategies that locally modulate oxygen availability or microbial competition at the advancing front, with potential effects on anaerobic niche formation and downstream community composition, including Muribaculaceae expansion. Time-lapse swarming assays are converted into boundary-resolved morphological masks by TexPol–Net, with Texture-Edge and Polar–Context attention, shared prototypes and mask assembly, defining observed morphological states. Morpher models their temporal evolution with a sequence backbone and Morphon memory to predict future morphologies for quantitative evaluation. Panels I–IV depict an application-oriented biological and experimental pipeline, whereas the modeling framework follows a two-stage formulation of boundary measurement and morphology-level forecasting.
To resolve these dynamics, we adopt a two-stage formulation that separates morphological measurement from temporal prediction. In Panel III (Fig. 1), time-lapse microscopy is converted into sequences of boundary-resolved morphological masks using TexPol–Net. Rather than serving as a generic segmentation backbone, TexPol–Net acts as a measurement operator that constrains the geometric degrees of freedom of the colony front, preserving thin protrusions and internal texture while suppressing illumination-induced variability. The resulting mask sequences constitute a compact and biologically interpretable representation of colony state across time.
Panel IV (Fig. 1) treats these observed morphologies as a temporal signal and models their evolution using Morpher. Operating directly in morphology space, Morpher infers latent dynamical states from the observed window and extrapolates future colony shapes beyond it. This formulation shifts the forecasting task from image synthesis to the prediction of boundary evolution, allowing local perturbations along the front to propagate coherently over long horizons. Autoregressive conditioning further permits predicted morphologies to influence subsequent growth states, reflecting the cumulative nature of swarming expansion.
By decoupling spatial characterization from temporal dynamics, the framework limits error accumulation from appearance noise and isolates the mechanisms governing shape evolution. This structure supports systematic evaluation of forecasting models across distinct morphogenetic regimes, from highly irregular fingered fronts to stabilized radial expansion, within a unified representation that preserves biologically relevant geometric structure.
2.2 Recovering colony fronts for accurate morphological forecasting
Accurate delineation of the advancing front is the prerequisite for morphological forecasting, as it encodes curvature, anisotropy, and the local direction of expansion. We therefore evaluated how different segmentation backbones reconstruct colony fronts across representative growth regimes, and quantitatively examined how boundary accuracy evolves as matching criteria become increasingly stringent (Fig. 3).
In the dense–finger regime, the front is highly corrugated and places stringent demands on boundary fidelity (Fig. 3b). YOLOv11 is able to follow the global outline, but tends to break slender fingers into shortened or discontinuous segments, effectively smoothing high-curvature protrusions and weakening the anisotropic structure of the advancing front. SAM-based models instead emphasize the central colony mass and suppress distal protrusions, causing the boundary to evolve toward a smoother, near-circular shape. In contrast, TexPol–Net remains aligned with the annotated front along both the outer envelope and fine branches, preserving uneven finger lengths and the radial bias characteristic of early expansion; YOLOv12 is the closest competitor in this regime.
[…]
Figure 3: TexPol–Net improves colony-front segmentation by coupling texture-sensitive boundary encoding with a geometry-aligned context prior. a, Texture-Polar Network (TexPol–Net) within a prototype-based instance segmentation pipeline. Texture-Edge Attention (TEA) is embedded in the backbone to preserve high-frequency boundary cues, while Polar–Context Attention (PCA) is interleaved in the PANet-style bidirectional neck to maintain polar consistency during multi-scale fusion. Dense heads at P3P3P5P5 predict class scores, boxes and mask coefficients; a lightweight Protonet generates k=32k{=}32 shared prototypes that are linearly combined into instance masks and post-processed by cropping and thresholding. b, Qualitative comparison on two representative swarming regimes. In the dense–finger phase (top), YOLOv11 [19] captures the coarse outline but smooths the corrugated front, shortening narrow fingers and reducing anisotropy. SAM [20] and SAM2 [34] favor the colony core, producing a boundary that approaches a near-circular shape. TexPol–Net preserves finger heterogeneity and radial bias, with YOLOv12 [41] as the closest competing backbone. In the near–concentric ring regime (bottom), all methods recover the circular structure, while TexPol–Net more precisely localizes the advancing boundary. c, Average precision as a function of IoU matching threshold, highlighting boundary-stringent performance at high IoU. d, Image-wise overlap statistics for IoU and Dice, summarizing accuracy and variability across images.
Accurate recovery of the advancing front is a structural prerequisite for morphological forecasting rather than a superficial preprocessing step. Errors introduced at the segmentation stage propagate directly into the temporal model, as smoothing finger-like protrusions, displacing the outer rim, or introducing spurious fragments alters the effective front trajectories available for learning. By preserving fine-scale protrusions while maintaining the global organization of the rim, TexPol–Net yields a geometrically consistent sequence of morphological masks in which successive fronts are linked through physically plausible deformations. This representation defines an appropriate input substrate for curvature-driven expansion and underpins the forecasting analyses presented in the following sections.
2.3 Forecasting the expanding front of swarming colonies
We conduct a systematic comparison between Morpher and representative state–of–the–art video prediction models, including MAU [6], MIM [46], PredRNN [44], PredRNNv2 [45], the original SimVP implementation which employs the TAU temporal unit [15, 40], and the improved SimVPv2 variant, in which TAU is replaced by the gSTA module [39]. These models have demonstrated strong performance on standard video prediction benchmarks and are widely regarded as effective in modeling temporal dependencies and spatial feature organization. However, unlike natural video sequences where temporal variation is dominated by local texture dynamics and intensity changes, swarming colony expansion is driven primarily by the propagation of morphological fronts, the generation and transmission of local geometric perturbations, and the temporal evolution of the global expansion velocity field. Consequently, the ability to accurately track front curvature, scale variation, and perturbation continuity is far more critical than minimizing pixel–wise reconstruction error. To ensure a fair comparison, all models were retrained from scratch under identical data partitions, observation/prediction splits, and evaluation protocols, and their outputs were evaluated using the mIoU, HD95, and ASSD metrics, as shown in Fig. 2.3a.
The results indicate that existing video prediction models exhibit clear limitations in this forecasting setting. MIM and SimVP+gSTA achieve relatively high region–level overlap, with mIoU of 89.32% and 90.52%, respectively, suggesting that they are able to preserve the overall expansion pattern to some extent. However, both models perform suboptimally in terms of HD95 and ASSD, indicating persistent discrepancies between the predicted and true colony fronts. In particular, PredRNN and PredRNNv2, although designed to enhance temporal memory and long–range dependency modeling, show consistently higher HD95 and ASSD values than the other models. This suggests that their hidden–state–accumulation–based recurrent formulation is susceptible to error amplification during extrapolation, making it difficult to maintain accurate front localization and stable representation of local geometric details over extended prediction horizons. In addition, SimVP+TAU demonstrates moderate performance across all three metrics, failing to strike an effective balance between region–level consistency and boundary precision.
In contrast, Morpher achieves the best performance across all evaluation metrics, with a mIoU of 95.42%, an HD95 of 10.61 pixels and an ASSD of 3.93 pixels, representing a 5.4% gain in region–level overlap and reductions of 42.0% and 55.7% in HD95 and ASSD relative to the strongest baseline, SimVP+gSTA. Although only the final 20% of each sequence requires extrapolative prediction, this window lies in the late stage of colony expansion, where the advancing interface is large and highly sensitive; even small deviations grow into disproportionately large discrepancies as the front progresses. The marked decreases in HD95 and ASSD show that Morpher not only aligns with the true expansion pattern but also preserves the geometry of the propagating front as it deforms. Each percentage point of mIoU thus corresponds to a tangible gain in propagation stability and in the maintenance of morphological integrity. By retaining the trajectory of the advancing edge with greater fidelity, Morpher produces forecasts that are more consistent with the underlying biological process of swarming expansion and capture features that matter for nutrient access, epithelial protection and microbial competition.
Figure 4: Failure modes of generic video prediction in swarming morphology forecasting and the role of morphology-aware representation. a, Performance comparison under an 80% observation / 20% prediction protocol. State-of-the-art video prediction models are evaluated under identical input–output settings for long-horizon forecasting of swarming colony expansion. While several methods attain reasonable region-level overlap, boundary-related errors remain pronounced across models. Morpher achieves higher mIoU together with consistently lower HD95 and ASSD, reflecting improved accuracy in front propagation and boundary geometry. b, Long-horizon qualitative forecasts on two representative sequences spanning distinct growth regimes. In the dense–finger phase (top), generic predictors either progressively smooth high-curvature fingers or fail to follow the true front dynamics, leading to increasing deviation from the ground truth. In the late-stage near–concentric ring expansion (bottom), maintaining global radial scale and subtle curvature becomes challenging for these models. Across both regimes, Morpher maintains a coherent and temporally consistent front geometry over extended prediction horizons, remaining closer to the true front. c, Effect of segmentation backbone on forecasting performance. Using TexPol–Net masks leads to more stable and accurate final-frame forecasts, whereas YOLOv12 [41], the strongest alternative backbone in this task, produces competitive results with residual boundary deviation and curvature attenuation. In the final predicted frame, TexPol–Net achieves higher IoU than YOLOv12 across both growth regimes, with improvements from 94.21% to 96.63% in the dense–finger phase and from 93.30% to 96.38% in the near–concentric ring expansion.
To further corroborate the quantitative results, we compare the predicted colony evolution across representative growth stages, as illustrated in Fig. 2.3b. During the dense–finger–instability–dominated expansion phase, MIM displays a clear tendency toward morphological smoothing over time: fine–scale protrusions at the advancing front gradually collapse into low–curvature boundaries, suppressing the inherent radial anisotropy and attenuating the transmission of local perturbations, which in turn compromises the accuracy of global expansion characterization over extended prediction horizons. The SimVP variants employing TAU or gSTA temporal units preserve the overall outline of the colony to a certain extent, but fail to maintain the dense–finger morphology, frequently exhibiting truncated lobe tips, blurred curvature transitions, and unstable front propagation rates; their predicted boundaries also become progressively rounded, reflecting insufficient geometric resolution and resulting in noticeable boundary drift. MAU and the recurrent architectures PredRNN and PredRNNv2 generate shapes that are superficially close to the ground truth in terms of global geometry, yet their temporal evolution is consistently slower, characterized by delayed dense–finger development, systematic underestimation of front propagation speed, and reduced amplitude of perturbation transmission, producing a temporal mismatch in both radial scale and finger–scale development. In contrast, Morpher yields stable and coherent predictions across growth regimes, maintaining both the spatial continuity of local perturbations and the global advancement and curvature modulation of the expanding front, while preserving a stable advancing interface without loss of curvature transmission or propagation continuity.
In the second sequence, the colony enters a late–stage expansion regime with a predominantly disc–like symmetry and a stabilized advancing front. In this context, the forecasting challenge shifts from preserving dense–finger morphology to maintaining the global radial scale and subtle boundary curvature modulation. MAU, MIM, PredRNN, and PredRNNv2 consistently underestimate the overall expansion radius, producing noticeably contracted predictions. This suggests that their temporal representations favor local smoothing and are unable to capture the long–range coordinated forces driving the sustained outward advancement of the expansion front and its slow, anisotropic curvature modulation. The SimVP model employing the gSTA temporal module better approximates the global scale, yet its predicted boundaries remain irregular and exhibit gradual drift over time, indicating limited geometric stability once high–frequency edge cues have diminished. The SimVP model using the TAU unit performs intermediately, but displays inconsistent boundary curvature and localized radial deviations across extended prediction intervals, reflecting difficulties in balancing temporal smoothness with geometric precision. In contrast, Morpher maintains both the disc–like radial organization and the fine–scale curvature variations throughout this stabilized growth phase. Its predicted boundaries closely follow the true expansion front without contraction, drift, or distortion, demonstrating that Morpher preserves stable global organization even when local structural cues weaken. Together with the observations from the dense–finger–dominated regime, these results indicate that Morpher captures the underlying evolutionary dynamics of swarming colonies rather than relying on short–term motion extrapolation or local texture continuity.
The limited performance of conventional video prediction models in this task stems from their emphasis on global pixel–level variation rather than on the dynamical organization of the expanding morphological front. Morpher departs from this paradigm by operating directly on the binary mask sequences produced by TexPol–Net, enabling temporal modeling to take place within a compact latent representation that encodes structural configuration rather than raw appearance. Within this latent space, Morphon works together with the autoregressive feedback loop to continuously update the representation based on the accumulated history of colony growth. This mechanism allows Morpher to directly capture the speed, directional bias, and spatial coordination of front propagation. Coupled with a multi–scale decoder that preserves continuity between successive morphological states, the model maintains coherent global organization and smooth boundary evolution even over long forecast horizons. As a result, Morpher produces long–range predictions that exhibit higher dynamical consistency and more faithful morphological progression than existing video prediction approaches.
2.4 Small improvements in front recovery yield large gains in forecasting
While the predictive performance of Morpher demonstrates its ability to capture temporal dependencies, the reliability of long–term forecasts hinges on the accuracy of the input representation—the segmentation masks delineating colony morphology. To assess this relationship, we compared Morpher coupled with our TexPol–Net segmentation backbone to the same forecasting framework driven by YOLOv12. YOLOv12 achieves the second–highest segmentation accuracy on our swarming colony dataset—second only to TexPol–Net—and thus serves as the strongest alternative baseline for this task. This comparison provides a clear and controlled evaluation of how the segmentation backbone affects subsequent temporal modeling. Both systems follow the identical two–stage pipeline illustrated in Fig. 1, using the same Morpher architecture, training settings, and evaluation criteria; the only difference lies in the source of the input segmentation masks.
Although YOLOv12 (91.81 mAP50:95) trails TexPol–Net (92.48 mAP50:95) by less than one percentage point in static segmentation accuracy, this seemingly minor gap translates into a disproportionately large effect on temporal forecasting. As illustrated in Fig. 2.3c, the slight boundary inaccuracies present in YOLOv12 masks accumulate through Morpher’s autoregressive updates, leading to a gradual drift of the colony front and a measurable loss of curvature fidelity over time—up to a 2.4–3.1 IoU-point gap after 20% forecasting. Forecasts initialized from YOLOv12 masks exhibit smoothed protrusions and reduced dense–finger morphology, consistent with progressive attenuation of high–frequency boundary features. In contrast, TexPol–Net masks preserve sharper front geometry and coherent internal texture cues, allowing Morpher to maintain stable propagation of the expanding rim with minimal geometric distortion.
Temporal coherence in morphological forecasting depends not only on the expressiveness of the temporal model but also on the structural fidelity of its spatial inputs. In swarming dynamics—where front curvature, anisotropy, and the organization of dense fingers jointly define the expansion process—even slight degradations in segmentation accuracy can disrupt the causal progression linking successive morphological states. TexPol–Net mitigates this sensitivity through its TEA and PCA mechanisms, which together sharpen fine–scale front delineation while preserving the global radial organization of the colony. These inductive priors yield geometrically consistent masks that capture the underlying biophysical architecture with high fidelity, enabling Morpher to propagate colony morphology smoothly and without structural drift. From a modeling standpoint, the results reveal that even a seemingly marginal one–percentage–point improvement in segmentation accuracy can induce a qualitative shift in system behavior—from cumulative error amplification to dynamical stability. The integration of TexPol–Net and Morpher thus forms a synergistic framework that unites spatial precision with temporal reasoning, enabling quantitative reconstruction of curvature–driven, long–range evolution in swarming fronts and, more broadly, extending predictive modeling to other forms of collective biological motion.
2.5 Temporal models and inference strategies for front forecasting
We used Morpher to systematically probe how temporal sequence models, inference strategies, and the Morphon module influence long–horizon forecasts of swarming colony morphology, as summarized in Fig. 2.5b. Four architectures—a simple RNN, GRU, LSTM, and Transformer encoder—were evaluated under parallel versus autoregressive decoding, with or without Morphon.
Forecasting quality was assessed using a set of complementary metrics capturing geometric accuracy and dynamical stability, including region–level overlap, boundary deviation, front–propagation dynamics, temporal consistency, and angular organization. The quantitative results underlying these comparisons are reported in detail in Table 2, which provides the full ablation statistics for each architectural and inference configuration.
A first observation is that autoregressive decoding consistently produced more faithful colony evolution than parallel one–shot decoding. Across architectures, switching from parallel to autoregressive inference improved either mIoU, boundary metrics, or both. For example, the Transformer with Morphon achieved 94.80% mIoU and an ASSD of 4.63 px in parallel mode, compared with 95.42% mIoU and 3.93 px ASSD under autoregressive decoding. Similar patterns were observed for GRU and LSTM, where autoregression reduced HD95 and ASSD and kept RMSE at low values. This behavior matches the biology of the system: swarming expansion proceeds through many small, incremental deformations of the front, and each future shape is tightly constrained by the immediately preceding configuration. Parallel decoding enforces all future frames in a single pass, which tends to collapse the diversity of plausible growth trajectories into a smoother, more symmetric contour and suppresses local protrusions. Conditioning each prediction on the previous frame allows the model to adjust the advancing front step by step, preserving fingering, local curvature, and realistic front propagation over long horizons.
Figure 5: Morpher enables geometry-consistent long-horizon forecasting of swarming colony morphology. a, Architecture of Morpher. Observed segmentation masks are mapped to a compact morphological latent sequence by a multi-scale spatial encoder. Future evolution is predicted autoregressively in latent space, with each decoded morphology fed back to condition subsequent steps. The Morphon module selectively retrieves informative past states via cross-attention and integrates them through a learnable gate, while a multi-scale decoder reconstructs spatial detail and preserves peripheral protrusions and front curvature. b, Systematic evaluation of temporal architectures and inference strategies under a fixed 80% observation / 20% prediction protocol. Across eight complementary metrics capturing geometric accuracy and dynamical stability, autoregressive inference tends to outperform parallel decoding, and the inclusion of Morphon improves boundary fidelity and temporal consistency across sequence models (RNN, LSTM, GRU and Transformer). c, Robustness of morphological forecasting to observation length. For autoregressive models with Morphon enabled, prediction accuracy and stability improve smoothly as the observation ratio increases from 50% to 90%, indicating that Morpher exploits additional temporal context without overfitting to a specific observation–prediction split.
The sequence models themselves formed a clear performance hierarchy that aligns with their representational capacity. The plain RNN was consistently weakest on geometry-sensitive metrics. Without Morphon in parallel mode, it reached 93.23% mIoU but suffered from large boundary errors (ASSD 6.02 px, $|\Delta\mathrm{NAS}|$ 19.54%), yielding overly round and contracted colonies. Even with autoregression and Morphon, ASSD remained above 4.3 px and anisotropy above 16%, indicating that local perturbations at the front were rapidly forgotten and the contour relaxed toward a smooth disk. Introducing gating improved this behavior. The GRU in parallel mode without Morphon already reduced ASSD to 5.17 px and $|\Delta\mathrm{NAS}|$ to 15.31%, and in its best configuration (autoregressive with Morphon) it further lowered ASSD to 4.06 px and achieved the smallest RMSE (2.06 px per frame) among all models, indicating particularly sharp tracking of front advancement and rapid local changes. The LSTM occupied an intermediate regime: it produced lower HD and HD95 than the RNN and stable RMSE values (2.26–2.84 px per frame), capturing global expansion reliably but reacting less strongly to fine, transient undulations of the front.
The Transformer–based Morpher provided the most balanced overall forecasts. In the autoregressive Morphon setting, it achieved the highest mIoU (95.42%), the lowest ASSD (3.93 px), and the smallest anisotropy deviation ($|\Delta\mathrm{NAS}|=13.13\%$), while keeping both the RMSE of front propagation and $|\Delta\mathrm{H}_{2}|$ at low values. These scores indicate that the model not only captured the colony’s radial expansion and average outline, but also maintained fingering, sectoring, and higher–order harmonic structure. Self–attention over the encoded history enables the Transformer to couple distant parts of the front and integrate long–range constraints, which is particularly important for reproducing radial symmetry–breaking and coordinated fingering across the colony.
Morphon provided an additional axis of improvement by encoding a persistent structural memory of past fronts. For every architecture and inference strategy, enabling Morphon increased mIoU and sharpened the boundary. For instance, in parallel mode the RNN improved from 93.23% to 94.22% mIoU and reduced ASSD from 6.02 px to 4.85 px when Morphon was switched on. Autoregressive GRU improved HD95 from 12.03 px without Morphon to 10.14 px with Morphon, and the Transformer under autoregression decreased ASSD from 5.26 px to 3.93 px when Morphon was enabled. Angular metrics showed the same tendency: in the strong Transformer configuration, $|\Delta\mathrm{NAS}|$ decreased from 16.83% without Morphon in parallel mode to 13.13% with Morphon in autoregressive mode, reflecting the combined effect of autoregression and Morphon in reinforcing angular fidelity. Across models, $|\Delta\mathrm{H}_{2}|$ remained below 2.0%, indicating that second–harmonic structure was generally well preserved, while Morphon primarily helped to maintain the finer anisotropic patterning of the front.
Accurate forecasting depends on both an appropriate inference paradigm and biologically informed memory. Autoregressive decoding is crucial for maintaining causal continuity of growth, while Morphon stabilizes curvature and anisotropy by carrying forward structural cues. Within this design, the Transformer–based Morpher offers the best compromise between boundary fidelity, propagation dynamics, and angular organization. The GRU emphasizes rapid front dynamics, the LSTM ensures smooth global expansion at the cost of some fine structure, and the plain RNN underestimates both fingering and symmetry breaking. These patterns support a mechanistic interpretation: models that can remember and attend to past colony structure are best suited to predict how microbial collectives remodel themselves in time.
2.6 Forecasting stability across observation ratios
We next examined how Morpher behaves as the amount of observed history is systematically varied. For each temporal backbone operating in its best–performing configuration (autoregressive decoding with Morphon), the observation fraction was increased from 50% to 90%. Model performance was then evaluated using the same eight metrics described above. The resulting trends are summarized in Fig. 2.5c, with the corresponding quantitative values reported in Table 3. This analysis tests whether Morpher is tuned to a narrow observation–prediction split or instead maintains stable forecasting behavior across a broad range of experimental protocols.
Geometry-based metrics showed a smooth and largely monotonic improvement as more of the past was revealed. For all four sequence models, mIoU increased from the high–80% range at 50% observation to above 96% at 90%, while HD95 and ASSD decreased correspondingly. The Transformer, for example, improved from 88.22% to 96.79% mIoU and reduced ASSD from 9.49 px to 2.75 px as the observation ratio increased from 50% to 90%. The relative ranking of sequence models remained essentially unchanged: at each observation level the Transformer offered the best compromise between overlap and boundary error, GRU and LSTM occupied an intermediate regime, and the plain RNN lagged slightly behind. This preservation of ordering indicates that the choice of temporal architecture is more important than the precise observation ratio for recovering the colony outline and front position.
Front–propagation dynamics, measured by RMSE of radial expansion rate, exhibited a similar saturation pattern. The largest gains occurred between 50% and 60% observation, after which RMSE values fluctuated within a narrow band of roughly 2.0–2.3 px per frame across models. In other words, once the model has seen slightly more than half of the trajectory, the incremental benefit of additional frames for predicting front–propagation speed is modest. TCI behaved in a comparably stable way but with slightly larger relative variation. TCI values clustered around 63–65% from 50% to 80% observation, with the LSTM reaching the highest TCI (65.32%) at 80%. At 90% observation the prediction window becomes too short for a meaningful fluctuation index, which is why TCI is not reported there. From a mathematical standpoint, both RMSE and TCI are based on temporal differences rather than static shapes; they are sensitive to small phase shifts in the timing of front accelerations and pauses, even when the underlying boundary is well aligned.
Angular metrics, particularly $|\Delta\mathrm{NAS}|$, varied more across observation ratios and did not follow a simple monotonic trend. For the Transformer, $|\Delta\mathrm{NAS}|$ decreased from 14.62% at 50% observation to 11.93% at 60%, then increased to 15.57% at 70% before dropping again to 13.13% at 80% and settling around 14% at 90%. The recurrent models showed the same qualitative behaviour, with a pronounced peak near 70–80%. This seemingly counterintuitive pattern—more observed history not always yielding smaller anisotropy error—arises from how NAS is defined: it aggregates angular contrasts over the entire prediction horizon. Increasing the observation ratio shortens this horizon and shifts it toward later times, where the colony has already developed complex sectoring and branch splitting. In this regime, even small local disagreements in when a branch emerges or merges can produce relatively large changes in the angular power spectrum. Mathematically, $|\Delta\mathrm{NAS}|$ behaves more like a measure of angular fluctuation intensity than a simple bias; it is highly sensitive to local phase errors of individual fingers, and therefore displays a non–monotonic dependence on the observation window.
The second–harmonic descriptor $|\Delta\mathrm{H}_{2}|$ showed a different but equally non–trivial trend. For the three recurrent architectures, $|\Delta\mathrm{H}_{2}|$ increased almost monotonically from $\sim\!1.1$–1.2% at 50% observation to around 2.2–2.3% at 90%. The Transformer followed the same overall tendency, with values around 1.18–1.35% at 50–60% observation and $\sim\!2.2\%$ at 90%, albeit with a local maximum at 70% and a mild dip at 80%. At first glance, this appears to contradict the intuition that more context should always improve harmonic structure. However, $|\Delta\mathrm{H}_{2}|$ isolates only the second angular mode of the front. As the colony matures, curvature energy is progressively redistributed into higher–order modes through branch splitting and tip instabilities. The prediction window at high observation ratios is therefore dominated by subtle rearrangements of fine–scale protrusions rather than by the initial quadrupolar shape. In this late regime, tiny mismatches in the amplitude or phase of low–frequency modes can produce larger absolute differences in the second harmonic, even when the full boundary and overall anisotropy are well captured. From a mathematical perspective, the metric is probing how accurately the model tracks a single Fourier component in a spectrum whose energy is shifting across modes; as the spectrum becomes more crowded, that task becomes harder rather than easier.
Morpher remains stable across observation ratios for coarse geometric descriptors and front–propagation speed, whereas the more sensitive temporal and angular metrics reveal where the remaining difficulty lies. Once roughly 60–70% of the trajectory has been observed, mIoU, HD–based distances, ASSD, and RMSE change only gradually with further increases in observation, and the relative ranking of sequence models remains fixed. In contrast, $|\Delta\mathrm{NAS}|$ and $|\Delta\mathrm{H}_{2}|$ respond strongly to how the forecasting window slices the late–time dynamics of sectoring and finger rearrangement, whereas TCI shows modest but noticeable sensitivity. This pattern indicates that changing the observation–prediction split does not alter the relative ranking of the sequence models in recovering global expansion, and that the main residual challenges are concentrated in the most fluctuation– and angle–sensitive aspects of the swarming front.
3 Discussion
Swarming colonies are often cited as emblematic examples of emergent collective behaviour, yet most analyses remain anchored in static descriptions of highly structured fronts. Here we recast the same imaging data into a predictive formulation of swarming dynamics. Rather than treating expansion as a sequence of images, the colony edge is modelled as a dynamical object evolving within a low–dimensional morphological state space. By assembling the SwarmEvo dataset, introducing TexPol–Net for boundary measurement, and developing Morpher as a morphology–level dynamical model with structural memory, the expansion of Enterobacter sp. SM3 becomes a forecasting problem defined at the level of geometry, rather than a descriptive account of shape.
This reframing rests on an explicit separation between measurement, state construction and dynamics. Time–lapse videos are first reduced to sequences of boundary–resolved masks, defining the measured observables of the system. TexPol–Net operates at this stage not simply as a segmentation network, but as a boundary measurement operator that stabilizes the morphological degrees of freedom of the colony front. By suppressing diffuse edges, preserving thin protrusions and decoupling the front from background artefacts, TexPol–Net defines a consistent coordinate system in which small geometric variations retain temporal coherence. Improvements in boundary accuracy therefore propagate nonlinearly into forecasting performance, as every measurement error reshapes the subsequent trajectory in state space. In this sense, TexPol–Net establishes the state variables on which dynamics can be meaningfully learned, rather than merely refining per–frame masks.
This sensitivity becomes most apparent when segmentation backbones are required to resolve near–equivalent masks. On SwarmEvo, YOLOv12 differs from TexPol–Net by less than one mAP50:95 point (91.81 vs. 92.48; Fig. 3c), yet this minor discrepancy at the measurement stage expands once the masks define the state variables for autoregressive forecasting. When Morpher is held fixed, boundary deviations in YOLOv12 inputs, including slight edge displacement and reduced high–curvature detail, propagate through successive rollouts, producing gradual front drift and a measurable loss of curvature that reaches a 2.4–3.1 IoU gap after 20% prediction (Fig. 2.3c). TexPol–Net masks, by contrast, sustain consistent front localization and curvature over extended horizons, limiting geometric degradation during temporal propagation (Fig. 3b). SAM/SAM2 accentuate this contrast by suppressing distal structure and biasing contours toward the colony core, effects that are already evident at the segmentation stage (Fig. 3). These comparisons indicate that once the colony edge is treated as the dynamical state, boundary fidelity directly governs whether temporal prediction remains stable or undergoes progressive error growth.
Morpher operates on this measurement–defined state space to model the evolution of the front. Unlike generic video prediction architectures, Morpher is constrained to propagate contours and to encode dynamics in terms of geometric quantities such as curvature and radial displacement (Fig. 2.5). This design choice places Morpher explicitly in the role of a dynamical model, rather than an image generator. Benchmarking against state–of–the–art sequence models shows that enforcing this geometric alignment improves both region–level overlap and boundary fidelity, while preserving anisotropic branching despite the absence of texture information. Ablations across backbones, inference strategies and the Morphon module reinforce a mechanistic interpretation of the model behaviour (Fig. 2.5b). Autoregressive decoding consistently yields more realistic front propagation than parallel one–shot prediction, reflecting the incremental nature of physical growth. Architectures equipped with attention and persistent structural memory, particularly the Transformer with Morphon, best preserve fingering and radial symmetry breaking, indicating that long–range couplings are integral to swarming dynamics.
Varying the observation ratio provides further insight into the stability of morphology–level forecasting (Fig. 2.5c). As the observed fraction increases from 50% to 90%, coarse geometric metrics such as mIoU, HD–based distances and ASSD improve and then saturate. Once slightly more than half of the trajectory is observed, additional frames contribute little to recovering the global outline or average advance speed. In contrast, temporal and angular descriptors reveal where prediction remains sensitive. TCI exhibits measurable dependence on the forecasting window, indicating that small phase shifts in accelerations and pauses perturb fluctuation profiles even when boundaries remain well aligned.
Angular metrics expose a more stringent regime. The non–monotonic dependence of $|\Delta\mathrm{NAS}|$ on observation ratio suggests that anisotropy errors are governed less by the amount of historical data than by how the prediction window intersects late–stage sectoring and branch rearrangements. In this regime, minor mismatches in timing or amplitude reshape the angular power spectrum even when global geometry is preserved. The behaviour of $|\Delta\mathrm{H}_{2}|$ sharpens this interpretation. Isolating the second harmonic renders the metric sensitive to the redistribution of curvature energy across Fourier modes. At high observation ratios, prediction targets are dominated by fine protrusions and tip instabilities, where small deviations in low–frequency content amplify into substantial harmonic differences. These results indicate that anisotropic fingering and harmonic organization represent the most demanding aspects of swarming to forecast, and they suggest a fundamental sensitivity of morphology–only dynamics to small temporal errors.
Several limitations delineate avenues for extension. SwarmEvo focuses on a single strain, Enterobacter sp. SM3. Although its behaviour spans multiple regimes, swarming encompasses a broad spectrum of species and substrates whose statistics may differ substantially. The present representation further collapses colonies into quasi–two–dimensional fronts. While such projections preserve the dominant geometric degrees of freedom and are standard in swarming studies, height and density variations enter only implicitly. Morpher is trained without explicit biophysical constraints; the absence of hand–crafted rules affords flexibility, yet extrapolation into sparsely sampled regimes may benefit from mechanistic priors. Finally, forecasting is treated as an open–loop problem. Closed–loop experiments in which predictions guide perturbations would provide a natural next step toward control.
Beyond swarming, this work establishes a general methodology for forecasting dynamics in systems where appearance is uninformative but geometry is structured. By formalizing expansion as a measurement–defined dynamical system—where boundary measurement defines the state space and forecasting reduces to trajectory evolution rather than image prediction—the framework clarifies how segmentation, state construction and temporal inference interact. Representing videos as trajectories in a geometry-aligned morphological state space enables principled comparison of sequence models, separates measurement error from dynamical uncertainty, and supports interpretable metrics linked directly to branching, anisotropy and harmonic organization. These principles extend to other forms of collective motility, biofilm spreading and tissue invasion, where dynamics unfold through shape rather than texture. By enabling long–horizon forecasts and identifying features that are intrinsically sensitive to timing, this framework transforms colony expansion into a quantitative, methodologically grounded basis for intervention, providing a foundation for predictive modelling and control of microbial collectives in complex environments.
4 Methods
4.1 Experimental workflow, imaging, and mask generation
The experimental workflow followed Fig. 1. A single colony of Enterobacter sp. SM3 was transferred from an LB agar plate into LB broth (10 g/L tryptone, 5 g/L yeast extract, and 5 g/L NaCl) and incubated at $37\,^{\circ}\mathrm{C}$ with shaking at 200 rpm for approximately 16 h to obtain a high-density culture. A $5\text{--}8\,\mu\mathrm{L}$ aliquot of the overnight culture was spotted onto the center of a freshly prepared swarming agar plate (LB medium supplemented with 0.5–0.8% agar), poured to a uniform thickness of 3–4 mm to promote consistent hydration and nutrient diffusion across plates. The inoculated plates were incubated at $30\,^{\circ}\mathrm{C}$ and approximately 90% relative humidity for 4–6 h to activate swarming motility, and then transferred to a time-lapse imaging chamber maintained between $28\text{--}37\,^{\circ}\mathrm{C}$ and 80–92% relative humidity. Humidity was kept below the condensation threshold to prevent droplet formation on the agar surface. Surface stiffness and inoculum volume were varied across experiments to introduce controlled physical differences and generate diverse colony expansion conditions.
Imaging was performed using a vertically mounted high-resolution digital camera with uniform LED illumination to minimize specular reflection and shadowing artifacts. Frames were recorded at one-minute intervals throughout colony expansion and continued until the colony front reached the plate boundary or no further measurable growth was observed. Sequences were stored at native spatial resolution with associated timestamps.
For downstream analysis, each sequence was segmented using TexPol–Net to obtain pixel-level colony masks. These mask sequences formed a morphology-consistent representation used for training and evaluating the forecasting models. Training and validation datasets were assembled from non-overlapping experimental runs collected under distinct growth conditions to assess generalization across colony expansion regimes.
4.2 Recovering front texture and geometry
Swarming colony images present two coupled challenges for convolutional backbones. At the micro-structural level, boundaries are uncertain and textures vary across scales, so high-frequency evidence is easily attenuated when features are aggregated over local receptive fields. At the meso- to macro-structural level, growth follows a disc-like geometry with coherent radial propagation, producing long-range dependencies that are poorly captured by standard kernels. Addressing both phenomena with a single mechanism tends to inflate complexity or compromise either edge fidelity or global organization.
The proposed design decouples the challenges of texture variation and geometric organization while maintaining a unified attention framework. The Texture-Edge Attention (TEA) module, illustrated in Figure 6a, is devised to enhance the stability and clarity of fine-scale features that are often blurred by conventional filters. It integrates three cooperative components: a local depthwise branch that preserves intra-channel spatial detail, a multi-dilated branch that captures textures over multiple scales, and an edge-sensitive branch initialized with a discrete Laplacian kernel to introduce a high-pass prior at the beginning of training. Channel-wise and spatial attention mechanisms further purify the fused representation by reducing redundancy and highlighting informative regions, while a conditional residual pathway ensures optimization stability when feature dimensions match. Detailed formulations and implementation details for both modules are presented in Section S2.
The Polar–Context Attention (PCA) module, illustrated in Figure 6b, complements TEA by embedding a radial geometric prior directly into the feature representation. The module first compresses feature channels to eliminate redundancy and then performs a polar transformation that maps features into $(\rho,\theta)$ space, where depthwise dilated convolutions operate along radial and angular directions consistent with the colony’s outward propagation. A large-kernel Cartesian branch provides broad contextual coverage, while a local branch preserves fine structural fidelity. Channel- and spatial-attention mechanisms further refine the fused representation, enhancing the coherence between local detail and global geometry. A conditional residual connection stabilizes training when feature dimensions coincide.
TEA and PCA provide complementary inductive biases: TEA sharpens edges and preserves high-frequency texture amid blurred, irregular boundaries, whereas PCA captures geometry-aligned, long-range dependencies characteristic of near–concentric–ring radial growth. Operating both modules at a compact internal width with shared normalization and activation yields a representation that is simultaneously edge-aware, scale-robust, and geometry-aligned, without incurring prohibitive computational cost.
4.3 TexPol–Net
Texture- and geometry-aware instance segmentation. To recover accurate colony fronts from swarming imagery, we developed the Texture-Polar Network (TexPol–Net), an instance segmentation architecture that integrates fine-scale texture encoding with geometry-aligned contextual reasoning. This design addresses two coupled challenges in swarming colonies: diffuse and low-contrast boundaries at the local scale, and near–concentric radial organization at the global scale. TexPol–Net draws inspiration from efficient one-stage detectors such as YOLO [41, 35] and the prototype-based mask formulation of YOLACT [3], while introducing texture- and geometry-aware mechanisms tailored to biological front segmentation (Fig. 3a).
Hierarchical backbone with complementary attentions. TexPol–Net employs a five-stage hierarchical convolutional backbone with progressive downsampling. At each stage, Texture-Edge Attention (TEA) and Polar–Context Attention (PCA) are jointly embedded to serve complementary roles. TEA operates predominantly on shallower representations to preserve boundary sharpness and local texture variation, whereas PCA increasingly emphasizes geometry-aligned contextual aggregation in deeper layers, enabling the capture of long-range radial propagation patterns.
Bidirectional multi-scale feature fusion. The neck adopts a PANet-style bidirectional fusion strategy [25] to integrate semantic and spatial information across scales. Top–down and bottom–up pathways are interleaved with PCA modules, ensuring that polar consistency is preserved during resolution recovery. A final TEA module is applied prior to prediction to recalibrate channel responses and reinforce fine-scale texture cues.
Dense prediction and prototype-based mask assembly. The prediction head employs fully convolutional branches on feature levels $P3$, $P4$, and $P5$ to predict class scores, bounding boxes, and instance-specific mask coefficients. Instance masks are generated using a YOLACT-style prototype formulation: a lightweight Protonet produces k=32k{=}32 shared prototypes, and each detection predicts a coefficient vector for linear combination. After non-maximum suppression, masks are generated, cropped, and thresholded within the retained bounding regions.
Loss. During training, supervision is applied to all prediction heads in each mini-batch. The overall objective is defined as a weighted linear combination of the bounding box regression, classification, distribution focal loss (DFL), and mask segmentation terms:
$\mathcal{L}_{\mathrm{seg}}=\lambda_{\mathrm{box}}\,\mathcal{L}_{\mathrm{box}}+\lambda_{\mathrm{cls}}\,\mathcal{L}_{\mathrm{cls}}+\lambda_{\mathrm{dfl}}\,\mathcal{L}_{\mathrm{dfl}}+\lambda_{\mathrm{mask}}\,\mathcal{L}_{\mathrm{mask}}.$
The box loss penalizes geometric misalignment between predicted and ground-truth boxes, the classification loss supervises category confidence, and DFL refines localization by regularizing the underlying discrete distribution. The segmentation loss enforces consistency between predicted and ground-truth instance masks. All loss terms are combined linearly with fixed weights, which are kept constant across experiments.
4.4 Morpher
Mask-space forecasting. To model the temporal evolution of swarming colony morphology, we developed Morpher, a forecasting network that operates on segmentation masks rather than raw image intensities. This choice follows the biology of swarming expansion, where growth is expressed through the geometry and advancement of the colony front. Mask-level representation isolates occupied regions and evolving boundaries, focusing the model on front curvature and large-scale expansion while reducing sensitivity to illumination and imaging variability (Fig. 2.5).
Spatial encoder and latent sequence. Given an input mask sequence, a multi-scale encoder extracts a hierarchy of feature maps $\{f_{1}^{(t)},f_{2}^{(t)},f_{3}^{(t)},f_{4}^{(t)},f_{5}^{(t)}\}$ at each time step via progressive downsampling $(H\!\to\!H/2\!\to\!H/4\!\to\!H/8\!\to\!H/16\!\to\!H/32)$. In parallel, each frame is compressed into a latent descriptor $z_{t}\in\mathbb{R}^{256}$, yielding a compact temporal representation for forecasting. For temporal modeling, the observed segment is represented as a latent tensor of shape $(B,T_{\mathrm{obs}},256)$, where the latent dimensionality is fixed by the shared spatial encoder. Sinusoidal time encodings are added to the latent sequence prior to the temporal module to provide a uniform representation of temporal order.
Autoregressive temporal evolution. Morpher forecasts in an autoregressive loop. After observing $T_{\mathrm{obs}}$ frames, the model predicts $\hat{z}_{T_{\mathrm{obs}}+1}$, decodes it into a mask, and re-encodes the prediction as the next input. This feedback couples future states to the geometry produced at the previous step and helps maintain boundary continuity over long horizons. Forecasting is evaluated under two inference regimes in matched settings. In the autoregressive regime, predictions are generated stepwise over the horizon, with each predicted mask fed back as the next input. In the parallel regime, the model predicts the full forecasting horizon conditioned on the observed segment and the terminal latent state.
Sequence model variants with matched capacity. The temporal module is instantiated using one of four architectures: a vanilla RNN [13], GRU [10], LSTM [16], or a Transformer encoder [42]. All variants share the same spatial encoder–decoder backbone and operate in the same latent space. Across architectures, temporal information is summarized into a single terminal latent state that serves as the sole conditioning variable for future prediction, enforcing a consistent past-to-state formulation. The recurrent variants use three recurrent layers, while the Transformer uses three stacked encoder blocks. No model introduces future tokens, decoder-side attention, or changes to the latent dimensionality.
Morphon memory integration. To regulate how past morphology influences future predictions, Morphon retrieves informative historical states from the latent history via cross-attention, using a learnable query derived from the current state. The retrieved representation is fused into the evolving temporal state through a learnable gate $\alpha\in(0,1)$, balancing historical memory against the current predictive trajectory.
Multi-scale decoding and feedback. A multi-scale decoder mirrors the encoder hierarchy to progressively restore spatial resolution while reinjecting intermediate features retained during encoding. This design preserves peripheral protrusions and fine-scale curvature patterns during reconstruction. During training, the predicted mask is re-encoded and fed back for subsequent steps, matching the inference-time autoregressive procedure.
Sequence structuring. Observation length $T_{\mathrm{obs}}$ and forecasting horizon are treated as tunable hyperparameters. Sequences are constructed by uniformly subsampling frames with a fixed stride and partitioned into observation and prediction segments under predefined observation–prediction ratios.
Loss. Supervision is applied at every predicted time step within the forecasting horizon. Let $T$ denote the number of predicted frames. The training objective is defined as a temporal average over the prediction window:
$\mathcal{L}_{\mathrm{pred}}=\frac{1}{T}\sum_{t=1}^{T}\Big(\lambda_{\mathrm{focal}}\,\mathcal{L}_{\mathrm{focal}}^{(t)}+\lambda_{\mathrm{softIoU}}\,\mathcal{L}_{\mathrm{softIoU}}^{(t)}+\lambda_{\mathrm{boundary}}\,\mathcal{L}_{\mathrm{boundary}}^{(t)}\Big).$
Focal loss addresses foreground–background imbalance by up-weighting hard pixels, while soft IoU enforces region-level overlap. A boundary-aware loss further penalizes discrepancies along the colony front; it is computed using a GPU-efficient morphological gradient approximation and evaluated at one-quarter of the original spatial resolution. Loss terms are averaged uniformly over the prediction horizon, and the relative weights are fixed across experiments.
Interpretability as a dynamical model. By expressing each predicted morphology as a causal progression from prior front states, Morpher provides a data-driven dynamical representation of swarming expansion, enabling analysis of curvature perturbation propagation over time.
4.5 Evaluation metrics for colony front segmentation
We evaluated segmentation performance using four complementary metrics: mAP50:95, the Dice coefficient, the number of learnable parameters, and inference speed measured in frames per second (FPS). Among these, mAP50:95 serves as the primary metric, as it provides a comprehensive measure of segmentation quality across a range of localization tolerances. Following the COCO evaluation protocol, AP is first computed at multiple Intersection–over–Union (IoU) thresholds, and mAP50:95 is obtained by averaging across IoU thresholds ranging from 0.50 to 0.95, with a step size of 0.05:
$\mathrm{mAP}_{50:95}=\frac{1}{10}\sum_{\tau\in\{0.50,0.55,\dots,0.95\}}\left[\frac{1}{101}\sum_{i=0}^{100}P\!\left(\tfrac{i}{100};\,\tau\right)\right],$
where $P(r;\tau)$ denotes the interpolated precision at recall level $r$ for IoU threshold $\tau$.
In addition to mAP, segmentation accuracy is quantified using the Dice coefficient, which measures the overlap between predicted and ground-truth masks. For a predicted mask $\hat{Y}$ and a ground-truth mask $Y$, the Dice coefficient is defined as
$\mathrm{Dice}=\frac{2\,|\hat{Y}\cap Y|}{|\hat{Y}|+|Y|},$
with values ranging from 0 to 1, where higher scores indicate better spatial agreement.
To assess model complexity, we report the total number of learnable parameters, which indicates the representational capacity and computational burden of the segmentation network. To evaluate practical efficiency, we additionally measure inference throughput in FPS, defined as the number of images the model can process per second under the same hardware configuration.
4.6 Evaluation metrics for morphological forecasting
Forecasting accuracy must be judged not only by per–frame agreement, but also by how faithfully the predicted colony advances and organizes its growth direction over time. Accordingly, our evaluation considers both the spatial fidelity of the predicted masks and boundaries, and the temporal consistency of the advancing front, including its radial expansion speed and the coherence of its directional variation around the colony rim. All symbols and operations used in the metrics below are defined explicitly.
To quantify mask fidelity over time, let $\widehat{Y}_{t}$ and $Y_{t}$ denote the predicted and true masks at time $t$, and let $\partial\widehat{Y}_{t}$ and $\partial Y_{t}$ denote their corresponding boundaries. We report the mean Intersection over Union (mIoU), defined as the temporal average of the IoU used in the AP protocol:
$\mathrm{mIoU}=\frac{1}{T}\sum_{t=1}^{T}\frac{\lvert\widehat{Y}_{t}\cap Y_{t}\rvert}{\lvert\widehat{Y}_{t}\cup Y_{t}\rvert}$.
To evaluate boundary placement in a manner sensitive to fine-scale morphological deviations, we use the symmetric Hausdorff distance,
$d_{\mathrm{H}}=\frac{1}{T}\sum_{t=1}^{T}\max\Big\{\max_{y\in\partial Y_{t}}\min_{\hat{y}\in\partial\widehat{Y}_{t}}\|y-\hat{y}\|,\;\max_{\hat{y}\in\partial\widehat{Y}_{t}}\min_{y\in\partial Y_{t}}\|\hat{y}-y\|\Big\},$
as well as its 95th–percentile variant,
$d_{\mathrm{H}_{95}}=\frac{1}{T}\sum_{t=1}^{T}\max\Big\{P_{95}\!\big(\min_{\hat{y}\in\partial\widehat{Y}_{t}}\|y-\hat{y}\|\big)_{y\in\partial Y_{t}},\;P_{95}\!\big(\min_{y\in\partial Y_{t}}\|\hat{y}-y\|\big)_{\hat{y}\in\partial\widehat{Y}_{t}}\Big\},$
which downweights sparse, extreme outliers arising from transient high–curvature growth. Finally, the average symmetric surface distance provides a measure of mean bidirectional nearest–boundary discrepancy,
$d_{\mathrm{ASSD}}=\frac{1}{T}\sum_{t=1}^{T}\frac{1}{\lvert\partial\widehat{Y}_{t}\rvert+\lvert\partial Y_{t}\rvert}\left(\sum_{y\in\partial Y_{t}}\min_{\hat{y}\in\partial\widehat{Y}_{t}}\|y-\hat{y}\|+\sum_{\hat{y}\in\partial\widehat{Y}_{t}}\min_{y\in\partial Y_{t}}\|\hat{y}-y\|\right).$
To evaluate whether the predicted colony reproduces the correct front propagation dynamics, we measure the radial expansion speed along $K$ uniformly sampled angular directions $\{\theta_{k}\}_{k=1}^{K}$ from the colony centroid. In our experiments, we set $K=720$, which corresponds to a sampling interval of $0.5^{\circ}$, which provides sufficiently fine angular resolution to capture the subtle, high–curvature perturbations that arise during dense–finger instabilities. Let $r(\theta_{k},t)$ and $\widehat{r}(\theta_{k},t)$ denote the ground–truth and predicted radial distances at time $t$. The corresponding expansion velocities are
$v(\theta_{k},t)=\frac{r(\theta_{k},t)-r(\theta_{k},t-\Delta t)}{\Delta t},\qquad\widehat{v}(\theta_{k},t)=\frac{\widehat{r}(\theta_{k},t)-\widehat{r}(\theta_{k},t-\Delta t)}{\Delta t}.$
The overall accuracy of front advancement is quantified by the mean root–mean–square error,
$\mathrm{RMSE}=\frac{1}{T-1}\sum_{t=2}^{T}\sqrt{\frac{1}{K}\sum_{k=1}^{K}\big(\widehat{v}(\theta_{k},t)-v(\theta_{k},t)\big)^{2}}.$
Beyond matching the average speed, we also assess whether the temporal fluctuation strength of the radial velocity field is preserved. To this end, we compute the Temporal Consistency Index (TCI) over sliding windows $w=1,\dots,W$ of fixed length $L=4$ radius frames, corresponding to three consecutive velocity steps. For each window and direction, we compare the temporal standard deviation of the predicted velocity trace, denoted $\sigma^{(w)}_{\widehat{v},k}$, with that of the ground–truth velocity trace, denoted $\sigma^{(w)}_{v,k}$. The directional consistency score is
$\mathrm{TCI}^{(w)}_{k}=1-\frac{\big|\sigma^{(w)}_{\widehat{v},k}-\sigma^{(w)}_{v,k}\big|}{\sigma^{(w)}_{\widehat{v},k}+\sigma^{(w)}_{v,k}+\varepsilon},$
and is evaluated only when $\sigma^{(w)}_{\widehat{v},k}+\sigma^{(w)}_{v,k}>\tau_{0}$, where $\tau_{0}=10^{-6}$ is a numerical tolerance that filters out angular directions with effectively no front motion, ensuring that TCI reflects meaningful temporal fluctuations rather than numerical artefacts. The final index aggregates all valid scores across windows and directions:
$\mathrm{TCI}=\frac{1}{W}\sum_{w=1}^{W}\frac{1}{\lvert\mathcal{K}_{w}\rvert}\sum_{k\in\mathcal{K}_{w}}\mathrm{TCI}^{(w)}_{k}\in[0,1],$
where $\mathcal{K}_{w}$ denotes the set of valid directions in window $w$; higher values indicate better preservation of intrinsic fluctuation dynamics.
To evaluate the organization of growth across angles, we examine the anisotropy of the radial velocity field. The normalized angular spread (NAS) is computed as the angular standard deviation divided by the angular mean of the velocity at each time point, where $\mathrm{NAS}_{t}$ refers to the value computed from the ground–truth velocity field and $\widehat{\mathrm{NAS}}_{t}$ refers to the value computed from the predicted one. We report their mean absolute deviation over time:
$\lvert\Delta\mathrm{NAS}\rvert=\frac{1}{T-1}\sum_{t=2}^{T}\big|\widehat{\mathrm{NAS}}_{t}-\mathrm{NAS}_{t}\big|.$
To further characterize directional patterning, we examine the angular Fourier spectrum of $v(\theta,t)$ and extract the normalized second–harmonic power,
$\mathrm{H}_{2,t}=\frac{\big|\mathcal{F}_{\theta}\{v(\theta,t)\}[2]\big|^{2}}{\sum_{m}\big|\mathcal{F}_{\theta}\{v(\theta,t)\}[m]\big|^{2}},\qquad\widehat{\mathrm{H}}_{2,t}=\frac{\big|\mathcal{F}_{\theta}\{\widehat{v}(\theta,t)\}[2]\big|^{2}}{\sum_{m}\big|\mathcal{F}_{\theta}\{\widehat{v}(\theta,t)\}[m]\big|^{2}},$
This quantity reflects the strength of the characteristic two–lobed angular mode. We define the Second–Harmonic Power Deviation as
$\lvert\Delta\mathrm{H}_{2}\rvert=\frac{1}{T-1}\sum_{t=2}^{T}\big|\widehat{\mathrm{H}}_{2,t}-\mathrm{H}_{2,t}\big|.$
4.7 Implementation of TexPol–Net
Framework and experimental setup. All models are trained using the Ultralytics YOLO framework, with the task configured for image segmentation. All experiments are conducted under identical training configurations and independently repeated to ensure fair comparability.
Training and optimization protocol. For training, the maximum number of epochs is set to 300, and an early stopping strategy is employed to mitigate overfitting once validation performance saturates. The batch size is set to 16, and all input images are resized to a fixed resolution of $640\times 640$, balancing computational efficiency and memory usage. Automatic mixed-precision (AMP) training is enabled throughout to improve training throughput and reduce memory consumption while maintaining numerical stability. The optimizer is set to stochastic gradient descent (SGD). The initial learning rate is set to $6\times 10^{-3}$ and is combined with a linear warm-up schedule over the first three epochs. After warm-up, the learning rate is gradually decayed to $1\%$ of its initial value. A weight decay of $5\times 10^{-4}$ and a momentum of $0.937$ are used to stabilize gradient updates and improve generalization. All experiments use a fixed random seed and enable deterministic training to ensure reproducibility.
4.8 Implementation of Morpher
Sequence construction and data preprocessing. Morpher operates exclusively on binary colony masks generated by a pretrained segmentation model and does not directly access raw image intensities. In our primary pipeline, TexPol–Net was used to generate these masks, while alternative segmentation models were used in comparative experiments. Each training sample therefore consisted of a temporally ordered sequence of segmentation masks representing colony occupancy and front geometry. Sequences were constructed by uniformly subsampling frames from the full temporal series using a fixed stride, yielding equal temporal spacing between adjacent frames, and the resulting sequence length $T$ was defined by this fixed-stride subsampling rule. Each sequence was partitioned into an observation segment and a prediction segment using observation–prediction ratios of $0.5/0.5$, $0.6/0.4$, $0.7/0.3$, $0.8/0.2$, and $0.9/0.1$. Within each experiment, all masks were generated using the same segmentation model to ensure a consistent morphological representation across both segmentation and forecasting stages. All masks were resized to a spatial resolution of $640\times 640$, which was used consistently across all experiments. The dataset was split into training and validation partitions at the growth-sequence level, with no colony contributing sequences to both partitions, and all quantitative results are reported on the validation split.
Training and optimization protocol. All variants were optimized with AdamW using an initial learning rate of $5\times 10^{-5}$ and weight decay of $1\times 10^{-4}$. Training was conducted for 300 epochs with a batch size of 2. All input mask sequences were resized to a fixed spatial resolution of $640\times 640$. A linear warm-up was applied over the first 10% of optimization steps, followed by cosine annealing. The global gradient norm was clipped to 1.0. Mixed-precision training was enabled throughout via automatic mixed precision with gradient scaling to improve computational throughput while maintaining numerical stability. Validation was performed at every epoch, and model selection was based on the checkpoint achieving the highest validation mIoU. No early stopping was applied. All experiments were conducted with fixed random seeds and deterministic backend settings to ensure reproducibility. TensorFloat-32 acceleration was enabled for matrix multiplications on supported hardware, while cuDNN was configured in deterministic mode.
4.9 Running of existing methods
Adjustable training settings were kept aligned across methods whenever applicable. All images were resized to $640\times 640$. Models trained with epoch-based schedules were optimized for 300 epochs, while models trained with iteration-based schedules explicitly report the corresponding iteration counts. For all video prediction models, the batch size was fixed at 2.
YOLOv11 and YOLOv12. YOLOv11 and YOLOv12 were trained and evaluated using the Ultralytics YOLO framework with the task configured for image segmentation. Early Stopping was enabled once validation performance saturated. A batch size of $16$ was used. AMP was enabled throughout training. Optimization used SGD with an initial learning rate of $6\times 10^{-3}$ and a linear warm-up over the first three epochs; the learning rate was then decayed to $1\%$ of its initial value. Weight decay and momentum were set to $5\times 10^{-4}$ and $0.937$, respectively.
SAM and SAM2. SAM and SAM2 were fine-tuned for segmentation under a unified training protocol. For SAM, training was conducted using an adaptive optimizer with weight decay (initial learning rate $8\times 10^{-4}$; weight decay $1\times 10^{-4}$). A warm-up phase of $250$ optimization steps was applied at the beginning of training, followed by stepwise learning-rate decays at iteration steps 60,00060{,}000 and 86,66686{,}666, each with a decay factor of $1/10$. The model was built on the SAM ViT-B backbone, and a selective freezing strategy was adopted: the image encoder and prompt encoder were frozen, while only the mask decoder was updated during training. For SAM2, training was formulated as a binary segmentation task with RGB images as input and binary masks as supervision. Images were normalized to the range $[0,1]$ using ToTensor(), while masks were processed using nearest-neighbor interpolation to preserve label integrity. Masks were explicitly binarized to restrict pixel values to $\{0,1\}$. Optimization was performed using Adam with an initial learning rate of $1\times 10^{-4}$ and no additional learning-rate scheduling. A batch size of $4$ was used, and BCEWithLogitsLoss was applied directly to the raw logits.
MAU. MAU was run using the official implementation. The model employed four recurrent layers with hidden dimension 64, convolutional filters of size 5 with stride 1 and patch size 1, and no layer normalization. The spatiotemporal relation size was set to 2 and the temporal decay parameter to $\tau=5$. Scheduled sampling was enabled, with the sampling probability linearly decayed from 1.0 to 0 over 50,000 iterations at a rate of $2\times 10^{-5}$. Training used the Adam optimizer with a learning rate of $5\times 10^{-4}$ and a OneCycle learning rate scheduler.
MIM. MIM was run using the official implementation built on the PredRNN framework. The model employed four recurrent layers with hidden dimensions of 128, convolutional filters of size 5 with stride 1 and patch size 4, and no layer normalization. Scheduled sampling followed the same linear decay strategy as above, while reverse scheduled sampling was disabled. Training used the Adam optimizer with a learning rate of $1\times 10^{-4}$ and a OneCycle learning rate scheduler; incomplete batches were dropped during training.
PredRNN and PredRNNv2. PredRNN-based models were run using the official implementations. Both models employed four recurrent layers with hidden dimensions of 128, convolutional filters of size 5 with stride 1 and patch size 2, and no layer normalization. Scheduled sampling was enabled with linear decay from 1.0 to 0 over 50,000 iterations at a rate of $2\times 10^{-5}$. Training used the Adam optimizer with a learning rate of $1\times 10^{-3}$ and a OneCycle learning rate scheduler. PredRNNv2 additionally enabled reverse scheduled sampling with transition steps at 25,000 and 50,000 iterations and an exponential coefficient of 5,000, and incorporated a decoupling loss with weight $\beta=0.01$.
SimVP and SimVPv2. SimVP-based baselines were run using the official implementation. The spatial encoder–decoder employed a channel width of 64 with four convolutional blocks ($N_{S}=4$), while temporal modeling used a hidden dimension of 256 with eight temporal blocks ($N_{T}=8$). SimVP used TAU units for temporal prediction, whereas SimVPv2 replaced TAU with gSTA modules. Training used the Adam optimizer with a learning rate of $1\times 10^{-3}$ and a OneCycle learning rate scheduler. Model selection followed the validation loss criterion defined in the configuration.
4.10 Swarming Morphogenesis Evolution dataset
The Swarming Morphogenesis Evolution (SwarmEvo) dataset consists of high-resolution time-lapse recordings of Enterobacter sp. SM3 acquired at a fixed spatial resolution of $1250\times 1250$ pixels. After augmentation, the dataset comprises 1,971 annotated sequences used for training and evaluating segmentation models, as well as 276 long time series derived from continuous recordings sampled at 1-min intervals, which serve as the basis for temporal modeling and multi-scale temporal downsampling. Data were collected across multiple agar plates and independent imaging sessions, introducing natural variability in growth dynamics and colony morphology across distinct morphogenetic regimes. Segmentation masks used for model training and evaluation were obtained through a dedicated segmentation pipeline and subsequently curated to ensure consistent boundary delineation, while temporal sequences were generated by propagating these masks across time to support forecasting tasks. Details of cultivation conditions, imaging setup, and mask generation are provided in Section 4.1.
Segmentation-level augmentation For training the segmentation model, augmentations were applied independently to each image–annotation pair. Photometric perturbations included linear intensity rescaling with offset, gamma correction, additive Gaussian noise, and sparse impulse-like pixel corruption. Geometric transformations were sampled per image and applied consistently to the image and its polygon annotations, including random in-plane rotation, isotropic scaling, translation constrained by the instance extent, and random horizontal or vertical flipping. A random cutout was further used to simulate partial occlusion; polygon annotations were updated by geometrically clipping the visible region and retaining valid connected components. After each transformation, polygon validity was enforced by automatic closure and self-intersection repair, and invalid or degenerate shapes were discarded.
Sequence-level augmentation. Data augmentation was applied at the sequence level and restricted to spatial transformations that preserve the underlying growth dynamics. For each sequence, a single set of affine transformation parameters was sampled and applied identically to all frames to maintain temporal coherence. The augmentation pipeline was limited to in-plane rotation, translation, and random horizontal and vertical flipping. No augmentation was applied selectively to specific temporal segments or across time. Temporal resolution was defined solely by fixed-stride subsampling, without stochastic temporal perturbations.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (NSFC, Grant No. 12202275) and the Shanghai Jiao Tong University Explore X Fund.
Supplementary Information
S1 Performance comparison with state–of–the–art video prediction models under an 80%–20% observation–prediction split.
Table 1: Performance comparison with state–of–the–art video prediction models under an 80%–20% observation–prediction split. This table benchmarks Morpher against leading video prediction architectures, including MAU, MIM, PredRNN variants, and SimVP–based models. All methods are evaluated under identical input–output protocols for long–term forecasting of swarming colony expansion. Morpher achieves substantially higher region–level overlap (mIoU) and lower boundary error (HD95, ASSD), indicating improved accuracy in front propagation and boundary–level morphology.
Model mIoU (%) $\uparrow$ HD95 (px) $\downarrow$ ASSD (px) $\downarrow$
MAU [6] 84.67 22.73 14.68
MIM [46] 89.32 20.17 10.30
PredRNN [44] 84.60 22.75 14.81
PredRNNv2 [45] 84.14 23.24 15.04
SimVP [15]+TAU [40] 86.87 23.19 12.47
SimVP [15]+gSTA [39] 90.52 18.28 8.87
Morpher (Ours) 95.42 10.61 3.93
S2 Texture-Edge Attention and Polar–Context Attention modules
Swarming colony images exhibit complex morphological organization characterized by uncertain boundaries, irregular shapes, and radially propagating texture patterns. These inherent properties pose significant challenges for CNNs, whose reliance on local receptive fields restricts their capacity to capture long-range spatial dependencies and global geometric structures, particularly the near–concentric–ring radial expansion typical of swarming growth. To overcome these limitations and enhance both fine-grained texture extraction and high-level semantic representation, we developed two specialized attention modules, the Texture-Edge Attention (TEA) and Polar–Context Attention (PCA), as illustrated in Figure 6.
The TEA module, as shown in Figure 6LABEL:fig:TEA, is designed to address blurred boundaries and multi-scale, high-frequency texture variability. It combines three cooperative paths: a local branch that preserves intra-channel spatial details, a multi-dilated path that ensures scale-robust texture encoding, and an edge-sensitive path initialized with a discrete Laplacian kernel to enhance boundary awareness. Channel-wise and spatial gating mechanisms further refine the fused representation by emphasizing informative structures while maintaining computational efficiency.
Let the input be $\mathbf{X}\!\in\!\mathbb{R}^{B\times C_{\mathrm{in}}\times H\times W}$ and the output be $\mathbf{Y}\!\in\!\mathbb{R}^{B\times C_{\mathrm{out}}\times H\times W}$, where $B$ is the batch size, $C_{\mathrm{in}}$ and $C_{\mathrm{out}}$ denote the number of input and output channels, and $H$ and $W$ represent spatial dimensions. To balance representation capacity and computational cost, an intermediate channel width is introduced as
$C_{h}=C_{\mathrm{out}}\cdot e,$ (1)
where $e\in(0,1]$ is the expansion ratio controlling internal dimensionality.
Local features are first extracted using a depthwise $3\times 3$ convolution to capture intra-channel spatial structures, followed by a $1\times 1$ pointwise projection to $C_{h}$ channels to ensure dimensional consistency. The normalized and activated local features are
$\mathbf{F}_{\mathrm{loc}}=\phi\!\left(\mathrm{GN}\!\left(\mathrm{Conv}_{1\times 1}\!\left(\mathrm{Conv}^{\mathrm{dw}}_{3\times 3}(\mathbf{X})\right)\right)\right),$ (2)
where $\mathrm{Conv}^{\mathrm{dw}}_{3\times 3}$ denotes depthwise convolution, $\mathrm{GN}$ represents group normalization, and $\phi$ is the SiLU activation.
A squeeze–and–excitation (SE) gate $\mathbf{f}\!\in\!\mathbb{R}^{B\times C_{h}\times 1\times 1}$ is computed via global average pooling (GAP) followed by two $1\times 1$ convolutions with nonlinearity and sigmoid activation:
$\mathbf{f}=\sigma\!\Big(\mathrm{Conv}_{1\times 1}\big(\phi(\mathrm{Conv}_{1\times 1}(\mathrm{GAP}(\mathbf{X})))\big)\Big).$ (3)
Channel gating is applied element-wise:
$\widetilde{\mathbf{F}}_{\mathrm{loc}}(c,h,w)=\mathbf{F}_{\mathrm{loc}}(c,h,w)\odot\mathbf{f}(c).$ (4)
To model textures across multiple scales, a multi-dilated branch applies depthwise convolutions with dilation factors $d_{k}\!\in\!\mathcal{D}=\{d_{1},d_{2},\ldots,d_{K}\}$:
$\mathbf{F}^{(d_{k})}_{\mathrm{tex}}=\phi\!\left(\mathrm{GN}\!\left(\mathrm{Conv}_{1\times 1}\!\Big(\mathrm{Conv}^{\mathrm{dw}}_{3\times 3,d_{k}}(\mathbf{X})\Big)\right)\right),$ (5)
Figure 6: Texture-Edge Attention (TEA) and Polar–Context Attention (PCA) modules. a, The TEA block enhances fine-scale texture fidelity and boundary sharpness through three cooperative branches: a local depthwise path for intra-channel spatial preservation, multi-dilated convolutions for scale-robust texture encoding, and an edge-sensitive Laplacian path that injects a high-pass prior. Channel and spatial gating further refine feature fusion, producing an edge-aware, redundancy-suppressed representation. b, The PCA block embeds a polar-aware geometric prior aligned with the radial growth of swarming colonies. Input features are first compressed and then processed by a local branch, a large-kernel Cartesian branch, and a polar-warped branch operating in $(\rho,\theta)$ coordinates. Depthwise dilated filters extract context along radial and angular axes, and subsequent channel– and spatial–attention gates yield a geometry-aligned output.
and concatenation of all paths yields
$\mathbf{F}_{\mathrm{tex}}=\mathrm{Concat}\big(\mathbf{F}^{(d_{1})}_{\mathrm{tex}},\mathbf{F}^{(d_{2})}_{\mathrm{tex}},\ldots,\mathbf{F}^{(d_{K})}_{\mathrm{tex}}\big)$, (6)
where K=3K{=}3 captures short–, medium–, and long–range textures.
An edge-aware branch initialized by the Laplacian kernel enhances boundary sensitivity:
$\mathbf{F}_{\mathrm{edge}}=\phi\!\left(\mathrm{GN}\!\left(\mathrm{Conv}_{1\times 1}\!\Big(\mathrm{Conv}^{\mathrm{dw,Lap}}_{3\times 3}(\mathbf{X})\Big)\right)\right),$ (7)
with initialization
$\mathbf{K}_{\mathrm{Lap}}=\begin{bmatrix}0&-1&0\\[-2.0pt] -1&4&-1\\[-2.0pt] 0&-1&0\end{bmatrix}$. (8)
The texture and edge features are concatenated and reweighted by an SE gate:
$\mathbf{F}_{\mathrm{tex+edge}}=\mathrm{Concat}(\mathbf{F}_{\mathrm{tex}},\,\mathbf{F}_{\mathrm{edge}}),$ (9)
$\widetilde{\mathbf{F}}_{\mathrm{tex+edge}}=\mathbf{F}_{\mathrm{tex+edge}}\odot\sigma\!\Big(\mathrm{Conv}_{1\times 1}\big(\phi(\mathrm{Conv}_{1\times 1}(\mathrm{GAP}(\mathbf{F}_{\mathrm{tex+edge}})))\big)\Big).$ (10)
The SE-weighted outputs are combined with local features:
$\mathbf{F}_{\mathrm{fuse}}=\mathrm{Concat}(\widetilde{\mathbf{F}}_{\mathrm{loc}},\,\widetilde{\mathbf{F}}_{\mathrm{tex+edge}}).$ (11)
A spatial attention gate $\mathbf{g}\!\in\!\mathbb{R}^{B\times 1\times H\times W}$ emphasizes salient regions:
$\mathbf{g}=\sigma\!\Big(\mathrm{Conv}_{1\times 1}\big(\phi(\mathrm{Conv}^{\mathrm{dw}}_{3\times 3}(\mathbf{F}_{\mathrm{fuse}}))\big)\Big),$ (12)
and is applied element-wise:
$\bar{\mathbf{F}}_{\mathrm{fuse}}(c,h,w)=\mathbf{F}_{\mathrm{fuse}}(c,h,w)\odot\mathbf{g}(h,w).$ (13)
The fused representation is projected to the output dimension:
$\mathbf{Y}^{\prime}=\phi\!\left(\mathrm{GN}\!\left(\mathrm{Conv}_{1\times 1}(\bar{\mathbf{F}}_{\mathrm{fuse}})\right)\right),$ (14)
with a conditional residual connection:
$\mathbf{Y}=\begin{cases}\mathbf{X}+\boldsymbol{\gamma}\odot\mathbf{Y}^{\prime},&C_{\mathrm{in}}=C_{\mathrm{out}},\\[3.0pt] \mathbf{Y}^{\prime},&C_{\mathrm{in}}\neq C_{\mathrm{out}},\end{cases}$ (15)
where $\boldsymbol{\gamma}\!\in\!\mathbb{R}^{C_{\mathrm{out}}}$ is a learnable scaling factor that ensures stability when the dimensions are the same.
While TEA focuses on boundary and texture fidelity, the PCA module illustrated in Figure 6LABEL:fig:PCA captures long-range dependencies and radial geometric organization inherent to swarming colonies. Conventional convolutional operators struggle with near–concentric–ring propagation whereas PCA embeds a polar-aware representation aligned with colony growth.
The module comprises three paths: a local branch for spatial detail, a large-kernel branch for contextual encoding, and a polar branch that transforms features into polar coordinates for radial modeling. The local branch follows the same $3\times 3$ depthwise–pointwise pattern described in Eq. 2, operating on $\mathbf{X}^{\prime}$:
$\mathbf{F}_{\mathrm{local}}=\phi\!\left(\mathrm{GN}\!\left(\mathrm{Conv}_{1\times 1}\!\big(\mathrm{Conv}^{\mathrm{dw}}_{3\times 3}(\mathbf{X}^{\prime})\big)\right)\right).$
Input $\mathbf{X}\!\in\!\mathbb{R}^{B\times C_{\mathrm{in}}\times H\times W}$ is first compressed by a $1\times 1$ convolution to $C_{h}$ channels, normalized, and activated to yield $\mathbf{X}^{\prime}$. The large-context branch employs a depthwise separable $7\times 7$ convolution:
$\mathbf{F}_{\mathrm{large}}=\phi\!\left(\mathrm{GN}\!\left(\mathrm{Conv}_{1\times 1}\!\Big(\mathrm{Conv}^{\mathrm{dw}}_{7\times 7}(\mathbf{X}^{\prime})\Big)\right)\right).$ (16)
Spatial indices $(h,w)$ are mapped to polar coordinates:
$\theta_{w}$ $=\tfrac{2\pi w}{W},$ $w$ $\in\{0,\ldots,W{-}1\}$ (17)
$\rho_{h}$ $=\tfrac{h}{H-1},$ $h$ $\in\{0,\ldots,H{-}1\}$ (18)
and then transformed to normalized Cartesian coordinates:
$u(h,w)$ $=\rho_{h}\cos\theta_{w}$ (19)
$v(h,w)$ $=\rho_{h}\sin\theta_{w}$ (20)
Bilinear interpolation provides the polar-warped feature map:
$\mathbf{X}^{\mathrm{pol}}_{b,c,h,w}=\mathcal{I}\!\big(\mathbf{X}^{\prime}_{b,c,:,:};\,u(h,w),v(h,w)\big),\quad b\!\in\!\{0,\ldots,B{-}1\},\;c\!\in\!\{0,\ldots,C_{h}{-}1\},$ (21)
where $\mathcal{I}$ denotes the bilinear interpolation operator sampling $\mathbf{X}^{\prime}$ at polar coordinates $(u,v)$.
A depthwise $3\times 3$ convolution with dilation $d_{\mathrm{pol}}{=}4$ extracts polar-domain context:
$\mathbf{F}_{\mathrm{polar}}=\phi\!\left(\mathrm{GN}\!\Big(\mathrm{Conv}_{1\times 1}\big(\mathrm{Conv}^{\mathrm{dw}}_{3\times 3,\,4}(\mathbf{X}^{\mathrm{pol}})\big)\Big)\right).$ (22)
Finally, outputs from the three branches are concatenated:
$\mathbf{F}_{\mathrm{cat}}=\mathrm{Concat}(\mathbf{F}_{\mathrm{local}},\,\mathbf{F}_{\mathrm{large}},\,\mathbf{F}_{\mathrm{polar}}),$ (23)
followed by channel SE and spatial attention (Eq. 313) to produce $\widetilde{\bar{\mathbf{F}}}_{\mathrm{cat}}$. Projection and conditional residual (Eq. 1415) yield the final PCA output. This design preserves fine structural details, integrates global context, and explicitly embeds radial priors, enabling robust modeling of colony expansion dynamics.
S3 Performance of Morpher under an 80%–20% observation–prediction split across temporal modeling and inference
paradigms
Table 2: Performance of Morpher under an 80%–20% observation–prediction split across temporal modeling and inference paradigms. Forecasting accuracy is evaluated across region-level overlap (mIoU), boundary accuracy (HD, HD95, ASSD), front–propagation dynamics (RMSE), temporal fluctuation consistency (TCI), and angular growth organization ($\lvert\Delta\mathrm{NAS}\rvert$, $\lvert\Delta\mathrm{H}_{2}\rvert$). Higher mIoU and TCI indicate superior forecasting performance, whereas lower HD-based distances, RMSE, $\lvert\Delta\mathrm{NAS}\rvert$, and $\lvert\Delta\mathrm{H}_{2}\rvert$ reflect improved geometric and dynamical fidelity. This table provides a mechanistic comparison by isolating the effects of temporal modeling, inference strategy, and the Morphon memory mechanism.
Seq. Model Inference Paradigm Morphon mIoU (%) $\uparrow$ HD (px) $\downarrow$ HD95 (px) $\downarrow$ ASSD (px) $\downarrow$ RMSE (px/frame) $\downarrow$ TCI (%) $\uparrow$ $\lvert\Delta\mathrm{NAS}\rvert$ (%) $\downarrow$ $\lvert\Delta\mathrm{H}_{2}\rvert$ (%) $\downarrow$
RNN Parallel 93.23 17.68 12.85 6.02 3.36 55.34 19.54 1.96
LSTM Parallel 93.24 17.44 12.65 5.85 2.84 60.48 15.99 1.87
GRU Parallel 93.96 17.14 12.51 5.17 3.02 54.80 15.31 1.79
Transformer Parallel 93.94 17.24 12.64 5.23 2.95 56.94 16.83 1.94
RNN Autoregr. 93.55 17.53 12.51 5.26 2.20 65.63 15.02 1.90
LSTM Autoregr. 94.07 16.92 11.85 5.34 2.60 63.14 17.80 1.91
GRU Autoregr. 94.20 17.25 12.03 5.01 2.39 64.10 19.02 1.84
Transformer Autoregr. 94.16 16.56 11.95 5.26 2.66 64.92 13.51 1.74
RNN Parallel 94.22 16.67 11.63 4.85 2.85 61.57 17.57 1.84
LSTM Parallel 94.44 15.97 10.90 4.59 2.76 59.29 14.72 1.69
GRU Parallel 94.58 15.86 11.46 4.81 2.68 63.14 15.56 1.76
Transformer Parallel 94.80 15.79 11.22 4.63 2.55 62.71 16.30 1.89
RNN Autoregr. 94.94 15.46 10.67 4.34 2.19 63.32 16.87 1.86
LSTM Autoregr. 95.01 15.32 10.77 4.20 2.26 65.32 15.03 1.79
GRU Autoregr. 95.29 15.01 10.14 4.06 2.06 64.31 15.45 1.89
Transformer Autoregr. 95.42 15.26 10.61 3.93 2.12 64.26 13.13 1.85
S4 Performance of Morpher under a series of observation–prediction splits across sequence models
Table 3: Performance of Morpher under a series of observation–prediction splits across sequence models. Results are reported for 50%, 60%, 70%, 80%, and 90% observation levels to assess how forecasting stability changes as more of the past is revealed. Metrics include region-level overlap (mIoU), boundary accuracy (HD, HD95, ASSD), front–propagation dynamics (RMSE), temporal fluctuation consistency (TCI), and angular growth organization ($\lvert\Delta\mathrm{NAS}\rvert$, $\lvert\Delta\mathrm{H}_{2}\rvert$). Higher mIoU and TCI indicate superior forecasting performance, whereas lower HD-based distances, RMSE, $\lvert\Delta\mathrm{NAS}\rvert$, and $\lvert\Delta\mathrm{H}_{2}\rvert$ reflect improved geometric and dynamical fidelity.
Observation (%) Seq. Model mIoU (%) $\uparrow$ HD $\downarrow$ HD95 $\downarrow$ ASSD $\downarrow$ RMSE (px/frame) $\downarrow$ TCI (%) $\uparrow$ $\lvert\Delta\mathrm{NAS}\rvert$ (%) $\downarrow$ $\lvert\Delta\mathrm{H}_{2}\rvert$ (%) $\downarrow$
50 RNN 87.88 25.38 20.92 10.19 2.56 62.77 13.42 1.14
LSTM 88.18 25.15 20.37 9.43 2.20 62.97 14.28 1.21
GRU 87.99 25.89 21.35 10.02 2.42 62.71 14.37 1.15
Transformer 88.22 23.64 19.12 9.49 2.28 63.79 14.62 1.18
60 RNN 92.20 19.76 15.35 6.29 2.10 64.09 12.04 1.33
LSTM 91.76 20.26 15.38 6.71 2.22 64.34 14.46 1.38
GRU 92.24 19.48 15.30 6.33 2.03 64.13 12.98 1.37
Transformer 92.64 18.21 14.18 5.96 1.96 63.66 11.93 1.35
70 RNN 93.18 18.02 13.34 5.64 2.29 64.89 17.55 1.54
LSTM 93.40 18.58 13.57 5.41 2.18 64.54 14.64 1.47
GRU 93.37 18.57 13.78 5.59 2.25 64.49 16.83 1.52
Transformer 93.80 16.88 12.18 4.95 2.12 62.42 15.57 2.25
80 RNN 94.94 15.46 10.67 4.34 2.19 63.32 16.87 1.86
LSTM 95.01 15.32 10.77 4.20 2.26 65.32 15.03 1.79
GRU 95.29 15.01 10.14 4.06 2.06 64.31 15.45 1.89
Transformer 95.42 15.26 10.61 3.93 2.12 64.26 13.13 1.85
90 RNN 96.02 12.48 8.36 3.13 2.24 13.24 2.20
LSTM 96.31 12.45 8.19 2.95 2.08 13.59 2.33
GRU 96.42 12.09 8.38 3.07 2.09 13.86 2.24
Transformer 96.79 11.20 7.91 2.75 2.07 13.96 2.21
  • No TCI is reported at 90% observation, because the remaining number of frames is insufficient to obtain a reliable estimate.

### v2

@@
 Microbial swarming on mucosal surfaces reshapes microbial communities and influences mucosal healing and antibiotic tolerance.
-Yet even with time-lapse microscopy and deep learning, analyses of swarming colonies remain descriptive and cannot forecast how their fronts reorganize in time.
-This limitation is significant because the advancing edge determines access to nutrients, host tissue and competing microbes.
-We recast the expansion of Enterobacter sp.
-SM3 swarms as a problem of morphological forecasting, and assemble SwarmEvo, a time-lapse dataset represented as boundary-resolved segmentations.
-TexPol–Net, a texture- and geometry-aware segmentation model, sharpens diffuse edges and preserves fingered fronts, creating a stable substrate for dynamics.
-On this representation, we develop Morpher, an autoregressive forecasting network with a “Morphon” memory that links local curvature to long-range temporal dependencies.
+Yet even with time-lapse microscopy and deep learning, analyses of swarming colonies remain largely descriptive and lack accurate forecasting of front reorganization.
+This limitation is significant because the advancing edge determines access to nutrients, host tissues, and competing microbes.
+The expansion of Enterobacter sp.
+SM3 swarms is recast as a problem of morphological forecasting, and SwarmEvo is assembled as a time-lapse dataset of boundary-resolved segmentations.
+TexPol–Net, a texture- and geometry-aware segmentation model, sharpens diffuse edges and preserves finger-like fronts, creating a stable substrate for dynamics.
+On this representation, Morpher is developed as an autoregressive forecasting network with a “Morphon” memory that links local curvature to long-range temporal dependencies.
 Morpher outperforms leading video-prediction models in maintaining front localization and anisotropic branching, and modest segmentation improvements yield noticeably more stable forecasts.
-Ablations across sequence models, inference strategies and observation ratios show that attention-based architectures with structural memory best preserve dense-finger propagation.
+Ablations across sequence models, inference strategies, and observation ratios show that attention-based architectures with structural memory best preserve branching propagation dynamics.
 By uniting geometry-aware segmentation with morphology-level forecasting, this framework turns swarming expansion into a predictive dynamical system, enabling quantitative interrogation and potential control of microbial collectives during mucosal repair and gut ecosystem engineering.
+Keywords: bacterial swarming; morphological forecasting; collective behavior; spatiotemporal dynamics; deep learning; predictive modeling
+Significance Statement Swarming bacteria reshape their environment through a moving front that governs access to oxygen, host tissue, and competing microbes.
+Yet this front has largely been described rather than predicted.
+We show that swarming trajectories are not organized by experimental conditions but by colony-specific dynamics, shifting prediction away from condition labels toward direct measurement of the front itself.
+By recasting the front as a geometric state whose evolution can be forecast, our framework links measurement fidelity to predictive stability.
+This establishes colony morphology as a quantitative, forward-looking variable and opens a path toward anticipating—and ultimately controlling—collective microbial behavior.
 1 Introduction
 Microbial communities form dense and continually reorganizing ecological networks within and around animal hosts [40, 10, 4, 69, 55, 53].
-Their spatial organization and collective movement govern population expansion, interspecies interactions and tissue homeostasis under both healthy and diseased conditions [17, 26, 22].
-Among these behaviors, collective surface motility on semi–solid substrates—classically termed swarming [1, 5]—produces continually reconfiguring fronts and patterned colony architectures that both reflect coordinated cellular behaviors and reshape local host environments [11, 49, 32].
-Yet despite extensive descriptive work, the spatiotemporal evolution of such structures remains poorly characterized.
-In particular, forecasting the future morphology of a swarming colony—one of the most dynamic forms of bacterial collective motion—is currently out of reach.
-A predictive framework that links present colony architecture to its future configurations would transform morphology from a descriptive endpoint into a quantitatively interrogable variable, enabling mechanistic studies of how microbial collectives respond to environmental and physiological cues.
-Attempts to analyze colony growth have historically been constrained by low imaging frequency and qualitative interpretation [24, 28].
-Classical computer–vision pipelines partially automated colony detection and quantification [4, 9, 37, 51], but they tend to fail when boundaries blur, textures reorganize or colonies overlap.
-More recent analyses of active suspensions and deformable colony–fluid interfaces have revealed turbulence–like flows and curvature–dependent edge dynamics [18, 48].
-However, these approaches often presume approximate symmetry or reduce colony evolution to low–dimensional order parameters, limiting their applicability to the irregular, anisotropic and burst–like propagation characteristic of swarming fronts.
-In parallel, deep learning has expanded biological image analysis, yielding convolutional neural network (CNN)–based colony detectors [14, 47], temporal classifiers of motility states [31] and hybrid pipelines that combine growth–curve inference with object detection [29].
-Advances in imaging, such as coherent time–lapse microscopy for early species identification [43] and engineered swarming biosensors [12], further underscore the morphological richness encoded in colony structure.
-Yet these developments remain essentially descriptive: they identify colonies, classify motility or extract bulk parameters, but they do not predict how a colony’s contour will evolve.
-Even models trained to infer motility type from single frames [23] collapse inherently dynamic processes into static labels and therefore offer limited insight into future behavior.
-A central obstacle is that swarming colonies violate the appearance continuity implicitly assumed by most natural–scene video models.
-Fronts propagate through intermittent bursts, transient asymmetries and rapid reorganizations of multi–scale texture.
-Curvature modulates local velocity; protrusions emerge and retract discontinuously; textures shift without preserving pixel–level coherence.
-Under these conditions, extrapolating visual appearance is neither stable nor biologically grounded.
-Predicting future morphology instead requires reasoning over geometric and biophysical representations of the colony boundary and its local environment.
-Such forecasts would enable in silico perturbation experiments, real–time anticipation of responses to surface biochemistry or antimicrobial stress, and strategies to steer collective migration in applications ranging from mucosal repair to microbial therapeutics and microbiome engineering.
-Predictions of front position and branching patterns implicitly reflected experimental responses to mucin gradients and antimicrobial stress, indicating that the learned representations encode biologically meaningful adaptation rather than merely geometric similarity.
-These questions crystallize in Enterobacter sp.
-SM3, a representative swarming commensal from the murine gut [33, 8].
-SM3 reshapes intestinal microbial organization and promotes mucosal healing, whereas mutants impaired in swarming lose these beneficial functions [11].
-More broadly, swarming has been implicated in antibiotic tolerance, regulation of virulence factors and robust colonization of host–associated surfaces [21, 30], with its dynamics strongly modulated by surface biochemical cues such as mucin [32].
-In inflamed intestinal environments, the advancing swarming front is the actionable interface: its future position specifies where oxygen is depleted and anaerobic niches emerge, shaping downstream microbial reorganization and recovery, as shown in Fig.
-1.
-Anticipating this position in advance therefore provides a principled basis for spatially and temporally targeted intervention, through pharmacological or environmental modulation of the local microenvironment.
-Despite this conceptual clarity, translating it into predictive capability remains an open challenge.
-No framework currently exists for forecasting the evolving contours of swarming colonies at the resolution of individual fronts.
-This gap limits our ability to quantify how environmental or physiological signals shape collective migration and to connect microscopic motility programs to macroscopic patterns in a predictive manner.
-Here we address this gap by formulating swarming colony expansion as a problem of morphological forecasting in a geometric state space.
-We first assemble the Swarming Morphogenesis Evolution (SwarmEvo) dataset, a curated collection of high–resolution time–lapse recordings of Enterobacter sp.
-SM3 spreading across semi–solid substrates under systematically varied conditions.
-Each recording is processed into boundary–resolved segmentation sequences that capture the evolving colony contour, providing both a biologically faithful account of swarming dynamics and a quantitative benchmark for evaluating whether computational models can maintain geometric fidelity over time.
-Building on this representation, we develop TexPol–Net, a segmentation framework that integrates texture- and geometry-aware priors to recover high–fidelity fronts even when boundaries are diffuse or heavily reorganized, thereby establishing a stable morphological state space for temporal prediction.
-We then introduce Morpher, a spatiotemporal model that treats forecasting as contour evolution within this state space, combining local geometric cues with long–range temporal dependencies to capture both smooth curvature modulations and rapid reorganization of protrusions.
-By establishing a geometric representation of colony expansion and a forecasting model capable of operating within it, SwarmEvo, TexPol–Net and Morpher together recast swarming from a descriptive phenomenon into a tractable dynamical system with predictable trajectories.
-This framework enables quantitative interrogation of how microbial collectives respond to environmental and physiological cues with a temporal resolution beyond the reach of current experimental tools, and it lays an analytical foundation for integrating morphological forecasting with real-time imaging and for interpreting and steering microbial collective behavior in vivo.
+Their spatial organization and collective movement shape population expansion, interspecies interactions, and tissue homeostasis in both health and disease [23, 39, 33, 20].
+Among these behaviors, collective surface motility on semi-solid substrates—classically termed swarming [1, 8, 66, 29, 3, 6]—produces continually reconfiguring fronts and colony architectures that both reflect coordinated cellular behavior and actively remodel the local host environment [14, 68, 45, 25].
+Although classic studies of swarm pattern formation, cell-state differentiation, and periodic colony expansion documented rich phenomenology [50, 49, 24, 28], they stopped short of treating front evolution as a predictive problem.
+We still lack a framework that can predict how a swarming colony will change shape over time, a capability that would turn morphology from a visual endpoint into a quantitative variable for probing how microbial collectives respond to environmental and physiological cues.
+This gap persists despite substantial progress in colony analysis.
+Early studies were limited by sparse imaging and qualitative interpretation [35, 41].
+Classical computer-vision pipelines improved colony detection and counting [7, 12, 54, 71, 2, 70], but they deteriorate when boundaries blur, textures reorganize, or colonies overlap.
+Physical analyses of active suspensions and deformable colony interfaces have revealed turbulence-like flows and curvature-dependent edge dynamics [26, 65], yet they often assume approximate symmetry or compress growth into low-dimensional summaries that cannot capture the irregular, anisotropic, and burst-like propagation of swarming fronts.
+Deep learning has broadened biological image analysis [72, 37, 19, 48] through CNN-based colony detectors [17, 64], temporal classifiers of motility states [44], and hybrid detection-growth pipelines [42].
+Advances in imaging, including coherent time-lapse microscopy for early species identification [60] and engineered swarming biosensors [15], further emphasize how much biological information is encoded in colony morphology.
+But these approaches remain fundamentally descriptive: they detect, classify, or summarize; they do not forecast contour evolution.
+Even single-frame predictors of motility type [34] collapse a dynamic process into a static label.
+A central reason is that swarming violates the appearance continuity assumed by most natural-scene video models.
+Fronts advance through intermittent bursts, transient asymmetries, and rapid multiscale reorganization.
+Curvature modulates local speed; protrusions emerge and retract discontinuously; texture shifts without preserving pixel-level coherence.
+Under these conditions, extrapolating appearance is neither stable nor biologically persuasive.
+What carries the future is the front itself: its geometry, branching structure, and temporal continuity.
+In spatially expanding microbial populations, the advancing frontier constitutes a thin, dynamically active region that governs large-scale structure while the interior remains effectively frozen [21].
+A useful predictive model must therefore resolve colony-specific morphology rather than rely on coarse labels, bulk summaries, or framewise appearance alone.
+Forecasting future shape in this setting is not an image-generation problem but a problem of front dynamics.
+This challenge is especially consequential in Enterobacter sp.
+SM3, a representative swarming commensal from the murine gut [47, 11, 27].
+SM3 reshapes intestinal microbial organization and promotes mucosal healing, whereas swarming-deficient mutants lose these beneficial effects [14].
+More broadly, swarming has been linked to antibiotic tolerance, virulence regulation, and robust colonization of host-associated surfaces [32, 43, 46], and its dynamics are strongly modulated by surface biochemical cues such as mucin [45].
+In inflamed intestinal environments, the advancing swarming front is the actionable interface: its future position determines where oxygen is depleted and anaerobic niches emerge, thereby shaping downstream community reorganization and recovery (Figure 1).
+Anticipating that front in advance would provide a principled basis for spatially and temporally targeted intervention.
+Yet no framework currently forecasts swarming morphology at the resolution of individual fronts, leaving a gap between microscopic motility programs and macroscopic pattern evolution.
+Here we address that gap by recasting swarming colony expansion as a problem of morphological forecasting in a geometric state space.
+We assemble the Swarming Morphogenesis Evolution (SwarmEvo) dataset, a time-lapse resource of Enterobacter sp.
+SM3 expanding across systematically varied semi-solid environments.
+We then recover boundary-resolved colony states with TexPol–Net, designed to preserve diffuse but biologically important front structures, and forecast their evolution with Morpher, a spatiotemporal model that treats prediction as contour dynamics rather than future-image synthesis.
+By linking boundary measurement to long-horizon front forecasting, this framework reframes swarming as a measurable and predictable dynamical process, opening a route toward quantitative interrogation—and ultimately control—of microbial collective behavior in living environments.
 2 Results
-2.1 A two-stage formulation of swarming morphology
-The swarming expansion of Enterobacter sp.
-SM3 is characterized by pronounced morphological plasticity.
-Under identical assay conditions, colonies may evolve through irregular, dense finger-like protrusions driven by curvature instabilities at the advancing front, or settle into near-concentric growth characterized by a stabilized rim and weak directional anisotropy.
-These recurring regimes define a structured spectrum of morphogenetic behavior, spanning fine-scale boundary fluctuations and colony-scale organization (Fig.
-1).
+2.1 A state-based formulation of swarming morphology
+Swarming expansion in Enterobacter sp.
+SM3 spans a reproducible morphological spectrum.
+Even under the same experimental conditions, colonies form two distinct morphogenetic regimes: an anisotropic branching regime (finger-like fronts) and a near-concentric regime (approximately isotropic expansion) (Figure 1).
+These regimes reflect structured morphogenetic states rather than frame-level visual fluctuations.
+We therefore formulate swarming as a two-stage problem: first, measuring the advancing front as a geometric state; second, forecasting the evolution of that state.
+Time-lapse images are converted into boundary-resolved masks that preserve protrusions while suppressing appearance variability, yielding a stable representation across time.
+Forecasting is then posed as front evolution rather than future-image synthesis.
+This shift moves the problem from appearance prediction to state evolution, with morphology serving as a compact state descriptor of colony dynamics.
 Figure 1: From swarming dynamics to predictive guidance.
-In a mouse model of colitis, Enterobacter sp.
-SM3 swarms along the inflamed mucosal surface.
-Under pharmacological, environmental, or microbiota-based modulation, forecasting the future position of the active swarming front may enable spatially targeted and temporally coordinated interventions.
-Such predictive capability could be coupled with experimental strategies that locally modulate oxygen availability or microbial competition at the advancing front, with potential effects on anaerobic niche formation and downstream community composition, including Muribaculaceae expansion.
-Time-lapse swarming assays are converted into boundary-resolved morphological masks by TexPol–Net, with Texture-Edge and Polar–Context attention, shared prototypes and mask assembly, defining observed morphological states.
-Morpher models their temporal evolution with a sequence backbone and Morphon memory to predict future morphologies for quantitative evaluation.
-Panels I–IV depict an application-oriented biological and experimental pipeline, whereas the modeling framework follows a two-stage formulation of boundary measurement and morphology-level forecasting.
-To resolve these dynamics, we adopt a two-stage formulation that separates morphological measurement from temporal prediction.
-In Panel III (Fig.
-1), time-lapse microscopy is converted into sequences of boundary-resolved morphological masks using TexPol–Net.
-Rather than serving as a generic segmentation backbone, TexPol–Net acts as a measurement operator that constrains the geometric degrees of freedom of the colony front, preserving thin protrusions and internal texture while suppressing illumination-induced variability.
-The resulting mask sequences constitute a compact and biologically interpretable representation of colony state across time.
-Panel IV (Fig.
-1) treats these observed morphologies as a temporal signal and models their evolution using Morpher.
-Operating directly in morphology space, Morpher infers latent dynamical states from the observed window and extrapolates future colony shapes beyond it.
-This formulation shifts the forecasting task from image synthesis to the prediction of boundary evolution, allowing local perturbations along the front to propagate coherently over long horizons.
-Autoregressive conditioning further permits predicted morphologies to influence subsequent growth states, reflecting the cumulative nature of swarming expansion.
-By decoupling spatial characterization from temporal dynamics, the framework limits error accumulation from appearance noise and isolates the mechanisms governing shape evolution.
-This structure supports systematic evaluation of forecasting models across distinct morphogenetic regimes, from highly irregular fingered fronts to stabilized radial expansion, within a unified representation that preserves biologically relevant geometric structure.
-2.2 Recovering colony fronts for accurate morphological forecasting
-Accurate delineation of the advancing front is the prerequisite for morphological forecasting, as it encodes curvature, anisotropy, and the local direction of expansion.
-We therefore evaluated how different segmentation backbones reconstruct colony fronts across representative growth regimes, and quantitatively examined how boundary accuracy evolves as matching criteria become increasingly stringent (Fig.
-3).
-In the dense–finger regime, the front is highly corrugated and places stringent demands on boundary fidelity (Fig.
-3b).
-YOLOv11 is able to follow the global outline, but tends to break slender fingers into shortened or discontinuous segments, effectively smoothing high-curvature protrusions and weakening the anisotropic structure of the advancing front.
-SAM-based models instead emphasize the central colony mass and suppress distal protrusions, causing the boundary to evolve toward a smoother, near-circular shape.
-In contrast, TexPol–Net remains aligned with the annotated front along both the outer envelope and fine branches, preserving uneven finger lengths and the radial bias characteristic of early expansion; YOLOv12 is the closest competitor in this regime.
+In colitis, Enterobacter sp.
+SM3 swarms along the inflamed mucosal surface, where the advancing front governs access to oxygen and microbial competition.
+Anticipating its future position enables spatially and temporally targeted intervention.
+Time-lapse assays are converted by TexPol–Net into boundary-resolved colony states, and Morpher forecasts their evolution.
+This framework links the biological problem to a two-stage formulation of boundary measurement and morphology-level prediction.
+2.2 Individual variability in swarming dynamics
+Before evaluating forecasting models, we asked whether swarming trajectories are cleanly organized by assay conditions.
+They are not.
+In PCA of trajectory-level perimeter and area features, colonies measured under different temperatures, humidities, and agar concentrations remain extensively overlapped rather than forming distinct clusters (Figure 2a,b).
+The same pattern persists at the distribution level: for both perimeter and area trajectories, within- and across-condition pairwise distances nearly coincide, with overlap coefficients remaining high (0.92–0.97) and Cliff’s $\delta$ close to zero across all three variable groupings (Figure 2c,d).
+Figure 2: Large individual variability and weak separability in swarming morphology.
+a,b PCA of trajectory-level perimeter and area features from 81 colonies across temperature, humidity, and agar concentration.
+Broad overlap indicates weak separability by nominal condition.
+c,d Pairwise-distance distributions for perimeter and area trajectories under the same groupings.
+Overlap between within- and across-condition distances shows that individual variation is comparable to condition-level differences.
+Swarming trajectories are not organized by nominal experimental conditions.
+Instead, variation is dominated by colony-specific dynamics, with within- and across-condition distances remaining statistically indistinguishable.
+This eliminates condition labels as a predictive axis and shifts the problem to geometry-resolved state evolution.
+2.3 Segmentation defines the predictive state
+Because forecasting depends on front geometry, we first evaluated how accurately different segmentation backbones reconstruct colony boundaries across representative growth regimes (Figure 3b).
+The clearest separation appears in the anisotropic branching regime, where boundary fidelity is most demanding.
+YOLOv11 captures the global outline but shortens or fragments slender fingers.
+SAM and SAM2 preserve the colony core while suppressing distal protrusions, driving the front toward a smoother, more circular contour.
+TexPol–Net remains aligned with both the outer envelope and fine branches, with YOLOv12 as the closest competitor.
+Figure 3: TexPol–Net improves colony-front segmentation by coupling texture-sensitive boundary encoding with a geometry-aligned context prior.
+a TexPol–Net architecture within a prototype-based instance segmentation pipeline.
+TEA in the backbone preserves boundary texture, and PCA in the bidirectional neck maintains polar consistency during multi-scale fusion.
+b Qualitative comparison on representative anisotropic branching and near-concentric regimes.
+TexPol–Net better preserves branch heterogeneity and boundary position than YOLOv11 [30], SAM [31], and SAM2 [51], with YOLOv12 [58] as the closest competitor.
+c AP as a function of IoU threshold on the SwarmEvo segmentation test set (78 colonies), highlighting stronger performance under boundary-stringent matching.
+d Image-wise IoU distribution on the same test set, summarizing accuracy across samples.
+e Image-wise Dice distribution on the same test set, reflecting consistency and variability of segmentation quality.
 […]
-Figure 3: TexPol–Net improves colony-front segmentation by coupling texture-sensitive boundary encoding with a geometry-aligned context prior.
-a, Texture-Polar Network (TexPol–Net) within a prototype-based instance segmentation pipeline.
-Texture-Edge Attention (TEA) is embedded in the backbone to preserve high-frequency boundary cues, while Polar–Context Attention (PCA) is interleaved in the PANet-style bidirectional neck to maintain polar consistency during multi-scale fusion.
-Dense heads at P3P3P5P5 predict class scores, boxes and mask coefficients; a lightweight Protonet generates k=32k{=}32 shared prototypes that are linearly combined into instance masks and post-processed by cropping and thresholding.
-b, Qualitative comparison on two representative swarming regimes.
-In the dense–finger phase (top), YOLOv11 [19] captures the coarse outline but smooths the corrugated front, shortening narrow fingers and reducing anisotropy.
-SAM [20] and SAM2 [34] favor the colony core, producing a boundary that approaches a near-circular shape.
-TexPol–Net preserves finger heterogeneity and radial bias, with YOLOv12 [41] as the closest competing backbone.
-In the near–concentric ring regime (bottom), all methods recover the circular structure, while TexPol–Net more precisely localizes the advancing boundary.
-c, Average precision as a function of IoU matching threshold, highlighting boundary-stringent performance at high IoU.
-d, Image-wise overlap statistics for IoU and Dice, summarizing accuracy and variability across images.
-Accurate recovery of the advancing front is a structural prerequisite for morphological forecasting rather than a superficial preprocessing step.
-Errors introduced at the segmentation stage propagate directly into the temporal model, as smoothing finger-like protrusions, displacing the outer rim, or introducing spurious fragments alters the effective front trajectories available for learning.
-By preserving fine-scale protrusions while maintaining the global organization of the rim, TexPol–Net yields a geometrically consistent sequence of morphological masks in which successive fronts are linked through physically plausible deformations.
-This representation defines an appropriate input substrate for curvature-driven expansion and underpins the forecasting analyses presented in the following sections.
-2.3 Forecasting the expanding front of swarming colonies
-We conduct a systematic comparison between Morpher and representative state–of–the–art video prediction models, including MAU [6], MIM [46], PredRNN [44], PredRNNv2 [45], the original SimVP implementation which employs the TAU temporal unit [15, 40], and the improved SimVPv2 variant, in which TAU is replaced by the gSTA module [39].
-These models have demonstrated strong performance on standard video prediction benchmarks and are widely regarded as effective in modeling temporal dependencies and spatial feature organization.
-However, unlike natural video sequences where temporal variation is dominated by local texture dynamics and intensity changes, swarming colony expansion is driven primarily by the propagation of morphological fronts, the generation and transmission of local geometric perturbations, and the temporal evolution of the global expansion velocity field.
-Consequently, the ability to accurately track front curvature, scale variation, and perturbation continuity is far more critical than minimizing pixel–wise reconstruction error.
-To ensure a fair comparison, all models were retrained from scratch under identical data partitions, observation/prediction splits, and evaluation protocols, and their outputs were evaluated using the mIoU, HD95, and ASSD metrics, as shown in Fig.
-2.3a.
-The results indicate that existing video prediction models exhibit clear limitations in this forecasting setting.
-MIM and SimVP+gSTA achieve relatively high region–level overlap, with mIoU of 89.32% and 90.52%, respectively, suggesting that they are able to preserve the overall expansion pattern to some extent.
-However, both models perform suboptimally in terms of HD95 and ASSD, indicating persistent discrepancies between the predicted and true colony fronts.
-In particular, PredRNN and PredRNNv2, although designed to enhance temporal memory and long–range dependency modeling, show consistently higher HD95 and ASSD values than the other models.
-This suggests that their hidden–state–accumulation–based recurrent formulation is susceptible to error amplification during extrapolation, making it difficult to maintain accurate front localization and stable representation of local geometric details over extended prediction horizons.
-In addition, SimVP+TAU demonstrates moderate performance across all three metrics, failing to strike an effective balance between region–level consistency and boundary precision.
-In contrast, Morpher achieves the best performance across all evaluation metrics, with a mIoU of 95.42%, an HD95 of 10.61 pixels and an ASSD of 3.93 pixels, representing a 5.4% gain in region–level overlap and reductions of 42.0% and 55.7% in HD95 and ASSD relative to the strongest baseline, SimVP+gSTA.
-Although only the final 20% of each sequence requires extrapolative prediction, this window lies in the late stage of colony expansion, where the advancing interface is large and highly sensitive; even small deviations grow into disproportionately large discrepancies as the front progresses.
-The marked decreases in HD95 and ASSD show that Morpher not only aligns with the true expansion pattern but also preserves the geometry of the propagating front as it deforms.
-Each percentage point of mIoU thus corresponds to a tangible gain in propagation stability and in the maintenance of morphological integrity.
-By retaining the trajectory of the advancing edge with greater fidelity, Morpher produces forecasts that are more consistent with the underlying biological process of swarming expansion and capture features that matter for nutrient access, epithelial protection and microbial competition.
+2.4 Limits of appearance-based prediction
+Morphology-level forecasting is ultimately judged by whether the active front is localized and whether its fine branches are preserved.
+Region overlap (mIoU) can remain high while the contour drifts; we therefore treat boundary-sensitive distances (HD95 and ASSD) as the primary indicators of predictive fidelity, and use overlap as a complementary check on coarse extent (Figure 4a).
+We compared Morpher with representative video prediction models, including MAU [9], MIM [63], PredRNN [61], PredRNNv2 [62], the original SimVP implementation with the TAU temporal unit [18, 57], and the improved SimVPv2 variant with the gSTA module [56], all retrained under identical splits and evaluated on the same 80% observation / 20% prediction protocol (Figure 4a).
+This ordering holds across all metrics.
+MIM and SimVP+gSTA preserve coarse regional extent to some degree, reaching mIoU values of 89.32% and 90.52%, but remain substantially worse in boundary accuracy.
+PredRNN and PredRNNv2 exhibit still larger HD95 and ASSD, indicating stronger drift during extrapolation.
+SimVP+TAU remains intermediate without resolving the overlap–boundary tradeoff.
 Figure 4: Failure modes of generic video prediction in swarming morphology forecasting and the role of morphology-aware representation.
-a, Performance comparison under an 80% observation / 20% prediction protocol.
-State-of-the-art video prediction models are evaluated under identical input–output settings for long-horizon forecasting of swarming colony expansion.
-While several methods attain reasonable region-level overlap, boundary-related errors remain pronounced across models.
-Morpher achieves higher mIoU together with consistently lower HD95 and ASSD, reflecting improved accuracy in front propagation and boundary geometry.
-b, Long-horizon qualitative forecasts on two representative sequences spanning distinct growth regimes.
-In the dense–finger phase (top), generic predictors either progressively smooth high-curvature fingers or fail to follow the true front dynamics, leading to increasing deviation from the ground truth.
-In the late-stage near–concentric ring expansion (bottom), maintaining global radial scale and subtle curvature becomes challenging for these models.
-Across both regimes, Morpher maintains a coherent and temporally consistent front geometry over extended prediction horizons, remaining closer to the true front.
-c, Effect of segmentation backbone on forecasting performance.
-Using TexPol–Net masks leads to more stable and accurate final-frame forecasts, whereas YOLOv12 [41], the strongest alternative backbone in this task, produces competitive results with residual boundary deviation and curvature attenuation.
-In the final predicted frame, TexPol–Net achieves higher IoU than YOLOv12 across both growth regimes, with improvements from 94.21% to 96.63% in the dense–finger phase and from 93.30% to 96.38% in the near–concentric ring expansion.
-To further corroborate the quantitative results, we compare the predicted colony evolution across representative growth stages, as illustrated in Fig.
-2.3b.
-During the dense–finger–instability–dominated expansion phase, MIM displays a clear tendency toward morphological smoothing over time: fine–scale protrusions at the advancing front gradually collapse into low–curvature boundaries, suppressing the inherent radial anisotropy and attenuating the transmission of local perturbations, which in turn compromises the accuracy of global expansion characterization over extended prediction horizons.
-The SimVP variants employing TAU or gSTA temporal units preserve the overall outline of the colony to a certain extent, but fail to maintain the dense–finger morphology, frequently exhibiting truncated lobe tips, blurred curvature transitions, and unstable front propagation rates; their predicted boundaries also become progressively rounded, reflecting insufficient geometric resolution and resulting in noticeable boundary drift.
-MAU and the recurrent architectures PredRNN and PredRNNv2 generate shapes that are superficially close to the ground truth in terms of global geometry, yet their temporal evolution is consistently slower, characterized by delayed dense–finger development, systematic underestimation of front propagation speed, and reduced amplitude of perturbation transmission, producing a temporal mismatch in both radial scale and finger–scale development.
-In contrast, Morpher yields stable and coherent predictions across growth regimes, maintaining both the spatial continuity of local perturbations and the global advancement and curvature modulation of the expanding front, while preserving a stable advancing interface without loss of curvature transmission or propagation continuity.
-In the second sequence, the colony enters a late–stage expansion regime with a predominantly disc–like symmetry and a stabilized advancing front.
-In this context, the forecasting challenge shifts from preserving dense–finger morphology to maintaining the global radial scale and subtle boundary curvature modulation.
-MAU, MIM, PredRNN, and PredRNNv2 consistently underestimate the overall expansion radius, producing noticeably contracted predictions.
-This suggests that their temporal representations favor local smoothing and are unable to capture the long–range coordinated forces driving the sustained outward advancement of the expansion front and its slow, anisotropic curvature modulation.
-The SimVP model employing the gSTA temporal module better approximates the global scale, yet its predicted boundaries remain irregular and exhibit gradual drift over time, indicating limited geometric stability once high–frequency edge cues have diminished.
-The SimVP model using the TAU unit performs intermediately, but displays inconsistent boundary curvature and localized radial deviations across extended prediction intervals, reflecting difficulties in balancing temporal smoothness with geometric precision.
-In contrast, Morpher maintains both the disc–like radial organization and the fine–scale curvature variations throughout this stabilized growth phase.
-Its predicted boundaries closely follow the true expansion front without contraction, drift, or distortion, demonstrating that Morpher preserves stable global organization even when local structural cues weaken.
-Together with the observations from the dense–finger–dominated regime, these results indicate that Morpher captures the underlying evolutionary dynamics of swarming colonies rather than relying on short–term motion extrapolation or local texture continuity.
-The limited performance of conventional video prediction models in this task stems from their emphasis on global pixel–level variation rather than on the dynamical organization of the expanding morphological front.
-Morpher departs from this paradigm by operating directly on the binary mask sequences produced by TexPol–Net, enabling temporal modeling to take place within a compact latent representation that encodes structural configuration rather than raw appearance.
-Within this latent space, Morphon works together with the autoregressive feedback loop to continuously update the representation based on the accumulated history of colony growth.
-This mechanism allows Morpher to directly capture the speed, directional bias, and spatial coordination of front propagation.
-Coupled with a multi–scale decoder that preserves continuity between successive morphological states, the model maintains coherent global organization and smooth boundary evolution even over long forecast horizons.
-As a result, Morpher produces long–range predictions that exhibit higher dynamical consistency and more faithful morphological progression than existing video prediction approaches.
-2.4 Small improvements in front recovery yield large gains in forecasting
-While the predictive performance of Morpher demonstrates its ability to capture temporal dependencies, the reliability of long–term forecasts hinges on the accuracy of the input representation—the segmentation masks delineating colony morphology.
-To assess this relationship, we compared Morpher coupled with our TexPol–Net segmentation backbone to the same forecasting framework driven by YOLOv12.
-YOLOv12 achieves the second–highest segmentation accuracy on our swarming colony dataset—second only to TexPol–Net—and thus serves as the strongest alternative baseline for this task.
-This comparison provides a clear and controlled evaluation of how the segmentation backbone affects subsequent temporal modeling.
-Both systems follow the identical two–stage pipeline illustrated in Fig.
-1, using the same Morpher architecture, training settings, and evaluation criteria; the only difference lies in the source of the input segmentation masks.
-Although YOLOv12 (91.81 mAP50:95) trails TexPol–Net (92.48 mAP50:95) by less than one percentage point in static segmentation accuracy, this seemingly minor gap translates into a disproportionately large effect on temporal forecasting.
-As illustrated in Fig.
-2.3c, the slight boundary inaccuracies present in YOLOv12 masks accumulate through Morpher’s autoregressive updates, leading to a gradual drift of the colony front and a measurable loss of curvature fidelity over time—up to a 2.4–3.1 IoU-point gap after 20% forecasting.
-Forecasts initialized from YOLOv12 masks exhibit smoothed protrusions and reduced dense–finger morphology, consistent with progressive attenuation of high–frequency boundary features.
-In contrast, TexPol–Net masks preserve sharper front geometry and coherent internal texture cues, allowing Morpher to maintain stable propagation of the expanding rim with minimal geometric distortion.
-Temporal coherence in morphological forecasting depends not only on the expressiveness of the temporal model but also on the structural fidelity of its spatial inputs.
-In swarming dynamics—where front curvature, anisotropy, and the organization of dense fingers jointly define the expansion process—even slight degradations in segmentation accuracy can disrupt the causal progression linking successive morphological states.
-TexPol–Net mitigates this sensitivity through its TEA and PCA mechanisms, which together sharpen fine–scale front delineation while preserving the global radial organization of the colony.
-These inductive priors yield geometrically consistent masks that capture the underlying biophysical architecture with high fidelity, enabling Morpher to propagate colony morphology smoothly and without structural drift.
-From a modeling standpoint, the results reveal that even a seemingly marginal one–percentage–point improvement in segmentation accuracy can induce a qualitative shift in system behavior—from cumulative error amplification to dynamical stability.
-The integration of TexPol–Net and Morpher thus forms a synergistic framework that unites spatial precision with temporal reasoning, enabling quantitative reconstruction of curvature–driven, long–range evolution in swarming fronts and, more broadly, extending predictive modeling to other forms of collective biological motion.
-2.5 Temporal models and inference strategies for front forecasting
-We used Morpher to systematically probe how temporal sequence models, inference strategies, and the Morphon module influence long–horizon forecasts of swarming colony morphology, as summarized in Fig.
-2.5b.
-Four architectures—a simple RNN, GRU, LSTM, and Transformer encoder—were evaluated under parallel versus autoregressive decoding, with or without Morphon.
-Forecasting quality was assessed using a set of complementary metrics capturing geometric accuracy and dynamical stability, including region–level overlap, boundary deviation, front–propagation dynamics, temporal consistency, and angular organization.
-The quantitative results underlying these comparisons are reported in detail in Table 2, which provides the full ablation statistics for each architectural and inference configuration.
-A first observation is that autoregressive decoding consistently produced more faithful colony evolution than parallel one–shot decoding.
-Across architectures, switching from parallel to autoregressive inference improved either mIoU, boundary metrics, or both.
-For example, the Transformer with Morphon achieved 94.80% mIoU and an ASSD of 4.63 px in parallel mode, compared with 95.42% mIoU and 3.93 px ASSD under autoregressive decoding.
-Similar patterns were observed for GRU and LSTM, where autoregression reduced HD95 and ASSD and kept RMSE at low values.
-This behavior matches the biology of the system: swarming expansion proceeds through many small, incremental deformations of the front, and each future shape is tightly constrained by the immediately preceding configuration.
-Parallel decoding enforces all future frames in a single pass, which tends to collapse the diversity of plausible growth trajectories into a smoother, more symmetric contour and suppresses local protrusions.
-Conditioning each prediction on the previous frame allows the model to adjust the advancing front step by step, preserving fingering, local curvature, and realistic front propagation over long horizons.
+a Performance under an 80% observation / 20% prediction protocol.
+Morpher achieves the best overall accuracy (95.42% mIoU, 10.61 px HD95, 3.93 px ASSD), indicating improved front localization and boundary fidelity.
+b Long-horizon forecasts on two representative sequences.
+Generic predictors smooth finger-like branches or accumulate boundary drift, whereas Morpher maintains a coherent front closer to the ground truth across regimes.
+c Effect of segmentation backbone on forecasting.
+TexPol–Net masks yield more accurate final-frame predictions than YOLOv12 [58], increasing IoU from 94.21% to 96.63% in the anisotropic branching regime and from 93.30% to 96.38% in the near-concentric regime.
+Morpher performs best on all three metrics, reaching 95.42% mIoU, 10.61 px HD95, and 3.93 px ASSD.
+Relative to the strongest baseline, SimVP+gSTA, this corresponds to a 5.4% gain in overlap and reductions of 42.0% and 55.7% in HD95 and ASSD.
+These gains are obtained under an 80% observation / 20% prediction protocol, where prediction is restricted to the late expansion stage, when the front is already extended and highly sensitive to small geometric deviations.
+The qualitative failures follow the same ordering (Figure 4b).
+In the anisotropic branching regime, MIM progressively smooths protrusions; SimVP variants retain a coarse outline but truncate lobe tips and blunt high-curvature sectors; MAU, PredRNN, and PredRNNv2 lag in propagation, yielding fronts that are spatially plausible but temporally delayed.
+In the late near-concentric regime, the same models underestimate radial extent or accumulate boundary drift.
+Morpher remains closest to the true front across both regimes, preserving local perturbations without losing global scale.
+The separation reflects a mismatch in what is being modeled: generic video predictors favor appearance continuity, whereas Morpher treats morphology itself as the evolving state.
+2.5 Segmentation errors accumulate during forecasting
+A causal link between measurement fidelity and forecast stability can be tested by a controlled swap: holding the temporal model fixed while changing only the segmentation backbone that defines the state.
+We therefore paired the same Morpher architecture with either TexPol–Net or YOLOv12.
+The static difference is small: TexPol–Net achieves 92.48% mAP50:95, whereas YOLOv12 reaches 91.81%.
+The temporal consequence is not.
+Over the final 20% prediction window, YOLOv12 masks incur a 2.4–3.1 IoU-point loss in final-frame prediction.
+This amplification is visible in both representative regimes (Figure 4c).
+In the anisotropic branching sequence, final-frame IoU increases from 94.21% to 96.63%; in the near-concentric sequence, from 93.30% to 96.38%.
+This establishes a causal amplification mechanism: sub-percent-level segmentation differences are sufficient to induce multi-point degradation in long-horizon prediction.
+The mechanism is intrinsic to autoregressive boundary evolution, in which small geometric biases are recursively fed back as state and thereby converted into cumulative drift.
+2.6 Temporal modeling of front dynamics
+Temporal prediction must preserve two coupled properties: where the boundary lands (geometric fidelity) and how it gets there (dynamical consistency).
+The evaluation therefore spans eight complementary metrics (Figure 5b), with a deliberate hierarchy for interpretation: boundary-sensitive distances (HD, HD95, ASSD) and overlap (mIoU) report localization; propagation RMSE and TCI test trajectory-level stability; angular deviations ($|\Delta\mathrm{NAS}|$ and $|\Delta\mathrm{H}_{2}|$) diagnose whether anisotropic branching is preserved beyond coarse extent.
 Figure 5: Morpher enables geometry-consistent long-horizon forecasting of swarming colony morphology.
-a, Architecture of Morpher.
-Observed segmentation masks are mapped to a compact morphological latent sequence by a multi-scale spatial encoder.
-Future evolution is predicted autoregressively in latent space, with each decoded morphology fed back to condition subsequent steps.
-The Morphon module selectively retrieves informative past states via cross-attention and integrates them through a learnable gate, while a multi-scale decoder reconstructs spatial detail and preserves peripheral protrusions and front curvature.
-b, Systematic evaluation of temporal architectures and inference strategies under a fixed 80% observation / 20% prediction protocol.
-Across eight complementary metrics capturing geometric accuracy and dynamical stability, autoregressive inference tends to outperform parallel decoding, and the inclusion of Morphon improves boundary fidelity and temporal consistency across sequence models (RNN, LSTM, GRU and Transformer).
-c, Robustness of morphological forecasting to observation length.
-For autoregressive models with Morphon enabled, prediction accuracy and stability improve smoothly as the observation ratio increases from 50% to 90%, indicating that Morpher exploits additional temporal context without overfitting to a specific observation–prediction split.
-The sequence models themselves formed a clear performance hierarchy that aligns with their representational capacity.
-The plain RNN was consistently weakest on geometry-sensitive metrics.
-Without Morphon in parallel mode, it reached 93.23% mIoU but suffered from large boundary errors (ASSD 6.02 px, $|\Delta\mathrm{NAS}|$ 19.54%), yielding overly round and contracted colonies.
-Even with autoregression and Morphon, ASSD remained above 4.3 px and anisotropy above 16%, indicating that local perturbations at the front were rapidly forgotten and the contour relaxed toward a smooth disk.
-Introducing gating improved this behavior.
-The GRU in parallel mode without Morphon already reduced ASSD to 5.17 px and $|\Delta\mathrm{NAS}|$ to 15.31%, and in its best configuration (autoregressive with Morphon) it further lowered ASSD to 4.06 px and achieved the smallest RMSE (2.06 px per frame) among all models, indicating particularly sharp tracking of front advancement and rapid local changes.
-The LSTM occupied an intermediate regime: it produced lower HD and HD95 than the RNN and stable RMSE values (2.26–2.84 px per frame), capturing global expansion reliably but reacting less strongly to fine, transient undulations of the front.
-The Transformer–based Morpher provided the most balanced overall forecasts.
-In the autoregressive Morphon setting, it achieved the highest mIoU (95.42%), the lowest ASSD (3.93 px), and the smallest anisotropy deviation ($|\Delta\mathrm{NAS}|=13.13\%$), while keeping both the RMSE of front propagation and $|\Delta\mathrm{H}_{2}|$ at low values.
-These scores indicate that the model not only captured the colony’s radial expansion and average outline, but also maintained fingering, sectoring, and higher–order harmonic structure.
-Self–attention over the encoded history enables the Transformer to couple distant parts of the front and integrate long–range constraints, which is particularly important for reproducing radial symmetry–breaking and coordinated fingering across the colony.
-Morphon provided an additional axis of improvement by encoding a persistent structural memory of past fronts.
-For every architecture and inference strategy, enabling Morphon increased mIoU and sharpened the boundary.
-For instance, in parallel mode the RNN improved from 93.23% to 94.22% mIoU and reduced ASSD from 6.02 px to 4.85 px when Morphon was switched on.
-Autoregressive GRU improved HD95 from 12.03 px without Morphon to 10.14 px with Morphon, and the Transformer under autoregression decreased ASSD from 5.26 px to 3.93 px when Morphon was enabled.
-Angular metrics showed the same tendency: in the strong Transformer configuration, $|\Delta\mathrm{NAS}|$ decreased from 16.83% without Morphon in parallel mode to 13.13% with Morphon in autoregressive mode, reflecting the combined effect of autoregression and Morphon in reinforcing angular fidelity.
-Across models, $|\Delta\mathrm{H}_{2}|$ remained below 2.0%, indicating that second–harmonic structure was generally well preserved, while Morphon primarily helped to maintain the finer anisotropic patterning of the front.
-Accurate forecasting depends on both an appropriate inference paradigm and biologically informed memory.
-Autoregressive decoding is crucial for maintaining causal continuity of growth, while Morphon stabilizes curvature and anisotropy by carrying forward structural cues.
-Within this design, the Transformer–based Morpher offers the best compromise between boundary fidelity, propagation dynamics, and angular organization.
-The GRU emphasizes rapid front dynamics, the LSTM ensures smooth global expansion at the cost of some fine structure, and the plain RNN underestimates both fingering and symmetry breaking.
-These patterns support a mechanistic interpretation: models that can remember and attend to past colony structure are best suited to predict how microbial collectives remodel themselves in time.
-2.6 Forecasting stability across observation ratios
-We next examined how Morpher behaves as the amount of observed history is systematically varied.
-For each temporal backbone operating in its best–performing configuration (autoregressive decoding with Morphon), the observation fraction was increased from 50% to 90%.
-Model performance was then evaluated using the same eight metrics described above.
-The resulting trends are summarized in Fig.
-2.5c, with the corresponding quantitative values reported in Table 3.
-This analysis tests whether Morpher is tuned to a narrow observation–prediction split or instead maintains stable forecasting behavior across a broad range of experimental protocols.
-Geometry-based metrics showed a smooth and largely monotonic improvement as more of the past was revealed.
-For all four sequence models, mIoU increased from the high–80% range at 50% observation to above 96% at 90%, while HD95 and ASSD decreased correspondingly.
-The Transformer, for example, improved from 88.22% to 96.79% mIoU and reduced ASSD from 9.49 px to 2.75 px as the observation ratio increased from 50% to 90%.
-The relative ranking of sequence models remained essentially unchanged: at each observation level the Transformer offered the best compromise between overlap and boundary error, GRU and LSTM occupied an intermediate regime, and the plain RNN lagged slightly behind.
-This preservation of ordering indicates that the choice of temporal architecture is more important than the precise observation ratio for recovering the colony outline and front position.
-Front–propagation dynamics, measured by RMSE of radial expansion rate, exhibited a similar saturation pattern.
-The largest gains occurred between 50% and 60% observation, after which RMSE values fluctuated within a narrow band of roughly 2.0–2.3 px per frame across models.
-In other words, once the model has seen slightly more than half of the trajectory, the incremental benefit of additional frames for predicting front–propagation speed is modest.
-TCI behaved in a comparably stable way but with slightly larger relative variation.
-TCI values clustered around 63–65% from 50% to 80% observation, with the LSTM reaching the highest TCI (65.32%) at 80%.
-At 90% observation the prediction window becomes too short for a meaningful fluctuation index, which is why TCI is not reported there.
-From a mathematical standpoint, both RMSE and TCI are based on temporal differences rather than static shapes; they are sensitive to small phase shifts in the timing of front accelerations and pauses, even when the underlying boundary is well aligned.
-Angular metrics, particularly $|\Delta\mathrm{NAS}|$, varied more across observation ratios and did not follow a simple monotonic trend.
-For the Transformer, $|\Delta\mathrm{NAS}|$ decreased from 14.62% at 50% observation to 11.93% at 60%, then increased to 15.57% at 70% before dropping again to 13.13% at 80% and settling around 14% at 90%.
-The recurrent models showed the same qualitative behaviour, with a pronounced peak near 70–80%.
-This seemingly counterintuitive pattern—more observed history not always yielding smaller anisotropy error—arises from how NAS is defined: it aggregates angular contrasts over the entire prediction horizon.
-Increasing the observation ratio shortens this horizon and shifts it toward later times, where the colony has already developed complex sectoring and branch splitting.
-In this regime, even small local disagreements in when a branch emerges or merges can produce relatively large changes in the angular power spectrum.
-Mathematically, $|\Delta\mathrm{NAS}|$ behaves more like a measure of angular fluctuation intensity than a simple bias; it is highly sensitive to local phase errors of individual fingers, and therefore displays a non–monotonic dependence on the observation window.
-The second–harmonic descriptor $|\Delta\mathrm{H}_{2}|$ showed a different but equally non–trivial trend.
-For the three recurrent architectures, $|\Delta\mathrm{H}_{2}|$ increased almost monotonically from $\sim\!1.1$–1.2% at 50% observation to around 2.2–2.3% at 90%.
-The Transformer followed the same overall tendency, with values around 1.18–1.35% at 50–60% observation and $\sim\!2.2\%$ at 90%, albeit with a local maximum at 70% and a mild dip at 80%.
-At first glance, this appears to contradict the intuition that more context should always improve harmonic structure.
-However, $|\Delta\mathrm{H}_{2}|$ isolates only the second angular mode of the front.
-As the colony matures, curvature energy is progressively redistributed into higher–order modes through branch splitting and tip instabilities.
-The prediction window at high observation ratios is therefore dominated by subtle rearrangements of fine–scale protrusions rather than by the initial quadrupolar shape.
-In this late regime, tiny mismatches in the amplitude or phase of low–frequency modes can produce larger absolute differences in the second harmonic, even when the full boundary and overall anisotropy are well captured.
-From a mathematical perspective, the metric is probing how accurately the model tracks a single Fourier component in a spectrum whose energy is shifting across modes; as the spectrum becomes more crowded, that task becomes harder rather than easier.
-Morpher remains stable across observation ratios for coarse geometric descriptors and front–propagation speed, whereas the more sensitive temporal and angular metrics reveal where the remaining difficulty lies.
-Once roughly 60–70% of the trajectory has been observed, mIoU, HD–based distances, ASSD, and RMSE change only gradually with further increases in observation, and the relative ranking of sequence models remains fixed.
-In contrast, $|\Delta\mathrm{NAS}|$ and $|\Delta\mathrm{H}_{2}|$ respond strongly to how the forecasting window slices the late–time dynamics of sectoring and finger rearrangement, whereas TCI shows modest but noticeable sensitivity.
-This pattern indicates that changing the observation–prediction split does not alter the relative ranking of the sequence models in recovering global expansion, and that the main residual challenges are concentrated in the most fluctuation– and angle–sensitive aspects of the swarming front.
+a Morpher architecture.
+Observed masks are encoded into a compact morphological latent sequence, and future evolution is predicted autoregressively with decoded states recursively conditioning subsequent steps.
+The Morphon module retrieves past states via cross-attention and integrates them through a learnable gate, while a multi-scale decoder preserves boundary detail and front geometry.
+b Evaluation of temporal architectures and inference strategies under an 80% observation / 20% prediction protocol.
+Autoregressive inference consistently outperforms parallel decoding, and Morphon improves boundary fidelity and temporal consistency across RNN, LSTM, GRU, and Transformer models.
+c Robustness to observation length.
+With Morphon enabled, prediction accuracy and stability improve as the observation ratio increases from 50% to 90%, indicating effective use of temporal context without overfitting.
+We analyzed temporal architectures (RNN, GRU, LSTM, and Transformer) under parallel and autoregressive decoding, with or without Morphon (Figure 5b).
+Autoregressive decoding consistently outperforms parallel prediction.
+For Transformer with Morphon, performance improves from 94.80% mIoU and 4.63 px ASSD in parallel mode to 95.42% and 3.93 px under autoregressive decoding.
+GRU and LSTM show the same directional improvement.
+Stepwise state updates preserve local geometric continuity, whereas one-shot prediction smooths the front.
+The backbones nonetheless separate along different axes of performance.
+The plain RNN is weakest: without Morphon in parallel mode, it reaches 93.23% mIoU, 6.02 px ASSD, and $|\Delta\mathrm{NAS}|=19.54\%$, producing overly round and contracted fronts.
+GRU is substantially stronger; even without Morphon in parallel mode, it reduces ASSD to 5.17 px and $|\Delta\mathrm{NAS}|$ to 15.31%, and in its best configuration it further lowers ASSD to 4.06 px and achieves the lowest propagation RMSE, 2.06 px per frame.
+LSTM is more stable than RNN and keeps RMSE between 2.26 and 2.84 px per frame, but reacts less strongly than GRU to fine front undulations.
+Critically, the model that minimizes propagation error is not the one that best preserves morphology.
+GRU achieves the lowest RMSE (2.06 px per frame).
+Under the same setting, it remains worse than Transformer in boundary fidelity and anisotropic structure preservation.
+In the autoregressive Morphon setting, Transformer yields the highest mIoU (95.42%), the lowest ASSD (3.93 px), and the smallest anisotropy deviation ($|\Delta\mathrm{NAS}|=13.13\%$), while keeping both propagation RMSE and $|\Delta\mathrm{H}_{2}|$ low.
+Accurate front advance and faithful morphology reconstruction are therefore not identical objectives.
+The former is largely a gated-memory problem, whereas the latter requires long-range retrieval of branch history and spatial configuration.
+Morphon improves every backbone, and its effect is diagnostic rather than cosmetic: it specifically targets the non-Markovian part of boundary evolution, where similar instantaneous curvatures can lead to different futures depending on earlier branch history.
+In parallel RNN, Morphon raises mIoU from 93.23% to 94.22% and reduces ASSD from 6.02 to 4.85 px.
+In autoregressive GRU, it reduces HD95 from 12.03 to 10.14 px.
+In Transformer, it lowers ASSD from 5.26 to 3.93 px and decreases $|\Delta\mathrm{NAS}|$ from 16.83% to 13.13%.
+Notably, $|\Delta\mathrm{H}_{2}|$ remains below 2.0% across models, suggesting that low-order harmonic structure is largely preserved even by weaker predictors.
+The main separation therefore lies not in coarse global shape, but in the preservation of finer anisotropic organization along the advancing front.
+2.7 Scaling with observation length
+The observation ratio was varied from 50% to 90% for each backbone in its best configuration, namely autoregressive decoding with Morphon (Figure 5c).
+We begin at 50% observation because the early phase is morphologically convergent across colonies (Figure 6d–f), and prediction only becomes well-posed once colony-specific geometry begins to diverge.
+Geometry-based metrics improve smoothly with longer history.
+For the Transformer, mIoU rises from 88.22% at 50% observation to 96.79% at 90%, while ASSD falls from 9.49 to 2.75 px.
+The relative ordering of the backbones remains unchanged across the full range, indicating that temporal architecture matters more than the exact observation–prediction split.
+Front-propagation RMSE follows the same direction but saturates earlier.
+The largest gain appears between 50% and 60% observation; beyond that, values remain within a narrow 2.0–2.3 px per frame band.
+TCI is similarly stable, clustering around 63–65% from 50% to 80% observation, with the highest value observed for LSTM at 80% observation (65.32%).
+At 90%, the prediction window is too short for a stable TCI estimate.
+Angular metrics expose a stricter regime and reveal a second constraint: more history does not guarantee better anisotropy preservation because the error is dominated by phase alignment of late-stage branch rearrangements.
+For the Transformer, $|\Delta\mathrm{NAS}|$ changes from 14.62% at 50% observation to 11.93% at 60%, rises to 15.57% at 70%, drops to 13.13% at 80%, and remains around 14% at 90%.
+$|\Delta\mathrm{H}_{2}|$ shows a different but equally late-stage-sensitive trend: for the Transformer it is 1.18–1.35% at 50–60% observation and approaches $\sim$2.2% at 90%.
+Additional history therefore yields diminishing returns for low-order shape statistics, while leaving phase-sensitive alignment of late front rearrangements as the dominant residual difficulty.
+2.8 Generalization of trajectory-level dynamics
+Out-of-sample stability was tested by leave-one-out cross-validation, evaluating each trajectory with models trained on all remaining trajectories (Figure 6).
+Performance approaches saturation as the training set expands, with only marginal gains once most trajectories are included (Figure 6a), indicating that the dominant modes of swarming variation relevant for forecasting are already well represented.
+Figure 6: Generalization, accuracy, and dynamical consistency of swarming morphology forecasting under data-limited conditions.
+Results are evaluated using leave-one-out cross-validation across 81 independent colonies.
+a Effect of training set size on forecasting performance (80% observation / 20% prediction).
+Metrics converge with increasing data, indicating that the observed variability is sufficiently captured.
+b Static forecasting accuracy (80% / 20%) measured by region- and boundary-based metrics.
+c Dynamical consistency (80% / 20%) relating propagation error, anisotropy deviation, and normalized performance.
+d–f Time-aligned trajectories of effective radius, perimeter, and front propagation velocity under a 50% observation / 50% prediction protocol, with individual trajectories and mean trends.
+g Prediction-horizon–aligned trajectories (50% / 50%) showing forecast evolution within the prediction window.
+h Prediction-horizon–aligned trajectories (80% / 20%) showing consistent behavior across observation regimes.
+Leave-one-out evaluation separates static accuracy from dynamical fidelity.
+Region- and boundary-based metrics remain concentrated in a high-performance regime (Figure 6b), whereas dynamical measures separate propagation error, anisotropy deviation, and harmonic distortion more clearly (Figure 6c).
+The model therefore generalizes well in mask overlap without collapsing the more sensitive dynamical dimensions into the same signal.
+This separation is preserved at the trajectory level.
+Under a 50% observation / 50% prediction protocol, predicted effective radius and perimeter closely track the true time-aligned trajectories, and front propagation velocity reproduces the transition from rapid early expansion to later stabilization (Figure 6d–f).
+Agreement remains after alignment by prediction horizon (Figure 6g).
+Under an 80% observation / 20% prediction split, the same dynamical organization is retained (Figure 6h): longer observation tightens prediction over a shorter horizon without changing the overall structure of the trajectories.
+Across settings, Morpher preserves not only colony extent but also the temporal organization of expansion.
 3 Discussion
-Swarming colonies are often cited as emblematic examples of emergent collective behaviour, yet most analyses remain anchored in static descriptions of highly structured fronts.
-Here we recast the same imaging data into a predictive formulation of swarming dynamics.
-Rather than treating expansion as a sequence of images, the colony edge is modelled as a dynamical object evolving within a low–dimensional morphological state space.
-By assembling the SwarmEvo dataset, introducing TexPol–Net for boundary measurement, and developing Morpher as a morphology–level dynamical model with structural memory, the expansion of Enterobacter sp.
-SM3 becomes a forecasting problem defined at the level of geometry, rather than a descriptive account of shape.
-This reframing rests on an explicit separation between measurement, state construction and dynamics.
-Time–lapse videos are first reduced to sequences of boundary–resolved masks, defining the measured observables of the system.
-TexPol–Net operates at this stage not simply as a segmentation network, but as a boundary measurement operator that stabilizes the morphological degrees of freedom of the colony front.
-By suppressing diffuse edges, preserving thin protrusions and decoupling the front from background artefacts, TexPol–Net defines a consistent coordinate system in which small geometric variations retain temporal coherence.
-Improvements in boundary accuracy therefore propagate nonlinearly into forecasting performance, as every measurement error reshapes the subsequent trajectory in state space.
-In this sense, TexPol–Net establishes the state variables on which dynamics can be meaningfully learned, rather than merely refining per–frame masks.
-This sensitivity becomes most apparent when segmentation backbones are required to resolve near–equivalent masks.
-On SwarmEvo, YOLOv12 differs from TexPol–Net by less than one mAP50:95 point (91.81 vs.
-92.48; Fig.
-3c), yet this minor discrepancy at the measurement stage expands once the masks define the state variables for autoregressive forecasting.
-When Morpher is held fixed, boundary deviations in YOLOv12 inputs, including slight edge displacement and reduced high–curvature detail, propagate through successive rollouts, producing gradual front drift and a measurable loss of curvature that reaches a 2.4–3.1 IoU gap after 20% prediction (Fig.
-2.3c).
-TexPol–Net masks, by contrast, sustain consistent front localization and curvature over extended horizons, limiting geometric degradation during temporal propagation (Fig.
-3b).
-SAM/SAM2 accentuate this contrast by suppressing distal structure and biasing contours toward the colony core, effects that are already evident at the segmentation stage (Fig.
-3).
-These comparisons indicate that once the colony edge is treated as the dynamical state, boundary fidelity directly governs whether temporal prediction remains stable or undergoes progressive error growth.
-Morpher operates on this measurement–defined state space to model the evolution of the front.
-Unlike generic video prediction architectures, Morpher is constrained to propagate contours and to encode dynamics in terms of geometric quantities such as curvature and radial displacement (Fig.
-2.5).
-This design choice places Morpher explicitly in the role of a dynamical model, rather than an image generator.
-Benchmarking against state–of–the–art sequence models shows that enforcing this geometric alignment improves both region–level overlap and boundary fidelity, while preserving anisotropic branching despite the absence of texture information.
-Ablations across backbones, inference strategies and the Morphon module reinforce a mechanistic interpretation of the model behaviour (Fig.
-2.5b).
-Autoregressive decoding consistently yields more realistic front propagation than parallel one–shot prediction, reflecting the incremental nature of physical growth.
-Architectures equipped with attention and persistent structural memory, particularly the Transformer with Morphon, best preserve fingering and radial symmetry breaking, indicating that long–range couplings are integral to swarming dynamics.
-Varying the observation ratio provides further insight into the stability of morphology–level forecasting (Fig.
-2.5c).
-As the observed fraction increases from 50% to 90%, coarse geometric metrics such as mIoU, HD–based distances and ASSD improve and then saturate.
-Once slightly more than half of the trajectory is observed, additional frames contribute little to recovering the global outline or average advance speed.
-In contrast, temporal and angular descriptors reveal where prediction remains sensitive.
-TCI exhibits measurable dependence on the forecasting window, indicating that small phase shifts in accelerations and pauses perturb fluctuation profiles even when boundaries remain well aligned.
-Angular metrics expose a more stringent regime.
-The non–monotonic dependence of $|\Delta\mathrm{NAS}|$ on observation ratio suggests that anisotropy errors are governed less by the amount of historical data than by how the prediction window intersects late–stage sectoring and branch rearrangements.
-In this regime, minor mismatches in timing or amplitude reshape the angular power spectrum even when global geometry is preserved.
-The behaviour of $|\Delta\mathrm{H}_{2}|$ sharpens this interpretation.
-Isolating the second harmonic renders the metric sensitive to the redistribution of curvature energy across Fourier modes.
-At high observation ratios, prediction targets are dominated by fine protrusions and tip instabilities, where small deviations in low–frequency content amplify into substantial harmonic differences.
-These results indicate that anisotropic fingering and harmonic organization represent the most demanding aspects of swarming to forecast, and they suggest a fundamental sensitivity of morphology–only dynamics to small temporal errors.
-Several limitations delineate avenues for extension.
-SwarmEvo focuses on a single strain, Enterobacter sp.
-SM3.
-Although its behaviour spans multiple regimes, swarming encompasses a broad spectrum of species and substrates whose statistics may differ substantially.
-The present representation further collapses colonies into quasi–two–dimensional fronts.
-While such projections preserve the dominant geometric degrees of freedom and are standard in swarming studies, height and density variations enter only implicitly.
-Morpher is trained without explicit biophysical constraints; the absence of hand–crafted rules affords flexibility, yet extrapolation into sparsely sampled regimes may benefit from mechanistic priors.
-Finally, forecasting is treated as an open–loop problem.
-Closed–loop experiments in which predictions guide perturbations would provide a natural next step toward control.
-Beyond swarming, this work establishes a general methodology for forecasting dynamics in systems where appearance is uninformative but geometry is structured.
-By formalizing expansion as a measurement–defined dynamical system—where boundary measurement defines the state space and forecasting reduces to trajectory evolution rather than image prediction—the framework clarifies how segmentation, state construction and temporal inference interact.
-Representing videos as trajectories in a geometry-aligned morphological state space enables principled comparison of sequence models, separates measurement error from dynamical uncertainty, and supports interpretable metrics linked directly to branching, anisotropy and harmonic organization.
-These principles extend to other forms of collective motility, biofilm spreading and tissue invasion, where dynamics unfold through shape rather than texture.
-By enabling long–horizon forecasts and identifying features that are intrinsically sensitive to timing, this framework transforms colony expansion into a quantitative, methodologically grounded basis for intervention, providing a foundation for predictive modelling and control of microbial collectives in complex environments.
+Swarming is a striking example of emergent collective behavior [66, 25], yet its advancing front remains treated as a descriptive pattern rather than a predictive dynamical state.
+Here we recast Enterobacter sp.
+SM3 expansion as a predictive problem defined at the colony edge.
+This shift is biologically consequential: in inflamed intestinal environments, the advancing front is the interface through which swarming reshapes oxygen availability, niche structure, and downstream community organization (Figure 1[14].
+The question is therefore not simply how a colony looks, but whether the future position and organization of its front can be inferred from its present state.
+A central result is that this future is not cleanly indexed by nominal assay conditions.
+Trajectories remain broadly overlapping across temperature, humidity, and agar groupings (Figure 2a,b), and within- and across-condition distances are only weakly separated (Figure 2c,d).
+Prediction therefore depends on colony-specific morphology rather than on coarse labels.
+This helps explain why appearance-based extrapolation fails: the relevant information is carried by front geometry, protrusions, and temporal continuity, not by condition identity or framewise visual similarity [65].
+That observation makes measurement decisive.
+TexPol–Net matters not simply because it improves segmentation, but because it defines the state on which dynamics are learned.
+Once the colony edge is treated as the evolving variable, small boundary errors are no longer cosmetic (Figure 3c–e).
+A sub-point difference in segmentation quality between TexPol–Net and YOLOv12 (92.48% vs.
+91.81% mAP50:95) expands into a 2.4–3.1 IoU loss after autoregressive rollout (Figure 3c and Figure 4c).
+Stable forecasting therefore requires boundary fidelity at the measurement stage.
+Within that state space, Morpher behaves, as schematized in Figure 5a, as a front-dynamics model rather than an image predictor.
+Its advantage over generic video architectures lies not only in overlap, but in preserving front localization and anisotropic branching over long horizons (Figure 4a,b).
+The temporal ablations sharpen this point.
+Autoregressive decoding consistently outperforms parallel prediction, indicating that swarming is better modeled as incremental state propagation than as one-shot future synthesis (Figure 5b).
+More importantly, the model that minimizes propagation error is not the one that best preserves morphology: GRU achieves the lowest RMSE, whereas Transformer with Morphon best maintains boundary fidelity and anisotropic structure (Figure 5b).
+The harder part of swarming prediction is therefore not coarse advance alone, but retention of branch history and late-stage front organization [67].
+The observation-ratio and leave-one-out analyses place that difficulty more precisely.
+Global geometric metrics improve and then saturate with longer observation, whereas angular descriptors remain sensitive to late rearrangements of the front (Figure 5c).
+At the same time, leave-one-out evaluation shows that the model preserves not only mask overlap but also the temporal organization of effective radius, perimeter, and propagation velocity across held-out trajectories (Figure 6b,d–h).
+Together, these results indicate that the dominant residual challenge lies in anisotropic branching and phase-sensitive reorganization, not in recovery of coarse colony extent.
+Several limitations define the next steps.
+While leave-one-out evaluation already indicates that the model captures transferable dynamics under limited data, SwarmEvo is currently restricted to a single strain, Enterobacter sp.
+SM3, with broader taxonomic and substrate diversity remaining as a natural extension to assess the generality of these dynamics.
+The present representation is quasi-two-dimensional, so height and density enter only indirectly.
+Morpher also operates without explicit biophysical constraints; incorporating mechanistic priors may further improve extrapolation in sparsely sampled regimes.
+Finally, forecasting remains open-loop.
+Closed-loop perturbation experiments guided by model predictions would provide the most direct test of whether predictive morphology can be translated into controllable intervention at the swarming front.
+More broadly, this framework suggests a route for living systems in which shape carries the essential dynamics [19].
+By separating measurement, state construction, and temporal evolution, it becomes possible to ask which aspects of collective behavior are predictable, which remain unstable, and which are most sensitive to perturbation.
+For swarming, this turns morphology from a descriptive readout into a dynamical variable that can be compared, forecasted, and ultimately acted upon.
+4 Conclusions
+Boundary-resolved morphology provides a workable state space for forecasting Enterobacter sp.
+SM3 swarming.
+SwarmEvo supplies the time-lapse trajectories, TexPol–Net recovers fronts with sufficient fidelity to define the predictive state, and Morpher advances those states through autoregressive rollout with structural memory.
+Under the 80% observation / 20% prediction protocol, Morpher achieves the best overall performance (mIoU 95.42%, HD95 10.61 px, ASSD 3.93 px), outperforming the strongest baseline by 5.4% in overlap and reducing HD95 and ASSD by 42.0% and 55.7%, respectively.
+Just as importantly, the study shows why forecasting is feasible only in a geometry-first formulation: once morphology is measured accurately, swarming expansion can be treated as a predictable dynamical process rather than an appearance-driven one.
+This establishes a basis for quantitative comparison of swarming trajectories across environments and for future predictive control of advancing microbial fronts.
 5 Methods
-4.1 Experimental workflow, imaging, and mask generation
-The experimental workflow followed Fig.
-1.
+5.1 Swarming assay, imaging, and dataset construction
+The experimental workflow followed Figure 1.
 A single colony of Enterobacter sp.
-SM3 was transferred from an LB agar plate into LB broth (10 g/L tryptone, 5 g/L yeast extract, and 5 g/L NaCl) and incubated at $37\,^{\circ}\mathrm{C}$ with shaking at 200 rpm for approximately 16 h to obtain a high-density culture.
-A $5\text{--}8\,\mu\mathrm{L}$ aliquot of the overnight culture was spotted onto the center of a freshly prepared swarming agar plate (LB medium supplemented with 0.5–0.8% agar), poured to a uniform thickness of 3–4 mm to promote consistent hydration and nutrient diffusion across plates.
-The inoculated plates were incubated at $30\,^{\circ}\mathrm{C}$ and approximately 90% relative humidity for 4–6 h to activate swarming motility, and then transferred to a time-lapse imaging chamber maintained between $28\text{--}37\,^{\circ}\mathrm{C}$ and 80–92% relative humidity.
-Humidity was kept below the condensation threshold to prevent droplet formation on the agar surface.
-Surface stiffness and inoculum volume were varied across experiments to introduce controlled physical differences and generate diverse colony expansion conditions.
-Imaging was performed using a vertically mounted high-resolution digital camera with uniform LED illumination to minimize specular reflection and shadowing artifacts.
-Frames were recorded at one-minute intervals throughout colony expansion and continued until the colony front reached the plate boundary or no further measurable growth was observed.
-Sequences were stored at native spatial resolution with associated timestamps.
-For downstream analysis, each sequence was segmented using TexPol–Net to obtain pixel-level colony masks.
-These mask sequences formed a morphology-consistent representation used for training and evaluating the forecasting models.
-Training and validation datasets were assembled from non-overlapping experimental runs collected under distinct growth conditions to assess generalization across colony expansion regimes.
-4.2 Recovering front texture and geometry
-Swarming colony images present two coupled challenges for convolutional backbones.
-At the micro-structural level, boundaries are uncertain and textures vary across scales, so high-frequency evidence is easily attenuated when features are aggregated over local receptive fields.
-At the meso- to macro-structural level, growth follows a disc-like geometry with coherent radial propagation, producing long-range dependencies that are poorly captured by standard kernels.
-Addressing both phenomena with a single mechanism tends to inflate complexity or compromise either edge fidelity or global organization.
-The proposed design decouples the challenges of texture variation and geometric organization while maintaining a unified attention framework.
-The Texture-Edge Attention (TEA) module, illustrated in Figure 6a, is devised to enhance the stability and clarity of fine-scale features that are often blurred by conventional filters.
-It integrates three cooperative components: a local depthwise branch that preserves intra-channel spatial detail, a multi-dilated branch that captures textures over multiple scales, and an edge-sensitive branch initialized with a discrete Laplacian kernel to introduce a high-pass prior at the beginning of training.
-Channel-wise and spatial attention mechanisms further purify the fused representation by reducing redundancy and highlighting informative regions, while a conditional residual pathway ensures optimization stability when feature dimensions match.
-Detailed formulations and implementation details for both modules are presented in Section S2.
-The Polar–Context Attention (PCA) module, illustrated in Figure 6b, complements TEA by embedding a radial geometric prior directly into the feature representation.
-The module first compresses feature channels to eliminate redundancy and then performs a polar transformation that maps features into $(\rho,\theta)$ space, where depthwise dilated convolutions operate along radial and angular directions consistent with the colony’s outward propagation.
-A large-kernel Cartesian branch provides broad contextual coverage, while a local branch preserves fine structural fidelity.
-Channel- and spatial-attention mechanisms further refine the fused representation, enhancing the coherence between local detail and global geometry.
-A conditional residual connection stabilizes training when feature dimensions coincide.
-TEA and PCA provide complementary inductive biases: TEA sharpens edges and preserves high-frequency texture amid blurred, irregular boundaries, whereas PCA captures geometry-aligned, long-range dependencies characteristic of near–concentric–ring radial growth.
-Operating both modules at a compact internal width with shared normalization and activation yields a representation that is simultaneously edge-aware, scale-robust, and geometry-aligned, without incurring prohibitive computational cost.
-4.3 TexPol–Net
-Texture- and geometry-aware instance segmentation.
-To recover accurate colony fronts from swarming imagery, we developed the Texture-Polar Network (TexPol–Net), an instance segmentation architecture that integrates fine-scale texture encoding with geometry-aligned contextual reasoning.
-This design addresses two coupled challenges in swarming colonies: diffuse and low-contrast boundaries at the local scale, and near–concentric radial organization at the global scale.
-TexPol–Net draws inspiration from efficient one-stage detectors such as YOLO [41, 35] and the prototype-based mask formulation of YOLACT [3], while introducing texture- and geometry-aware mechanisms tailored to biological front segmentation (Fig.
-3a).
-Hierarchical backbone with complementary attentions.
-TexPol–Net employs a five-stage hierarchical convolutional backbone with progressive downsampling.
-At each stage, Texture-Edge Attention (TEA) and Polar–Context Attention (PCA) are jointly embedded to serve complementary roles.
-TEA operates predominantly on shallower representations to preserve boundary sharpness and local texture variation, whereas PCA increasingly emphasizes geometry-aligned contextual aggregation in deeper layers, enabling the capture of long-range radial propagation patterns.
-Bidirectional multi-scale feature fusion.
-The neck adopts a PANet-style bidirectional fusion strategy [25] to integrate semantic and spatial information across scales.
-Top–down and bottom–up pathways are interleaved with PCA modules, ensuring that polar consistency is preserved during resolution recovery.
-A final TEA module is applied prior to prediction to recalibrate channel responses and reinforce fine-scale texture cues.
-Dense prediction and prototype-based mask assembly.
-The prediction head employs fully convolutional branches on feature levels $P3$, $P4$, and $P5$ to predict class scores, bounding boxes, and instance-specific mask coefficients.
-Instance masks are generated using a YOLACT-style prototype formulation: a lightweight Protonet produces k=32k{=}32 shared prototypes, and each detection predicts a coefficient vector for linear combination.
-After non-maximum suppression, masks are generated, cropped, and thresholded within the retained bounding regions.
-Loss.
-During training, supervision is applied to all prediction heads in each mini-batch.
-The overall objective is defined as a weighted linear combination of the bounding box regression, classification, distribution focal loss (DFL), and mask segmentation terms:
-$\mathcal{L}_{\mathrm{seg}}=\lambda_{\mathrm{box}}\,\mathcal{L}_{\mathrm{box}}+\lambda_{\mathrm{cls}}\,\mathcal{L}_{\mathrm{cls}}+\lambda_{\mathrm{dfl}}\,\mathcal{L}_{\mathrm{dfl}}+\lambda_{\mathrm{mask}}\,\mathcal{L}_{\mathrm{mask}}.$
-The box loss penalizes geometric misalignment between predicted and ground-truth boxes, the classification loss supervises category confidence, and DFL refines localization by regularizing the underlying discrete distribution.
-The segmentation loss enforces consistency between predicted and ground-truth instance masks.
-All loss terms are combined linearly with fixed weights, which are kept constant across experiments.
-4.4 Morpher
-Mask-space forecasting.
-To model the temporal evolution of swarming colony morphology, we developed Morpher, a forecasting network that operates on segmentation masks rather than raw image intensities.
-This choice follows the biology of swarming expansion, where growth is expressed through the geometry and advancement of the colony front.
-Mask-level representation isolates occupied regions and evolving boundaries, focusing the model on front curvature and large-scale expansion while reducing sensitivity to illumination and imaging variability (Fig.
-2.5).
-Spatial encoder and latent sequence.
-Given an input mask sequence, a multi-scale encoder extracts a hierarchy of feature maps $\{f_{1}^{(t)},f_{2}^{(t)},f_{3}^{(t)},f_{4}^{(t)},f_{5}^{(t)}\}$ at each time step via progressive downsampling $(H\!\to\!H/2\!\to\!H/4\!\to\!H/8\!\to\!H/16\!\to\!H/32)$.
-In parallel, each frame is compressed into a latent descriptor $z_{t}\in\mathbb{R}^{256}$, yielding a compact temporal representation for forecasting.
-For temporal modeling, the observed segment is represented as a latent tensor of shape $(B,T_{\mathrm{obs}},256)$, where the latent dimensionality is fixed by the shared spatial encoder.
-Sinusoidal time encodings are added to the latent sequence prior to the temporal module to provide a uniform representation of temporal order.
-Autoregressive temporal evolution.
-Morpher forecasts in an autoregressive loop.
-After observing $T_{\mathrm{obs}}$ frames, the model predicts $\hat{z}_{T_{\mathrm{obs}}+1}$, decodes it into a mask, and re-encodes the prediction as the next input.
-This feedback couples future states to the geometry produced at the previous step and helps maintain boundary continuity over long horizons.
-Forecasting is evaluated under two inference regimes in matched settings.
-In the autoregressive regime, predictions are generated stepwise over the horizon, with each predicted mask fed back as the next input.
-In the parallel regime, the model predicts the full forecasting horizon conditioned on the observed segment and the terminal latent state.
-Sequence model variants with matched capacity.
-The temporal module is instantiated using one of four architectures: a vanilla RNN [13], GRU [10], LSTM [16], or a Transformer encoder [42].
-All variants share the same spatial encoder–decoder backbone and operate in the same latent space.
-Across architectures, temporal information is summarized into a single terminal latent state that serves as the sole conditioning variable for future prediction, enforcing a consistent past-to-state formulation.
-The recurrent variants use three recurrent layers, while the Transformer uses three stacked encoder blocks.
-No model introduces future tokens, decoder-side attention, or changes to the latent dimensionality.
-Morphon memory integration.
-To regulate how past morphology influences future predictions, Morphon retrieves informative historical states from the latent history via cross-attention, using a learnable query derived from the current state.
-The retrieved representation is fused into the evolving temporal state through a learnable gate $\alpha\in(0,1)$, balancing historical memory against the current predictive trajectory.
-Multi-scale decoding and feedback.
-A multi-scale decoder mirrors the encoder hierarchy to progressively restore spatial resolution while reinjecting intermediate features retained during encoding.
-This design preserves peripheral protrusions and fine-scale curvature patterns during reconstruction.
-During training, the predicted mask is re-encoded and fed back for subsequent steps, matching the inference-time autoregressive procedure.
-Sequence structuring.
-Observation length $T_{\mathrm{obs}}$ and forecasting horizon are treated as tunable hyperparameters.
-Sequences are constructed by uniformly subsampling frames with a fixed stride and partitioned into observation and prediction segments under predefined observation–prediction ratios.
-Loss.
-Supervision is applied at every predicted time step within the forecasting horizon.
-Let $T$ denote the number of predicted frames.
-The training objective is defined as a temporal average over the prediction window:
-$\mathcal{L}_{\mathrm{pred}}=\frac{1}{T}\sum_{t=1}^{T}\Big(\lambda_{\mathrm{focal}}\,\mathcal{L}_{\mathrm{focal}}^{(t)}+\lambda_{\mathrm{softIoU}}\,\mathcal{L}_{\mathrm{softIoU}}^{(t)}+\lambda_{\mathrm{boundary}}\,\mathcal{L}_{\mathrm{boundary}}^{(t)}\Big).$
-Focal loss addresses foreground–background imbalance by up-weighting hard pixels, while soft IoU enforces region-level overlap.
-A boundary-aware loss further penalizes discrepancies along the colony front; it is computed using a GPU-efficient morphological gradient approximation and evaluated at one-quarter of the original spatial resolution.
-Loss terms are averaged uniformly over the prediction horizon, and the relative weights are fixed across experiments.
-Interpretability as a dynamical model.
-By expressing each predicted morphology as a causal progression from prior front states, Morpher provides a data-driven dynamical representation of swarming expansion, enabling analysis of curvature perturbation propagation over time.
-4.5 Evaluation metrics for colony front segmentation
-We evaluated segmentation performance using four complementary metrics: mAP50:95, the Dice coefficient, the number of learnable parameters, and inference speed measured in frames per second (FPS).
-Among these, mAP50:95 serves as the primary metric, as it provides a comprehensive measure of segmentation quality across a range of localization tolerances.
-Following the COCO evaluation protocol, AP is first computed at multiple Intersection–over–Union (IoU) thresholds, and mAP50:95 is obtained by averaging across IoU thresholds ranging from 0.50 to 0.95, with a step size of 0.05:
-$\mathrm{mAP}_{50:95}=\frac{1}{10}\sum_{\tau\in\{0.50,0.55,\dots,0.95\}}\left[\frac{1}{101}\sum_{i=0}^{100}P\!\left(\tfrac{i}{100};\,\tau\right)\right],$
-where $P(r;\tau)$ denotes the interpolated precision at recall level $r$ for IoU threshold $\tau$.
-In addition to mAP, segmentation accuracy is quantified using the Dice coefficient, which measures the overlap between predicted and ground-truth masks.
-For a predicted mask $\hat{Y}$ and a ground-truth mask $Y$, the Dice coefficient is defined as
-$\mathrm{Dice}=\frac{2\,|\hat{Y}\cap Y|}{|\hat{Y}|+|Y|},$
-with values ranging from 0 to 1, where higher scores indicate better spatial agreement.
-To assess model complexity, we report the total number of learnable parameters, which indicates the representational capacity and computational burden of the segmentation network.
-To evaluate practical efficiency, we additionally measure inference throughput in FPS, defined as the number of images the model can process per second under the same hardware configuration.
-4.6 Evaluation metrics for morphological forecasting
-Forecasting accuracy must be judged not only by per–frame agreement, but also by how faithfully the predicted colony advances and organizes its growth direction over time.
-Accordingly, our evaluation considers both the spatial fidelity of the predicted masks and boundaries, and the temporal consistency of the advancing front, including its radial expansion speed and the coherence of its directional variation around the colony rim.
-All symbols and operations used in the metrics below are defined explicitly.
-To quantify mask fidelity over time, let $\widehat{Y}_{t}$ and $Y_{t}$ denote the predicted and true masks at time $t$, and let $\partial\widehat{Y}_{t}$ and $\partial Y_{t}$ denote their corresponding boundaries.
-We report the mean Intersection over Union (mIoU), defined as the temporal average of the IoU used in the AP protocol:
-$\mathrm{mIoU}=\frac{1}{T}\sum_{t=1}^{T}\frac{\lvert\widehat{Y}_{t}\cap Y_{t}\rvert}{\lvert\widehat{Y}_{t}\cup Y_{t}\rvert}$.
-To evaluate boundary placement in a manner sensitive to fine-scale morphological deviations, we use the symmetric Hausdorff distance,
-$d_{\mathrm{H}}=\frac{1}{T}\sum_{t=1}^{T}\max\Big\{\max_{y\in\partial Y_{t}}\min_{\hat{y}\in\partial\widehat{Y}_{t}}\|y-\hat{y}\|,\;\max_{\hat{y}\in\partial\widehat{Y}_{t}}\min_{y\in\partial Y_{t}}\|\hat{y}-y\|\Big\},$
-as well as its 95th–percentile variant,
-$d_{\mathrm{H}_{95}}=\frac{1}{T}\sum_{t=1}^{T}\max\Big\{P_{95}\!\big(\min_{\hat{y}\in\partial\widehat{Y}_{t}}\|y-\hat{y}\|\big)_{y\in\partial Y_{t}},\;P_{95}\!\big(\min_{y\in\partial Y_{t}}\|\hat{y}-y\|\big)_{\hat{y}\in\partial\widehat{Y}_{t}}\Big\},$
-which downweights sparse, extreme outliers arising from transient high–curvature growth.
-Finally, the average symmetric surface distance provides a measure of mean bidirectional nearest–boundary discrepancy,
-$d_{\mathrm{ASSD}}=\frac{1}{T}\sum_{t=1}^{T}\frac{1}{\lvert\partial\widehat{Y}_{t}\rvert+\lvert\partial Y_{t}\rvert}\left(\sum_{y\in\partial Y_{t}}\min_{\hat{y}\in\partial\widehat{Y}_{t}}\|y-\hat{y}\|+\sum_{\hat{y}\in\partial\widehat{Y}_{t}}\min_{y\in\partial Y_{t}}\|\hat{y}-y\|\right).$
-To evaluate whether the predicted colony reproduces the correct front propagation dynamics, we measure the radial expansion speed along $K$ uniformly sampled angular directions $\{\theta_{k}\}_{k=1}^{K}$ from the colony centroid.
-In our experiments, we set $K=720$, which corresponds to a sampling interval of $0.5^{\circ}$, which provides sufficiently fine angular resolution to capture the subtle, high–curvature perturbations that arise during dense–finger instabilities.
-Let $r(\theta_{k},t)$ and $\widehat{r}(\theta_{k},t)$ denote the ground–truth and predicted radial distances at time $t$.
-The corresponding expansion velocities are
-$v(\theta_{k},t)=\frac{r(\theta_{k},t)-r(\theta_{k},t-\Delta t)}{\Delta t},\qquad\widehat{v}(\theta_{k},t)=\frac{\widehat{r}(\theta_{k},t)-\widehat{r}(\theta_{k},t-\Delta t)}{\Delta t}.$
-The overall accuracy of front advancement is quantified by the mean root–mean–square error,
-$\mathrm{RMSE}=\frac{1}{T-1}\sum_{t=2}^{T}\sqrt{\frac{1}{K}\sum_{k=1}^{K}\big(\widehat{v}(\theta_{k},t)-v(\theta_{k},t)\big)^{2}}.$
-Beyond matching the average speed, we also assess whether the temporal fluctuation strength of the radial velocity field is preserved.
-To this end, we compute the Temporal Consistency Index (TCI) over sliding windows $w=1,\dots,W$ of fixed length $L=4$ radius frames, corresponding to three consecutive velocity steps.
-For each window and direction, we compare the temporal standard deviation of the predicted velocity trace, denoted $\sigma^{(w)}_{\widehat{v},k}$, with that of the ground–truth velocity trace, denoted $\sigma^{(w)}_{v,k}$.
-The directional consistency score is
-$\mathrm{TCI}^{(w)}_{k}=1-\frac{\big|\sigma^{(w)}_{\widehat{v},k}-\sigma^{(w)}_{v,k}\big|}{\sigma^{(w)}_{\widehat{v},k}+\sigma^{(w)}_{v,k}+\varepsilon},$
-and is evaluated only when $\sigma^{(w)}_{\widehat{v},k}+\sigma^{(w)}_{v,k}>\tau_{0}$, where $\tau_{0}=10^{-6}$ is a numerical tolerance that filters out angular directions with effectively no front motion, ensuring that TCI reflects meaningful temporal fluctuations rather than numerical artefacts.
-The final index aggregates all valid scores across windows and directions:
-$\mathrm{TCI}=\frac{1}{W}\sum_{w=1}^{W}\frac{1}{\lvert\mathcal{K}_{w}\rvert}\sum_{k\in\mathcal{K}_{w}}\mathrm{TCI}^{(w)}_{k}\in[0,1],$
-where $\mathcal{K}_{w}$ denotes the set of valid directions in window $w$; higher values indicate better preservation of intrinsic fluctuation dynamics.
-To evaluate the organization of growth across angles, we examine the anisotropy of the radial velocity field.
-The normalized angular spread (NAS) is computed as the angular standard deviation divided by the angular mean of the velocity at each time point, where $\mathrm{NAS}_{t}$ refers to the value computed from the ground–truth velocity field and $\widehat{\mathrm{NAS}}_{t}$ refers to the value computed from the predicted one.
-We report their mean absolute deviation over time:
-$\lvert\Delta\mathrm{NAS}\rvert=\frac{1}{T-1}\sum_{t=2}^{T}\big|\widehat{\mathrm{NAS}}_{t}-\mathrm{NAS}_{t}\big|.$
-To further characterize directional patterning, we examine the angular Fourier spectrum of $v(\theta,t)$ and extract the normalized second–harmonic power,
-$\mathrm{H}_{2,t}=\frac{\big|\mathcal{F}_{\theta}\{v(\theta,t)\}[2]\big|^{2}}{\sum_{m}\big|\mathcal{F}_{\theta}\{v(\theta,t)\}[m]\big|^{2}},\qquad\widehat{\mathrm{H}}_{2,t}=\frac{\big|\mathcal{F}_{\theta}\{\widehat{v}(\theta,t)\}[2]\big|^{2}}{\sum_{m}\big|\mathcal{F}_{\theta}\{\widehat{v}(\theta,t)\}[m]\big|^{2}},$
-This quantity reflects the strength of the characteristic two–lobed angular mode. We define the Second–Harmonic Power Deviation as
-$\lvert\Delta\mathrm{H}_{2}\rvert=\frac{1}{T-1}\sum_{t=2}^{T}\big|\widehat{\mathrm{H}}_{2,t}-\mathrm{H}_{2,t}\big|.$
-4.7 Implementation of TexPol–Net
-Framework and experimental setup.
-All models are trained using the Ultralytics YOLO framework, with the task configured for image segmentation.
-All experiments are conducted under identical training configurations and independently repeated to ensure fair comparability.
-Training and optimization protocol.
-For training, the maximum number of epochs is set to 300, and an early stopping strategy is employed to mitigate overfitting once validation performance saturates.
-The batch size is set to 16, and all input images are resized to a fixed resolution of $640\times 640$, balancing computational efficiency and memory usage.
-Automatic mixed-precision (AMP) training is enabled throughout to improve training throughput and reduce memory consumption while maintaining numerical stability.
-The optimizer is set to stochastic gradient descent (SGD).
-The initial learning rate is set to $6\times 10^{-3}$ and is combined with a linear warm-up schedule over the first three epochs.
-After warm-up, the learning rate is gradually decayed to $1\%$ of its initial value.
-A weight decay of $5\times 10^{-4}$ and a momentum of $0.937$ are used to stabilize gradient updates and improve generalization.
-All experiments use a fixed random seed and enable deterministic training to ensure reproducibility.
-4.8 Implementation of Morpher
-Sequence construction and data preprocessing.
-Morpher operates exclusively on binary colony masks generated by a pretrained segmentation model and does not directly access raw image intensities.
-In our primary pipeline, TexPol–Net was used to generate these masks, while alternative segmentation models were used in comparative experiments.
-Each training sample therefore consisted of a temporally ordered sequence of segmentation masks representing colony occupancy and front geometry.
-Sequences were constructed by uniformly subsampling frames from the full temporal series using a fixed stride, yielding equal temporal spacing between adjacent frames, and the resulting sequence length $T$ was defined by this fixed-stride subsampling rule.
-Each sequence was partitioned into an observation segment and a prediction segment using observation–prediction ratios of $0.5/0.5$, $0.6/0.4$, $0.7/0.3$, $0.8/0.2$, and $0.9/0.1$.
-Within each experiment, all masks were generated using the same segmentation model to ensure a consistent morphological representation across both segmentation and forecasting stages.
-All masks were resized to a spatial resolution of $640\times 640$, which was used consistently across all experiments.
-The dataset was split into training and validation partitions at the growth-sequence level, with no colony contributing sequences to both partitions, and all quantitative results are reported on the validation split.
-Training and optimization protocol.
-All variants were optimized with AdamW using an initial learning rate of $5\times 10^{-5}$ and weight decay of $1\times 10^{-4}$.
-Training was conducted for 300 epochs with a batch size of 2.
-All input mask sequences were resized to a fixed spatial resolution of $640\times 640$.
-A linear warm-up was applied over the first 10% of optimization steps, followed by cosine annealing.
-The global gradient norm was clipped to 1.0.
-Mixed-precision training was enabled throughout via automatic mixed precision with gradient scaling to improve computational throughput while maintaining numerical stability.
-Validation was performed at every epoch, and model selection was based on the checkpoint achieving the highest validation mIoU.
+SM3 was transferred from an LB agar plate into LB broth (10 g/L tryptone, 5 g/L yeast extract, and 5 g/L NaCl) and cultured overnight at $37\,^{\circ}\mathrm{C}$ with shaking at 200 rpm.
+A $5\text{--}8\,\mu\mathrm{L}$ aliquot was inoculated at the center of a freshly prepared swarming plate (LB with 0.5%, 0.6%, or 0.7% agar; plate thickness 3–4 mm), incubated at $30\,^{\circ}\mathrm{C}$ and $\sim 90\%$ relative humidity for 4–6 h to activate swarming, and then transferred to a time-lapse imaging chamber maintained at $27\,^{\circ}\mathrm{C}$, $30\,^{\circ}\mathrm{C}$, or $33\,^{\circ}\mathrm{C}$ and $86\%$, $90\%$, or $94\%$ relative humidity.
+These temperature, humidity, and agar conditions span the permissive regime of SM3 swarming [47, 11, 27] and were varied across discrete levels to probe model generalization within physiologically relevant growth states.
+Humidity was kept below the condensation threshold to avoid droplet formation on the agar surface.
+Experimental conditions were varied to generate distinct expansion regimes.
+Images were acquired with a vertically mounted high-resolution digital camera under uniform LED illumination.
+Frames were recorded every minute until the colony reached the plate boundary or no further measurable expansion was observed.
+The resulting Swarming Morphogenesis Evolution (SwarmEvo) dataset comprised 1,971 annotated images for segmentation and 276 long time series for temporal modeling.
+All recordings were stored at native resolution ($1250\times 1250$ px) with timestamps.
+For downstream analysis, each sequence was converted into boundary-resolved colony masks using TexPol–Net.
+These masks defined the state space for forecasting.
+Training and validation partitions were split at the sequence level, with no colony contributing trajectories to both partitions.
+Exact dataset composition and augmentation procedures are provided in the Supplementary Information.
+5.2 TexPol–Net for boundary-resolved state construction
+TexPol–Net was designed to recover colony fronts under two coupled conditions: diffuse, low-contrast boundaries at the local scale and near-concentric radial organization at the colony scale.
+The network therefore combines two complementary modules.
+Texture–Edge Attention (TEA) preserves fine boundary texture through local depthwise filtering, multi-scale dilated texture encoding, and an edge-sensitive high-pass prior.
+Polar–Context Attention (PCA) embeds a geometry-aligned prior by combining local features, large-kernel Cartesian context, and a polar branch operating in $(\rho,\theta)$ coordinates.
+Full module formulations and implementation details are provided in the Supplementary Information.
+As shown in Figure 3a, the full architecture follows a prototype-based one-stage instance segmentation design [52, 5].
+A five-stage hierarchical convolutional backbone progressively downsamples the input while interleaving TEA and PCA blocks, allowing fine boundary cues and geometry-aligned context to be encoded jointly.
+The resulting multi-scale features are fused by a PANet-style bidirectional neck [36, 38], in which PCA is retained during top-down and bottom-up aggregation to preserve polar consistency across resolutions.
+Prediction is performed at feature levels $P3$, $P4$, and $P5$ through dense heads for class scores, bounding boxes, and instance-specific mask coefficients.
+In parallel, a lightweight Protonet produces $k=32$ shared prototypes, which are linearly combined with the predicted coefficients and then cropped and thresholded to produce final colony masks.
+Training uses a composite segmentation objective,
+$\mathcal{L}_{\mathrm{seg}}=\lambda_{b}\mathcal{L}_{b}+\lambda_{c}\mathcal{L}_{c}+\lambda_{d}\mathcal{L}_{d}+\lambda_{m}\mathcal{L}_{m}.$
+Here $\mathcal{L}_{b}$, $\mathcal{L}_{c}$, $\mathcal{L}_{d}$, and $\mathcal{L}_{m}$ denote box, classification, distribution focal, and mask losses, respectively.
+All weights were fixed across experiments.
+5.3 Morpher for morphology forecasting
+Morpher forecasts swarming expansion in mask space rather than image space.
+This formulation follows the biology of the problem: what propagates is the front, not visual appearance.
+Each input mask sequence is encoded by a shared multi-scale spatial encoder into framewise latent descriptors $z_{t}\in\mathbb{R}^{256}$, together with intermediate feature maps used later for decoding.
+Sinusoidal temporal encodings are added before temporal modeling.
+As summarized in Figure 5a, Morpher consists of four coupled components: a multi-scale spatial encoder, a temporal sequence model, a Morphon memory block, and a multi-scale decoder.
+Observed masks are first compressed into a compact latent sequence, while encoder-side intermediate features are retained for later reconstruction.
+Future evolution is then predicted in latent space.
+In the autoregressive setting, each decoded prediction is re-encoded and fed back as the next input, so that the forecast proceeds through stepwise updates rather than one-shot generation.
+Morphon operates on the observed latent history by cross-attention with a learnable query derived from the aggregated observation state, and injects the retrieved structural memory through a learnable gate $\alpha\in(0,1)$.
+A multi-scale decoder finally reconstructs the predicted mask while reinjecting encoder features to preserve peripheral protrusions and fine curvature.
+Forecasting is performed autoregressively unless otherwise stated.
+After observing $T_{\mathrm{obs}}$ frames, the model predicts the next latent state, decodes it into a mask, re-encodes that prediction, and feeds it back for subsequent steps.
+This stepwise rollout couples future predictions to the geometry produced at the previous step and stabilizes boundary continuity over long horizons.
+Parallel prediction is used only in matched ablations.
+The temporal module is instantiated as one of four matched-capacity sequence models: vanilla RNN [16], GRU [13], LSTM [22], or Transformer encoder [59].
+All variants share the same encoder–decoder backbone, the same latent dimensionality, and the same past-to-state formulation.
+The recurrent variants use three recurrent layers; the Transformer uses three encoder blocks.
+Prediction is supervised at each forecast step by a temporally averaged objective,
+$\mathcal{L}_{\mathrm{pred}}=\frac{1}{T}\sum_{t=1}^{T}\mathcal{L}^{(t)},\quad\mathcal{L}^{(t)}=\lambda_{f}\mathcal{L}_{f}+\lambda_{i}\mathcal{L}_{i}+\lambda_{b}\mathcal{L}_{b}.$
+Here $\mathcal{L}_{f}$, $\mathcal{L}_{i}$, and $\mathcal{L}_{b}$ denote focal, soft IoU, and boundary losses, respectively.
+All weights were fixed across experiments.
+Implementation details, sequence construction, and baseline-matched comparisons are provided in the Supplementary Information.
+5.4 Training and implementation
+For TexPol–Net, all images were resized to $640\times 640$.
+Training used the Ultralytics YOLO framework with SGD ($6\times 10^{-3}$ initial learning rate, three-epoch linear warm-up, decay to 1% of the initial rate, momentum =0.937=0.937, weight decay $5\times 10^{-4}$), batch size =16=16, mixed-precision training, and early stopping once validation performance saturated.
+For Morpher, binary masks generated by a single segmentation model were uniformly subsampled with a fixed stride and partitioned into observation and prediction segments under ratios of $0.5/0.5$, $0.6/0.4$, $0.7/0.3$, $0.8/0.2$, and $0.9/0.1$.
+All masks were resized to $640\times 640$.
+Training used AdamW ($5\times 10^{-5}$ initial learning rate, weight decay $10^{-4}$), batch size =2=2, 300 epochs, 10% warm-up followed by cosine annealing, gradient clipping at 1.0, and mixed precision.
+Model selection was based on the highest validation mIoU.
 No early stopping was applied.
-All experiments were conducted with fixed random seeds and deterministic backend settings to ensure reproducibility.
-TensorFloat-32 acceleration was enabled for matrix multiplications on supported hardware, while cuDNN was configured in deterministic mode.
-4.9 Running of existing methods
-Adjustable training settings were kept aligned across methods whenever applicable.
-All images were resized to $640\times 640$.
-Models trained with epoch-based schedules were optimized for 300 epochs, while models trained with iteration-based schedules explicitly report the corresponding iteration counts.
-For all video prediction models, the batch size was fixed at 2.
-YOLOv11 and YOLOv12.
-YOLOv11 and YOLOv12 were trained and evaluated using the Ultralytics YOLO framework with the task configured for image segmentation.
-Early Stopping was enabled once validation performance saturated.
-A batch size of $16$ was used.
-AMP was enabled throughout training.
-Optimization used SGD with an initial learning rate of $6\times 10^{-3}$ and a linear warm-up over the first three epochs; the learning rate was then decayed to $1\%$ of its initial value.
-Weight decay and momentum were set to $5\times 10^{-4}$ and $0.937$, respectively.
-SAM and SAM2.
-SAM and SAM2 were fine-tuned for segmentation under a unified training protocol.
-For SAM, training was conducted using an adaptive optimizer with weight decay (initial learning rate $8\times 10^{-4}$; weight decay $1\times 10^{-4}$).
-A warm-up phase of $250$ optimization steps was applied at the beginning of training, followed by stepwise learning-rate decays at iteration steps 60,00060{,}000 and 86,66686{,}666, each with a decay factor of $1/10$.
-The model was built on the SAM ViT-B backbone, and a selective freezing strategy was adopted: the image encoder and prompt encoder were frozen, while only the mask decoder was updated during training.
-For SAM2, training was formulated as a binary segmentation task with RGB images as input and binary masks as supervision.
-Images were normalized to the range $[0,1]$ using ToTensor(), while masks were processed using nearest-neighbor interpolation to preserve label integrity.
-Masks were explicitly binarized to restrict pixel values to $\{0,1\}$.
-Optimization was performed using Adam with an initial learning rate of $1\times 10^{-4}$ and no additional learning-rate scheduling.
-A batch size of $4$ was used, and BCEWithLogitsLoss was applied directly to the raw logits.
-MAU.
-MAU was run using the official implementation.
-The model employed four recurrent layers with hidden dimension 64, convolutional filters of size 5 with stride 1 and patch size 1, and no layer normalization.
-The spatiotemporal relation size was set to 2 and the temporal decay parameter to $\tau=5$.
-Scheduled sampling was enabled, with the sampling probability linearly decayed from 1.0 to 0 over 50,000 iterations at a rate of $2\times 10^{-5}$.
-Training used the Adam optimizer with a learning rate of $5\times 10^{-4}$ and a OneCycle learning rate scheduler.
-MIM.
-MIM was run using the official implementation built on the PredRNN framework.
-The model employed four recurrent layers with hidden dimensions of 128, convolutional filters of size 5 with stride 1 and patch size 4, and no layer normalization.
-Scheduled sampling followed the same linear decay strategy as above, while reverse scheduled sampling was disabled.
-Training used the Adam optimizer with a learning rate of $1\times 10^{-4}$ and a OneCycle learning rate scheduler; incomplete batches were dropped during training.
-PredRNN and PredRNNv2.
-PredRNN-based models were run using the official implementations.
-Both models employed four recurrent layers with hidden dimensions of 128, convolutional filters of size 5 with stride 1 and patch size 2, and no layer normalization.
-Scheduled sampling was enabled with linear decay from 1.0 to 0 over 50,000 iterations at a rate of $2\times 10^{-5}$.
-Training used the Adam optimizer with a learning rate of $1\times 10^{-3}$ and a OneCycle learning rate scheduler.
-PredRNNv2 additionally enabled reverse scheduled sampling with transition steps at 25,000 and 50,000 iterations and an exponential coefficient of 5,000, and incorporated a decoupling loss with weight $\beta=0.01$.
-SimVP and SimVPv2.
-SimVP-based baselines were run using the official implementation.
-The spatial encoder–decoder employed a channel width of 64 with four convolutional blocks ($N_{S}=4$), while temporal modeling used a hidden dimension of 256 with eight temporal blocks ($N_{T}=8$).
-SimVP used TAU units for temporal prediction, whereas SimVPv2 replaced TAU with gSTA modules.
-Training used the Adam optimizer with a learning rate of $1\times 10^{-3}$ and a OneCycle learning rate scheduler.
-Model selection followed the validation loss criterion defined in the configuration.
-4.10 Swarming Morphogenesis Evolution dataset
-The Swarming Morphogenesis Evolution (SwarmEvo) dataset consists of high-resolution time-lapse recordings of Enterobacter sp.
-SM3 acquired at a fixed spatial resolution of $1250\times 1250$ pixels.
-After augmentation, the dataset comprises 1,971 annotated sequences used for training and evaluating segmentation models, as well as 276 long time series derived from continuous recordings sampled at 1-min intervals, which serve as the basis for temporal modeling and multi-scale temporal downsampling.
-Data were collected across multiple agar plates and independent imaging sessions, introducing natural variability in growth dynamics and colony morphology across distinct morphogenetic regimes.
-Segmentation masks used for model training and evaluation were obtained through a dedicated segmentation pipeline and subsequently curated to ensure consistent boundary delineation, while temporal sequences were generated by propagating these masks across time to support forecasting tasks.
-Details of cultivation conditions, imaging setup, and mask generation are provided in Section 4.1.
-Segmentation-level augmentation For training the segmentation model, augmentations were applied independently to each image–annotation pair.
-Photometric perturbations included linear intensity rescaling with offset, gamma correction, additive Gaussian noise, and sparse impulse-like pixel corruption.
-Geometric transformations were sampled per image and applied consistently to the image and its polygon annotations, including random in-plane rotation, isotropic scaling, translation constrained by the instance extent, and random horizontal or vertical flipping.
-A random cutout was further used to simulate partial occlusion; polygon annotations were updated by geometrically clipping the visible region and retaining valid connected components.
-After each transformation, polygon validity was enforced by automatic closure and self-intersection repair, and invalid or degenerate shapes were discarded.
-Sequence-level augmentation.
-Data augmentation was applied at the sequence level and restricted to spatial transformations that preserve the underlying growth dynamics.
-For each sequence, a single set of affine transformation parameters was sampled and applied identically to all frames to maintain temporal coherence.
-The augmentation pipeline was limited to in-plane rotation, translation, and random horizontal and vertical flipping.
-No augmentation was applied selectively to specific temporal segments or across time.
-Temporal resolution was defined solely by fixed-stride subsampling, without stochastic temporal perturbations.
+All experiments used fixed random seeds and deterministic backend settings; TensorFloat-32 acceleration was enabled where available.
+All baseline models were retrained under matched preprocessing and comparable train/validation splits.
+Model-specific implementation details, including optimizer schedules and architecture-specific settings for YOLOv11/12, SAM/SAM2, MAU, MIM, PredRNN, PredRNNv2, SimVP+TAU, and SimVPv2+gSTA, are provided in the Supplementary Information.
+5.5 Evaluation
+Segmentation was evaluated by mAP50:95, image-wise IoU, and Dice coefficient.
+Forecasting was evaluated at two levels.
+Spatial fidelity was assessed by mIoU, HD, HD95, and ASSD, which quantify overlap and boundary localization.
+Dynamical fidelity was assessed by radial-velocity RMSE and the Temporal Consistency Index (TCI), which measure front propagation and temporal fluctuation preservation.
+Angular organization was quantified by $|\Delta\mathrm{NAS}|$ and $|\Delta\mathrm{H}_{2}|$, which measure deviations in anisotropic growth and second-harmonic angular structure.
+In the main text, mAP50:95 serves as the primary segmentation metric, while HD95 and ASSD serve as the primary forecasting metrics because region overlap can remain high even when the contour drifts.
+Formal definitions of all metrics are provided in the Supplementary Information.
 Acknowledgements
 This work was supported by the National Natural Science Foundation of China (NSFC, Grant No. 12202275) and the Shanghai Jiao Tong University Explore X Fund.
+Conflict of Interest
+The authors declare no conflict of interest.
+Data availability
+All datasets used in this study are publicly available.
+The Swarming Morphogenesis Evolution (SwarmEvo) dataset, including both the segmentation and temporal prediction subsets, is hosted at https://huggingface.co/datasets/ShengyouDuan/SwarmEvo.
+The dataset provides polygon-based segmentation annotations and time-lapse sequences for bacterial swarming experiments, and is released to support reproducibility and further research in morphology-aware modeling.
+Code availability
+The complete implementation of the proposed framework, including TexPol–Net for morphology-aware segmentation and Morpher for autoregressive temporal forecasting, is publicly available at https://github.com/ShengyouDuan/From_shape_to_fate__making_bacterial_swarming_expansion_predictable.
+The repository contains all training and evaluation code required to reproduce the results reported in this work.
 Supplementary Information
-S1 Performance comparison with state–of–the–art video prediction models under an 80%–20% observation–prediction split.
-Table 1: Performance comparison with state–of–the–art video prediction models under an 80%–20% observation–prediction split.
-This table benchmarks Morpher against leading video prediction architectures, including MAU, MIM, PredRNN variants, and SimVP–based models.
-All methods are evaluated under identical input–output protocols for long–term forecasting of swarming colony expansion.
-Morpher achieves substantially higher region–level overlap (mIoU) and lower boundary error (HD95, ASSD), indicating improved accuracy in front propagation and boundary–level morphology.
-Model mIoU (%) $\uparrow$ HD95 (px) $\downarrow$ ASSD (px) $\downarrow$
-MAU [6] 84.67 22.73 14.68
-MIM [46] 89.32 20.17 10.30
-PredRNN [44] 84.60 22.75 14.81
-PredRNNv2 [45] 84.14 23.24 15.04
-SimVP [15]+TAU [40] 86.87 23.19 12.47
-SimVP [15]+gSTA [39] 90.52 18.28 8.87
-Morpher (Ours) 95.42 10.61 3.93
-S2 Texture-Edge Attention and Polar–Context Attention modules
+S1 Texture–Edge Attention and Polar–Context Attention modules
 Swarming colony images exhibit complex morphological organization characterized by uncertain boundaries, irregular shapes, and radially propagating texture patterns.
-These inherent properties pose significant challenges for CNNs, whose reliance on local receptive fields restricts their capacity to capture long-range spatial dependencies and global geometric structures, particularly the near–concentric–ring radial expansion typical of swarming growth.
-To overcome these limitations and enhance both fine-grained texture extraction and high-level semantic representation, we developed two specialized attention modules, the Texture-Edge Attention (TEA) and Polar–Context Attention (PCA), as illustrated in Figure 6.
-The TEA module, as shown in Figure 6LABEL:fig:TEA, is designed to address blurred boundaries and multi-scale, high-frequency texture variability.
+These inherent properties pose significant challenges for CNNs, whose reliance on local receptive fields restricts their capacity to capture long-range spatial dependencies and global geometric structures, particularly the near-concentric-ring radial expansion typical of swarming growth.
+To overcome these limitations and enhance both fine-grained texture extraction and high-level semantic representation, we developed two specialized attention modules, the Texture–Edge Attention (TEA) and Polar–Context Attention (PCA), as illustrated in Figure 7.
+The TEA module, as shown in Figure 7a, is designed to address blurred boundaries and multi-scale, high-frequency texture variability.
 It combines three cooperative paths: a local branch that preserves intra-channel spatial details, a multi-dilated path that ensures scale-robust texture encoding, and an edge-sensitive path initialized with a discrete Laplacian kernel to enhance boundary awareness.
@@
 where $\mathrm{Conv}^{\mathrm{dw}}_{3\times 3}$ denotes depthwise convolution, $\mathrm{GN}$ represents group normalization, and $\phi$ is the SiLU activation.
-A squeeze–and–excitation (SE) gate $\mathbf{f}\!\in\!\mathbb{R}^{B\times C_{h}\times 1\times 1}$ is computed via global average pooling (GAP) followed by two $1\times 1$ convolutions with nonlinearity and sigmoid activation:
+A squeeze-and-excitation (SE) gate $\mathbf{f}\!\in\!\mathbb{R}^{B\times C_{h}\times 1\times 1}$ is computed via global average pooling (GAP) followed by two $1\times 1$ convolutions with nonlinearity and sigmoid activation:
 $\mathbf{f}=\sigma\!\Big(\mathrm{Conv}_{1\times 1}\big(\phi(\mathrm{Conv}_{1\times 1}(\mathrm{GAP}(\mathbf{X})))\big)\Big).$ (3)
@@
 $\mathbf{F}^{(d_{k})}_{\mathrm{tex}}=\phi\!\left(\mathrm{GN}\!\left(\mathrm{Conv}_{1\times 1}\!\Big(\mathrm{Conv}^{\mathrm{dw}}_{3\times 3,d_{k}}(\mathbf{X})\Big)\right)\right),$ (5)
-Figure 6: Texture-Edge Attention (TEA) and Polar–Context Attention (PCA) modules.
+Figure 7: Texture–Edge Attention (TEA) and Polar–Context Attention (PCA) modules.
 a, The TEA block enhances fine-scale texture fidelity and boundary sharpness through three cooperative branches: a local depthwise path for intra-channel spatial preservation, multi-dilated convolutions for scale-robust texture encoding, and an edge-sensitive Laplacian path that injects a high-pass prior.
@@
 Input features are first compressed and then processed by a local branch, a large-kernel Cartesian branch, and a polar-warped branch operating in $(\rho,\theta)$ coordinates.
-Depthwise dilated filters extract context along radial and angular axes, and subsequent channel– and spatial–attention gates yield a geometry-aligned output.
-and concatenation of all paths yields
+Depthwise dilated filters extract context along radial and angular axes, and subsequent channel- and spatial-attention gates yield a geometry-aligned output.
+The concatenation of all paths yields
 $\mathbf{F}_{\mathrm{tex}}=\mathrm{Concat}\big(\mathbf{F}^{(d_{1})}_{\mathrm{tex}},\mathbf{F}^{(d_{2})}_{\mathrm{tex}},\ldots,\mathbf{F}^{(d_{K})}_{\mathrm{tex}}\big)$, (6)
-where K=3K{=}3 captures short–, medium–, and long–range textures.
+where K=3K{=}3 captures short-, medium-, and long-range textures.
 An edge-aware branch initialized by the Laplacian kernel enhances boundary sensitivity:
@@
 where $\boldsymbol{\gamma}\!\in\!\mathbb{R}^{C_{\mathrm{out}}}$ is a learnable scaling factor that ensures stability when the dimensions are the same.
-While TEA focuses on boundary and texture fidelity, the PCA module illustrated in Figure 6LABEL:fig:PCA captures long-range dependencies and radial geometric organization inherent to swarming colonies.
-Conventional convolutional operators struggle with near–concentric–ring propagation whereas PCA embeds a polar-aware representation aligned with colony growth.
+While TEA focuses on boundary and texture fidelity, the PCA module illustrated in Figure 7b captures long-range dependencies and radial geometric organization inherent to swarming colonies.
+Conventional convolutional operators struggle with near-concentric-ring propagation whereas PCA embeds a polar-aware representation aligned with colony growth.
 The module comprises three paths: a local branch for spatial detail, a large-kernel branch for contextual encoding, and a polar branch that transforms features into polar coordinates for radial modeling.
-The local branch follows the same $3\times 3$ depthwise–pointwise pattern described in Eq.
+The local branch follows the same $3\times 3$ depthwise-pointwise pattern described in Eq.
 2, operating on $\mathbf{X}^{\prime}$:
@@
 This design preserves fine structural details, integrates global context, and explicitly embeds radial priors, enabling robust modeling of colony expansion dynamics.
-S3 Performance of Morpher under an 80%–20% observation–prediction split across temporal modeling and inference
-paradigms
+S2 Swarming Morphogenesis Evolution dataset
+The Swarming Morphogenesis Evolution (SwarmEvo) dataset consists of high-resolution time-lapse recordings of Enterobacter sp.
+SM3 acquired at a fixed spatial resolution of $1250\times 1250$ px.
+Swarming colonies were initiated by inoculating a $5\text{--}8\,\mu\mathrm{L}$ aliquot from an overnight culture (grown at $37\,^{\circ}\mathrm{C}$, 200 rpm in LB broth) onto the center of freshly prepared LB agar plates with concentrations of 0.5%, 0.6%, or 0.7% (plate thickness 3–4 mm).
+Plates were first incubated at $30\,^{\circ}\mathrm{C}$ and approximately 90% relative humidity for 4–6 h to activate swarming, and subsequently transferred to a time-lapse imaging chamber with controlled environmental conditions.
+During imaging, temperature was set to $27\,^{\circ}\mathrm{C}$, $30\,^{\circ}\mathrm{C}$, or $33\,^{\circ}\mathrm{C}$, and relative humidity to 86%, 90%, or 94%, spanning the permissive regime of SM3 swarming.
+Humidity was maintained below the condensation threshold to prevent droplet formation on the agar surface.
+These controlled variations in agar concentration, temperature, and humidity produced distinct expansion regimes and were used to probe model generalization across physiologically relevant growth states.
+After augmentation, the dataset comprises 1,971 annotated samples used for training and evaluating segmentation models, as well as 276 long time series derived from continuous recordings sampled at 1-min intervals, which serve as the basis for temporal modeling and multi-scale temporal downsampling.
+Data were collected across multiple agar plates and independent imaging sessions, introducing natural variability in growth dynamics and colony morphology.
+Segmentation masks used for model training and evaluation were obtained through a dedicated segmentation pipeline and subsequently curated to ensure consistent boundary delineation, while temporal sequences were generated by propagating these masks across time to support forecasting tasks.
+Segmentation-level augmentation.
+For training the segmentation model, augmentations were applied independently to each image–annotation pair.
+Photometric perturbations included linear intensity rescaling with offset, gamma correction, additive Gaussian noise, and sparse impulse-like pixel corruption.
+Geometric transformations were sampled per image and applied consistently to the image and its polygon annotations, including random in-plane rotation, isotropic scaling, translation constrained by the instance extent, and random horizontal or vertical flipping.
+A random cutout was further used to simulate partial occlusion; polygon annotations were updated by geometrically clipping the visible region and retaining valid connected components.
+After each transformation, polygon validity was enforced by automatic closure and self-intersection repair, and invalid or degenerate shapes were discarded.
+Sequence-level augmentation.
+Data augmentation was applied at the sequence level and restricted to spatial transformations that preserve the underlying growth dynamics.
+For each sequence, a single set of affine transformation parameters was sampled and applied identically to all frames to maintain temporal coherence.
+The augmentation pipeline was limited to in-plane rotation, translation, and random horizontal and vertical flipping.
+No augmentation was applied selectively to specific temporal segments or across time.
+Temporal resolution was defined solely by fixed-stride subsampling, without stochastic temporal perturbations.
+S3 Implementation of TexPol–Net
+Framework and experimental setup.
+All models were trained using the Ultralytics YOLO framework, with the task configured for image segmentation.
+All experiments were conducted under identical training configurations and independently repeated to ensure fair comparability.
+Training and optimization protocol.
+For training, the maximum number of epochs was set to 300, and an early stopping strategy was employed to mitigate overfitting once validation performance saturated.
+The batch size was set to 16, and all input images were resized to a fixed resolution of $640\times 640$, balancing computational efficiency and memory usage.
+Automatic mixed-precision training was enabled throughout to improve training throughput and reduce memory consumption while maintaining numerical stability.
+Optimization used stochastic gradient descent with an initial learning rate of $6\times 10^{-3}$, combined with a linear warm-up schedule over the first three epochs.
+After warm-up, the learning rate was gradually decayed to $1\%$ of its initial value.
+Weight decay and momentum were set to $5\times 10^{-4}$ and $0.937$, respectively.
+All experiments used a fixed random seed and deterministic training settings to ensure reproducibility.
+S4 Implementation of Morpher
+Sequence construction and data preprocessing.
+Morpher operated exclusively on binary colony masks generated by a pretrained segmentation model and did not directly access raw image intensities.
+In the primary pipeline, TexPol–Net was used to generate these masks, while alternative segmentation models were used in comparative experiments.
+Each training sample therefore consisted of a temporally ordered sequence of segmentation masks representing colony occupancy and front geometry.
+Sequences were constructed by uniformly subsampling frames from the full temporal series using a fixed stride, yielding equal temporal spacing between adjacent frames, and the resulting sequence length $T$ was defined by this fixed-stride subsampling rule.
+Each sequence was partitioned into an observation segment and a prediction segment using observation–prediction ratios of $0.5/0.5$, $0.6/0.4$, $0.7/0.3$, $0.8/0.2$, and $0.9/0.1$.
+Within each experiment, all masks were generated using the same segmentation model to ensure a consistent morphological representation across both segmentation and forecasting stages.
+All masks were resized to a spatial resolution of $640\times 640$, which was used consistently across all experiments.
+The dataset was split into training and validation partitions at the growth-sequence level, with no colony contributing sequences to both partitions, and all quantitative results were reported on the validation split.
+Training and optimization protocol.
+All variants were optimized with AdamW using an initial learning rate of $5\times 10^{-5}$ and weight decay of $10^{-4}$.
+Training was conducted for 300 epochs with a batch size of 2.
+A linear warm-up was applied over the first $10\%$ of optimization steps, followed by cosine annealing.
+The global gradient norm was clipped to 1.0.
+Mixed-precision training was enabled throughout via automatic mixed precision with gradient scaling to improve computational throughput while maintaining numerical stability.
+Validation was performed at every epoch, and model selection was based on the checkpoint achieving the highest validation mIoU.
+No early stopping was applied.
+All experiments were conducted with fixed random seeds and deterministic backend settings to ensure reproducibility.
+TensorFloat-32 acceleration was enabled for matrix multiplications on supported hardware, while cuDNN was configured in deterministic mode.
+S5 Running of existing methods
+Adjustable training settings were kept aligned across methods whenever applicable.
+All images were resized to $640\times 640$.
+Models trained with epoch-based schedules were optimized for 300 epochs, while models trained with iteration-based schedules explicitly report the corresponding iteration counts.
+For all video prediction models, the batch size was fixed at 2.
+YOLOv11 and YOLOv12.
+YOLOv11 and YOLOv12 were trained and evaluated using the Ultralytics YOLO framework with the task configured for image segmentation.
+Early stopping was enabled once validation performance saturated.
+A batch size of 16 was used.
+AMP was enabled throughout training.
+Optimization used SGD with an initial learning rate of $6\times 10^{-3}$ and a linear warm-up over the first three epochs; the learning rate was then decayed to $1\%$ of its initial value.
+Weight decay and momentum were set to $5\times 10^{-4}$ and $0.937$, respectively.
+SAM and SAM2.
+SAM and SAM2 were fine-tuned for segmentation under a unified training protocol.
+For SAM, training used AdamW with an initial learning rate of $8\times 10^{-4}$ and weight decay $10^{-4}$.
+A warm-up phase of 250 optimization steps was applied at the beginning of training, followed by stepwise learning-rate decays at 60,000 and 86,666 iterations, each with a decay factor of $1/10$.
+The model was built on the SAM ViT-B backbone, and a selective freezing strategy was adopted: the image encoder and prompt encoder were frozen, while only the mask decoder was updated during training.
+For SAM2, training was formulated as a binary segmentation task with RGB images as input and binary masks as supervision.
+Images were converted to tensors and normalized to $[0,1]$, while masks were resized using nearest-neighbor interpolation and explicitly binarized.
+Optimization used Adam with an initial learning rate of $10^{-4}$, batch size 4, and binary cross-entropy loss with logits, without an additional learning-rate schedule.
+MAU.
+MAU was run using the official implementation.
+The model employed four recurrent layers with hidden dimension 64, convolutional filters of size 5 with stride 1 and patch size 1, and no layer normalization.
+The spatiotemporal relation size was set to 2 and the temporal decay parameter to $\tau=5$.
+Scheduled sampling was enabled, with the sampling probability linearly decayed from 1.0 to 0 over 50,000 iterations at a rate of $2\times 10^{-5}$.
+Training used Adam with a learning rate of $5\times 10^{-4}$ and a OneCycle learning-rate scheduler.
+MIM.
+MIM was run using the official implementation built on the PredRNN framework.
+The model employed four recurrent layers with hidden dimensions of 128, convolutional filters of size 5 with stride 1 and patch size 4, and no layer normalization.
+Scheduled sampling followed the same linear decay strategy as above, while reverse scheduled sampling was disabled.
+Training used Adam with a learning rate of $10^{-4}$ and a OneCycle learning-rate scheduler; incomplete batches were dropped during training.
+PredRNN and PredRNNv2.
+PredRNN-based models were run using the official implementations.
+Both models employed four recurrent layers with hidden dimensions of 128, convolutional filters of size 5 with stride 1 and patch size 2, and no layer normalization.
+Scheduled sampling was enabled with linear decay from 1.0 to 0 over 50,000 iterations at a rate of $2\times 10^{-5}$.
+Training used Adam with a learning rate of $10^{-3}$ and a OneCycle learning-rate scheduler.
+PredRNNv2 additionally enabled reverse scheduled sampling with transition steps at 25,000 and 50,000 iterations and an exponential coefficient of 5,000, and incorporated a decoupling loss with weight $\beta=0.01$.
+SimVP and SimVPv2.
+SimVP-based baselines were run using the official implementation.
+The spatial encoder–decoder employed a channel width of 64 with four convolutional blocks ($N_{S}=4$), while temporal modeling used a hidden dimension of 256 with eight temporal blocks ($N_{T}=8$).
+SimVP used TAU units for temporal prediction, whereas SimVPv2 replaced TAU with gSTA modules.
+Training used Adam with a learning rate of $10^{-3}$ and a OneCycle learning-rate scheduler.
+Model selection followed the validation loss criterion defined in the configuration.
+S6 Evaluation metrics for colony front segmentation
+Segmentation performance was evaluated using three complementary metrics: mAP50:95, image-wise IoU, and Dice coefficient.
+mAP50:95 served as the primary segmentation metric because it summarizes performance across a range of localization tolerances.
+Following the COCO protocol,
+$\mathrm{mAP}_{50:95}=\frac{1}{10}\sum_{\tau\in\{0.50,0.55,\dots,0.95\}}\left[\frac{1}{101}\sum_{i=0}^{100}P\!\left(\frac{i}{100};\tau\right)\right],$ (24)
+where $P(r;\tau)$ denotes the interpolated precision at recall level $r$ under IoU threshold $\tau$. Dice was computed as
+$\mathrm{Dice}=\frac{2\,|\hat{Y}\cap Y|}{|\hat{Y}|+|Y|},$ (25)
+where $\hat{Y}$ and $Y$ denote predicted and ground-truth masks.
+S7 Evaluation metrics for morphological forecasting
+Forecasting accuracy must be judged not only by per-frame agreement, but also by how faithfully the predicted colony advances and organizes its growth direction over time.
+Accordingly, evaluation considered both the spatial fidelity of the predicted masks and boundaries, and the temporal consistency of the advancing front, including its radial expansion speed and the coherence of its directional variation around the colony rim.
+To quantify mask fidelity over time, let $\widehat{Y}_{t}$ and $Y_{t}$ denote the predicted and true masks at time $t$, and let $\partial\widehat{Y}_{t}$ and $\partial Y_{t}$ denote their corresponding boundaries.
+The mean Intersection over Union was defined as
+$\mathrm{mIoU}=\frac{1}{T}\sum_{t=1}^{T}\frac{|\widehat{Y}_{t}\cap Y_{t}|}{|\widehat{Y}_{t}\cup Y_{t}|}.$ (26)
+Boundary placement was evaluated using the symmetric Hausdorff distance,
+$d_{\mathrm{H}}=\frac{1}{T}\sum_{t=1}^{T}\max\Big\{\max_{y\in\partial Y_{t}}\min_{\hat{y}\in\partial\widehat{Y}_{t}}\|y-\hat{y}\|,\;\max_{\hat{y}\in\partial\widehat{Y}_{t}}\min_{y\in\partial Y_{t}}\|\hat{y}-y\|\Big\},$ (27)
+its 95th-percentile variant,
+$d_{\mathrm{H}_{95}}=\frac{1}{T}\sum_{t=1}^{T}\max\Big\{P_{95}\!\big(\min_{\hat{y}\in\partial\widehat{Y}_{t}}\|y-\hat{y}\|\big)_{y\in\partial Y_{t}},\;P_{95}\!\big(\min_{y\in\partial Y_{t}}\|\hat{y}-y\|\big)_{\hat{y}\in\partial\widehat{Y}_{t}}\Big\},$ (28)
+and the average symmetric surface distance,
+$d_{\mathrm{ASSD}}=\frac{1}{T}\sum_{t=1}^{T}\frac{1}{|\partial\widehat{Y}_{t}|+|\partial Y_{t}|}\left(\sum_{y\in\partial Y_{t}}\min_{\hat{y}\in\partial\widehat{Y}_{t}}\|y-\hat{y}\|+\sum_{\hat{y}\in\partial\widehat{Y}_{t}}\min_{y\in\partial Y_{t}}\|\hat{y}-y\|\right).$ (29)
+To evaluate whether the predicted colony reproduced the correct front propagation dynamics, radial expansion speed was measured along $K$ uniformly sampled angular directions $\{\theta_{k}\}_{k=1}^{K}$ from the colony centroid.
+In the experiments, $K=720$, corresponding to a sampling interval of $0.5^{\circ}$.
+Let $r(\theta_{k},t)$ and $\widehat{r}(\theta_{k},t)$ denote the ground-truth and predicted radial distances at time $t$.
+The corresponding expansion velocities were
+$v(\theta_{k},t)=\frac{r(\theta_{k},t)-r(\theta_{k},t-\Delta t)}{\Delta t},\qquad\widehat{v}(\theta_{k},t)=\frac{\widehat{r}(\theta_{k},t)-\widehat{r}(\theta_{k},t-\Delta t)}{\Delta t}.$ (30)
+The overall accuracy of front advancement was quantified by
+$\mathrm{RMSE}=\frac{1}{T-1}\sum_{t=2}^{T}\sqrt{\frac{1}{K}\sum_{k=1}^{K}\big(\widehat{v}(\theta_{k},t)-v(\theta_{k},t)\big)^{2}}.$ (31)
+Temporal fluctuation preservation was quantified by the Temporal Consistency Index (TCI) over sliding windows $w=1,\dots,W$ of fixed length $L=4$ radius frames, corresponding to three consecutive velocity steps.
+For each window and direction, $\sigma^{(w)}_{\widehat{v},k}$ and $\sigma^{(w)}_{v,k}$ denote the temporal standard deviations of predicted and ground-truth velocity traces.
+The directional consistency score was
+$\mathrm{TCI}^{(w)}_{k}=1-\frac{\big|\sigma^{(w)}_{\widehat{v},k}-\sigma^{(w)}_{v,k}\big|}{\sigma^{(w)}_{\widehat{v},k}+\sigma^{(w)}_{v,k}+\varepsilon},$ (32)
+and was evaluated only when $\sigma^{(w)}_{\widehat{v},k}+\sigma^{(w)}_{v,k}>\tau_{0}$, where $\tau_{0}=10^{-6}$ filters directions with effectively no motion. The final index was
+$\mathrm{TCI}=\frac{1}{W}\sum_{w=1}^{W}\frac{1}{|\mathcal{K}_{w}|}\sum_{k\in\mathcal{K}_{w}}\mathrm{TCI}^{(w)}_{k}$, (33)
+where $\mathcal{K}_{w}$ denotes the set of valid directions in window $w$.
+To evaluate the organization of growth across angles, the normalized angular spread (NAS) was computed from the angular standard deviation divided by the angular mean of the velocity field.
+We report the mean absolute deviation over time,
+$|\Delta\mathrm{NAS}|=\frac{1}{T-1}\sum_{t=2}^{T}\big|\widehat{\mathrm{NAS}}_{t}-\mathrm{NAS}_{t}\big|.$ (34)
+To further characterize directional patterning, we examined the angular Fourier spectrum of $v(\theta,t)$ and extracted the normalized second-harmonic power,
+$\mathrm{H}_{2,t}\mathrm{H}_{2,t}=\frac{\big|\mathcal{F}_{\theta}\{v(\theta,t)\}[2]\big|^{2}}{\sum_{m}\big|\mathcal{F}_{\theta}\{v(\theta,t)\}[m]\big|^{2}},\qquad\widehat{\mathrm{H}}_{2,t}=\frac{\big|\mathcal{F}_{\theta}\{\widehat{v}(\theta,t)\}[2]\big|^{2}}{\sum_{m}\big|\mathcal{F}_{\theta}\{\widehat{v}(\theta,t)\}[m]\big|^{2}}.$ (35)
+The corresponding deviation was defined as
+$|\Delta\mathrm{H}_{2}|=\frac{1}{T-1}\sum_{t=2}^{T}\big|\widehat{\mathrm{H}}_{2,t}-\mathrm{H}_{2,t}\big|.$ (36)
+S8 Performance comparison with state-of-the-art video prediction models under an 80%–20% observation–prediction split
+Table 1: Performance comparison with state-of-the-art video prediction models under an 80%–20% observation–prediction split.
+This table benchmarks Morpher against leading video prediction architectures, including MAU, MIM, PredRNN variants, and SimVP-based models.
+All methods are evaluated under identical input–output protocols for long-term forecasting of swarming colony expansion.
+Morpher achieves substantially higher region-level overlap (mIoU) and lower boundary error (HD95, ASSD), indicating improved accuracy in front propagation and boundary-level morphology.
+Model mIoU (%) $\uparrow$ HD95 (px) $\downarrow$ ASSD (px) $\downarrow$
+MAU 84.67 22.73 14.68
+MIM 89.32 20.17 10.30
+PredRNN 84.60 22.75 14.81
+PredRNNv2 84.14 23.24 15.04
+SimVP+TAU 86.87 23.19 12.47
+SimVP+gSTA 90.52 18.28 8.87
+Morpher (Ours) 95.42 10.61 3.93
+S9 Performance of Morpher under an 80%–20% observation–prediction split across temporal modeling and inference paradigms
 Table 2: Performance of Morpher under an 80%–20% observation–prediction split across temporal modeling and inference paradigms.
-Forecasting accuracy is evaluated across region-level overlap (mIoU), boundary accuracy (HD, HD95, ASSD), front–propagation dynamics (RMSE), temporal fluctuation consistency (TCI), and angular growth organization ($\lvert\Delta\mathrm{NAS}\rvert$, $\lvert\Delta\mathrm{H}_{2}\rvert$).
+Forecasting accuracy is evaluated across region-level overlap (mIoU), boundary accuracy (HD, HD95, ASSD), front-propagation dynamics (RMSE), temporal fluctuation consistency (TCI), and angular growth organization ($\lvert\Delta\mathrm{NAS}\rvert$, $\lvert\Delta\mathrm{H}_{2}\rvert$).
 Higher mIoU and TCI indicate superior forecasting performance, whereas lower HD-based distances, RMSE, $\lvert\Delta\mathrm{NAS}\rvert$, and $\lvert\Delta\mathrm{H}_{2}\rvert$ reflect improved geometric and dynamical fidelity.
@@
 Transformer Parallel 93.94 17.24 12.64 5.23 2.95 56.94 16.83 1.94
-RNN Autoregr. 93.55 17.53 12.51 5.26 2.20 65.63 15.02 1.90
-LSTM Autoregr. 94.07 16.92 11.85 5.34 2.60 63.14 17.80 1.91
-GRU Autoregr. 94.20 17.25 12.03 5.01 2.39 64.10 19.02 1.84
-Transformer Autoregr. 94.16 16.56 11.95 5.26 2.66 64.92 13.51 1.74
+RNN Autoregressive 93.55 17.53 12.51 5.26 2.20 65.63 15.02 1.90
+LSTM Autoregressive 94.07 16.92 11.85 5.34 2.60 63.14 17.80 1.91
+GRU Autoregressive 94.20 17.25 12.03 5.01 2.39 64.10 19.02 1.84
+Transformer Autoregressive 94.16 16.56 11.95 5.26 2.66 64.92 13.51 1.74
 RNN Parallel 94.22 16.67 11.63 4.85 2.85 61.57 17.57 1.84
@@
 Transformer Parallel 94.80 15.79 11.22 4.63 2.55 62.71 16.30 1.89
-RNN Autoregr. 94.94 15.46 10.67 4.34 2.19 63.32 16.87 1.86
-LSTM Autoregr. 95.01 15.32 10.77 4.20 2.26 65.32 15.03 1.79
-GRU Autoregr. 95.29 15.01 10.14 4.06 2.06 64.31 15.45 1.89
-Transformer Autoregr. 95.42 15.26 10.61 3.93 2.12 64.26 13.13 1.85
-S4 Performance of Morpher under a series of observation–prediction splits across sequence models
+RNN Autoregressive 94.94 15.46 10.67 4.34 2.19 63.32 16.87 1.86
+LSTM Autoregressive 95.01 15.32 10.77 4.20 2.26 65.32 15.03 1.79
+GRU Autoregressive 95.29 15.01 10.14 4.06 2.06 64.31 15.45 1.89
+Transformer Autoregressive 95.42 15.26 10.61 3.93 2.12 64.26 13.13 1.85
+S10 Performance of Morpher under a series of observation–prediction splits across sequence models
 Table 3: Performance of Morpher under a series of observation–prediction splits across sequence models.
 Results are reported for 50%, 60%, 70%, 80%, and 90% observation levels to assess how forecasting stability changes as more of the past is revealed.
-Metrics include region-level overlap (mIoU), boundary accuracy (HD, HD95, ASSD), front–propagation dynamics (RMSE), temporal fluctuation consistency (TCI), and angular growth organization ($\lvert\Delta\mathrm{NAS}\rvert$, $\lvert\Delta\mathrm{H}_{2}\rvert$).
+Metrics include region-level overlap (mIoU), boundary accuracy (HD, HD95, ASSD), front-propagation dynamics (RMSE), temporal fluctuation consistency (TCI), and angular growth organization ($\lvert\Delta\mathrm{NAS}\rvert$, $\lvert\Delta\mathrm{H}_{2}\rvert$).
 Higher mIoU and TCI indicate superior forecasting performance, whereas lower HD-based distances, RMSE, $\lvert\Delta\mathrm{NAS}\rvert$, and $\lvert\Delta\mathrm{H}_{2}\rvert$ reflect improved geometric and dynamical fidelity.